### v1

# Should You Use Your Large Language Model to Explore or Exploit?

Abstract
We evaluate the ability of the current generation of large language models (LLMs) to help a decision-making agent facing an exploration-exploitation tradeoff. We use LLMs to explore and exploit in silos in various (contextual) bandit tasks. We find that while the current LLMs often struggle to exploit, in-context mitigations may be used to substantially improve performance for small-scale tasks. However even then, LLMs perform worse than a simple linear regression. On the other hand, we find that LLMs do help at exploring large action spaces with inherent semantics, by suggesting suitable candidates to explore.
1 Introduction
There has been significant interest in the machine learning community to apply recent advances in generative AI and large language models (LLMs) to solve important decision-making problems. Early work in this direction has already produced impressive agentic behavior in both virtual [e.g.,  41, 32] and physical-world environments [e.g.,  7]. However, significant challenges remain.
Beyond generalization (needed for supervised learning), decision-making under uncertainty requires two additional capabilities: exploitation (making the best decision given the current data) and exploration (trying new options for long-term benefit). Balancing the two has been a subject of an enormous literature, see books [e.g.,  see books 38, 25, 3].
A recent line of work [e.g.,  23, 31] evaluates the ability of LLMs to balance exploration and exploitation entirely in-context, i.e.,  by specifying the problem description, parameters, and history in the LLM prompt. Focused on simple tasks in reinforcement learning, these results have been mixed. Both papers show that LLMs fail to solve these tasks adequately out-of-the-box, but they can be prompted to do so, e.g. by providing various summary statistics in-context, or by fine-tuning on data from algorithmic baselines on similar problem instances. Unfortunately, such mitigations do not readily extend beyond simple settings. For example, succinct summary statistics do not exist for many decision-making tasks, and fine-tuning for specific tasks may be prohibitive due to cost or lack of sufficient training data.
Motivated by these observations, we study the ability of Gpt-4 [2], Gpt-4o [18], and Gpt-3.5 [8] to explore and exploit in-context in silos, with an eye towards leveraging a pre-trained LLM (and the inductive bias therein) as a part of a larger decision-making agent. We focus on (contextual) bandits, as a standard abstraction for the explore-exploit tradeoff.
In Section 3, we evaluate LLMs as exploitation oracles for contextual bandits. Given a history of (context, action, reward) tuples, the LLM is tasked with identifying the best action to take given a new context. Our results here are mixed. We show that LLMs can effectively exploit in-context for small-sized problems, but their performance degrades when the problem becomes moderately sized. We find that in-context summary techniques are useful for improving performance, but LLMs with these mitigations still perform worse than a simple linear regression baseline.
In Section 4, we evaluate LLMs as exploration oracles which suggest candidate actions within a large action space. To do so, we introduce a text-based multi-armed bandit task, where actions correspond to free-text answers to an open-ended question, and rewards are driven by the distance from some preselected answer in an embedding space (where the embedding is computed exogenously). Given the high dimensionality of the action space, traditional bandit approaches are inapplicable. However, an LLM can generate a small set of candidate actions which can then be used to instantiate an off-the-shelf bandit algorithm. We experiment with several prompting strategies, and find that they all lead to relatively good exploration. Finally, we repeat our experiments on a larger-scale bandit task based on paper titles and abstracts from arXiv (where the goal is to find a suitable title for a given abstract), with similar findings.
2 Related Work and Background
Our results belong to a small, but quickly growing line of work on using pre-trained LLMs for in-context reinforcement learning (RL). Coda-Forno et al. [13], Krishnamurthy et al. [23], Nie et al. [31], Monea et al. [30], Xia et al. [43], Park et al. [33], Wu et al. [42] evaluate the ability of LLMs to solve various multi-armed bandit and contextual bandit tasks, and find that the current generation of LLMs largely fail to solve these tasks in-context. Indeed, positive findings are restricted to very simple tasks and/or require substantial mitigations (which in turn do not readily extend beyond simple settings). Xia et al. [43] use LLMs to solve dueling bandit tasks, and Park et al. [33] also evaluate the ability of LLMs to learn in games. While our paper is primarily concerned with whether LLMs succeed as algorithms, several others [e.g.,  36, 17, 14] use in-context bandits (and many other tasks) to study whether LLMs exhibit human-like behavior/biases in decision-making.
A broader literature on in-context learning [8] aims to solve various tasks by providing all relevant information in the LLM prompt. The work on exemplar selection (selecting examples and other information to present in-context) [e.g.,  19, 48, 44, 40] is relevant to our exploitation experiments.
A growing line of work aims to use LLMs as a part of a larger decision-making agent [e.g.,  28, 50, 49]. Our exploration experiments take inspiration from the work on using LLMs as “action priors” inside of a larger RL algorithm [46, 10, 47, 16]. Much of this work falls under the proposer-verifier framework of Snell et al. [39], where an LLM proposes several possible sequences from which a verifier selects suitable candidates. In comparison, our goal is a more systematic evaluation of LLMs’ abilities to explore large action spaces, in isolation from other components of the decision-making task.
Finally, a parallel line of work trains transformers from scratch to solve various RL tasks [e.g.,  24, 26, 34, 45, 27].
Multi-armed bandits. We consider tasks based on multi-armed bandits (MAB) and contextual bandits (CB), well-studied special cases of RL that abstract the explore-exploit tradeoff, see books [38, 25] for background. In MAB, there are $T$ rounds and $K$ arms. In each round $t\in[T]$, the learner chooses an action (arm) $a_{t}\in[K]$ and observes a noisy reward $r_{t}$ drawn from some sub-Gaussian reward distribution for this arm. The reward distribution, and particularly its mean $\mu(a_{t})$, are unknown to the algorithm. In CB, the learner additionally observes a context $z_{t}$ before each round $t$, and the expected reward $\mu(z_{t},a_{t})$ depends on both the context and the arm. The learner’s goal is to balance exploration and exploitation to maximize cumulative reward.
An “exploitation oracle" (which optimizes for the current round given the history) naturally plugs into standard bandit algorithms such as Epsilon-Greedy, Explore-then-Commit, and Follow-The-Perturbed-Leader. Typical implementations in CB involve model-based (e.g.,  linear) regression or cost-sensitive classification [38, Ch.8]. Designing CB exploitation oracles for their own sake, a.k.a. offline policy optimization, is well-studied [starting from, e.g.,  6, 15].
An “exploration oracle" can suggest a set of candidate actions to discretize a large action space and initialize an off-the-shelf bandit algorithm. Ideally, the oracle would "zoom in" on a lower-dimensional subspace with better-performing actions. Our intuition for exploration oracles is grounded in the study of bandits with Lipschitz-like structure [38, Ch.4]. In particular, the size vs. quality tradeoff in the discretization and the importance of “zooming in" are key issues in this literature [e.g.,  22, 20, 21, 9].
3 LLMs as Exploitation Oracles
In this section, we evaluate the ability of LLMs to exploit in decision-making tasks with statistical uncertainty on the outcomes. We present LLMs with in-context exploit tasks inspired by multi-armed bandits (MAB) and contextual bandits (CB). In a CB exploit task, an LLM is given a history consisting of context-arm-reward tuples, and is instructed to take the best arm given the current history and the current context. A MAB exploit task is the same, but without contexts. These tasks are generated from some parameterized distributions called exploit puzzles. Our visualizations plot the fraction of correct answers against the task difficulty.
3.1 MAB Exploit Puzzles
Figure 1: MAB exploit puzzle for Gpt-4 (left), Gpt-4 with CoT (middle), and Gpt-3.5 with CoT (right), all with "buttons" prompt. The following conventions apply to all figures in this section. Each line corresponds to a particular value of $T$ and plots $\mathrm{FracCorrect}(\epsilon,T)$ against empirical gap $\epsilon$ on the X-axis. The shaded band around the line represents a $95\%$ confidence interval. The dashed line is the number of tasks (“runs") with empirical gap at most $\epsilon$; the corresponding Y-scale is on the right.
[…]
Once an exploit task is generated, we measure its difficulty via empirical gap: the difference between the largest and second-largest average reward $\bar{r}_{t}(a):=\frac{1}{T}\sum_{t\in[T]}r_{t}(a)$ among all arms $a$. Intuitively, the smaller its empirical gap, the more difficult is the puzzle. Let $S(\epsilon,T)$ be the set of all tasks with empirical gap at most $\epsilon$ and history size $T$.
We measure an LLM’s performance over a given set $S$ of tasks as the fraction of tasks for which the LLM returns a “correct answer": an arm with the largest empirical reward; denoted by $\mathrm{FracCorrect}(S)$. We are interested in how $\mathrm{FracCorrect}$ varies depending on the difficulty level. Hence, we plot $\mathrm{FracCorrect}(\epsilon,T):=\mathrm{FracCorrect}\left(\,S(\epsilon,T)\,\right)$ against empirical gap $\epsilon$.
We find that Gpt-4 and Gpt-3.5 do not perform well on these MAB exploit puzzles, see Figure 1. Performance tends to degrade (1) as the history length $T$ increases and (2) as the empirical gap decreases. While Gpt-4 generally performs much better than Gpt-3.5, we found that prompting the LLM to use chain-of-thought (CoT) reasoning provided a slight boost for Gpt-3.5, while hurting the performance of Gpt-4. We found that performance was similar across all of our prompt designs (see Appendix A for more plots).
3.2 CB Exploit Puzzles and Mitigations
Figure 2: Gpt-4 succeeds on a small CB exploit puzzle (left) and fails on a slightly larger one (right).
Figure 3: CB exploit puzzle with $d=K=2$ and $T=4000$: mitigations help substantially. Gpt-4 without CoT (left) and Gpt-4 with CoT (right). Note that providing the full history with this $T$ vastly exceeds the context window for Gpt-4, Gpt-4o, and Gpt-3.5.
While the history in $K$-armed bandits can be summarized using $2K$ numbers (the average reward for each arm and number of times it has been played), such succinct summary statistics may not be readily available (or even exist) in more complicated decision-making tasks such as CB. [^2] This consideration also motivates MAB exploit puzzles with raw (non-summarized) history, as a simpler special case of the general scenario when succinct summarization is unavailable.
We focus on linear CB, a well-studied CB model [starting from 29, 12, 1] in which the expected reward of each arm is linear in the context. Specifically, given context $z\in\mathbb{R}^{d}$ and arm $a$, the expected reward is $\mu(z,a)=\langle z,\theta^{*}_{a}\rangle$, for some fixed (but unknown) parameters $\theta^{*}_{a}\in\mathbb{R}^{d}$.
We consider a CB exploit puzzle parameterized by the number of arms $K$, context dimension $d$, and history size $T$. The tasks are constructed as follows. We sample parameters $\theta_{a}\in[-1,1]^{d}$ and $\gamma_{a}\in[-0.25,0.25]$ independently and u.a.r. for each arm $a$. Given context $z\in\mathbb{R}^{d}$, expected reward for arm $a$ is $\mu(z,a)=\langle z,\theta_{a}\rangle+\gamma_{a}$. We generate a history of $T$ rounds $t\in[T]$. Contexts $z_{t}$ are sampled independently and u.a.r. from $[-1,1]^{d}$. For simplicity, the history contains rewards of all arms $a$ in each round $t$, where the reward is generated as $\mu(z_{t},a)$ plus an independent unit-variance Gaussian. Given the history and a new context $z_{T+1}$ (drawn in the same way), the LLM is asked to select the action which appears best. This gives one exploit task. We generate $N$ tasks for the same $K,T,d$.
We present these exploit tasks to the LLM in-context using a modified “buttons” prompt, where contexts corresponds to “numbers on a screen" which affect the payoffs for pressing each button. The prompt does not mention linearity of the CB instance (because such model-based information is typically unavailable in applications).
For a particular exploit task, we now define a “correct answer" as an arm $a$ which maximizes expected reward $\mu(z_{T+1},a)$.  [^3] Note that it is unclear how to define an ”empirically best arm” given a CB history and the current context. As before, $\mathrm{FracCorrect}(S)$ is the fraction of correct answers in a given set $S$ of tasks.
Likewise, the task difficulty is not easily defined in terms of the realized rewards. Instead, we focus on the effective gap: the difference in expected reward between the best and second-best arm given the current context $z=z_{T+1}$. That is, the difference between the largest and second-largest number among $\mu(a,z)$, $a\in[K]$. Intuitively, smaller effective gap corresponds to increased difficulty.
We are interested in how $\mathrm{FracCorrect}$ varies with effective gap. In each plot, we fix the number of arms ($K$) and dimension ($d$), and let $S(\epsilon,T)$ be the set of all tasks with given $K,d,T$ and effective gap at most $\epsilon$. We plot $\mathrm{FracCorrect}(\epsilon,T):=\mathrm{FracCorrect}\left(\,S(\epsilon,T)\,\right)$ against $\epsilon$.
Figure 4: CB exploit puzzle with $d=K=5$ and $T=1000$: mitigations perform badly, but (mostly) much better than the no-mitigation baseline. Gpt-4o without CoT.
[…]
Motivated by these observations, we implement several natural mitigations inspired by the literature on exemplar selection for in-context learning (discussed in Related Work). The mitigations are as follows:
1.   $k$-nearest: Among the observed contexts, consider the distinct $k$ contexts closest to $z_{T+1}$, according to the $\ell_{2}$ metric. Limit the history reported in the prompt to (the rounds with) these $k$ contexts.
2.   $k$-means: Run an off-the-shelf algorithm for $k$-means clustering on contexts $\left\{\,z_{1}\,,\ \ldots\ ,z_{T}\,\right\}$, obtaining $k$ centroid contexts $z^{*}_{i}$ and their respective clusters $Z^{*}_{i}$, $i\in[k]$. For each centroid $z^{*}_{i}$ and each arm $a$, let $\bar{r}(z^{*}_{i},a)$ be the average reward for this arm over all rounds $t$ with contexts $z_{t}\in Z^{*}_{i}$. Report $\left(\,z^{*}_{i},\,a,\,\bar{r}(z^{*}_{i},a)\,\right)$ as a context-arm-reward triple.
3.   $k$-means, $k^{\prime}<k$-nearest: First, run the $k$-means mitigation. Report $\left(\,z^{*},\,a,\,\bar{r}(z^{*},a)\,\right)$ as a context-arm-reward triple, for each arm $a$ and each centeroid context $z^{*}$ among the $k^{\prime}$ centroids closest to $z_{T+1}$ (according to the $\ell_{2}$ metric).
In all three mitigations, we do not explain the “mitigation strategy" in the prompt. That is, we present the reported context-arm-reward tuples as if it were the entire history, not mentioning nearest neighbors, clustering, or averaging.
Figure 3 visualizes the performance of these mitigations on a slightly larger (but still relatively small) puzzle with $K=d=2$ and $T=4000$. We use Gpt-4 with and without CoT. We compare the mitigations against linear regression baseline (which is effectively an "upper bound", as the underlying CB instance is linear). Without CoT prompting, we find that all three mitigations achieve $\mathrm{FracCorrect}$ around $80\%-90\%$, although this dips to around $60\%-85\%$ when using CoT. In addition to (potentially) improving performance, mitigations can also offer a practical way to solve decision-making tasks using LLMs when the history is large; when $T=4,000$, our prompt vastly exceeds the context window of all models we had access to.
Figure 5: Gpt-4o on the text-based CB exploit puzzle: some mitigations help, but are outperformed by linear regression.
However, current LLMs struggle to exploit on even moderately-sized problems, even with these mitigations. In Figure 4, we plot the performance of Gpt-4o with mitigations under various hyperparameters. While $k$-means (left) is almost as bad as random guessing, the $k$-nearest mitigations (center) achieve about $50\%$ $\mathrm{FracCorrect}$, and the $k$-means, $k^{\prime}$-nearest mitigations (right) obtain approximately $60\%$ $\mathrm{FracCorrect}$. Both our $k$-nearest and $k$-means, $k^{\prime}$-nearest mitigations significantly out-perform unmitigated Gpt-4o, but fall significantly short of the linear baseline.
3.3 CB Exploit Puzzles (text-based)
As a robustness check, we repeat our CB experiments on a text-based exploit puzzle. In this puzzle, contexts are items in a room (e.g. animals, objects on a table), and actions have an associated semantic meaning (e.g. eat the food item, leave the room). Rewards are still presented numerically, and are non-linear functions of both the context and action. See Appendix A for full details on our experimental setup.
Figure 5 shows the performance of Gpt-4o (with mitigations) on this puzzle. While the reward function is non-linear (and thus the linear baseline only achieves  $70\%$ $\mathrm{FracCorrect}$), we find that all configurations are still significantly out-performed by the linear baseline.
4 LLMs as Exploration Oracles
Figure 6: Algorithm’s average expected reward $\overline{\mathtt{rew}}(\text{task},K)$ (averaged over rounds and over runs), against $K$, the number of candidates. Each line corresponds to a prompting strategy or the $\mathtt{Random}$ baseline. The shaded regions represent a $95\%$ confidence interval.
Figure 7: Arm histogram for one particular question (“What is the meaning of freedom?”, left) and averaged over all questions (right). We consider $K=10$ suggestions generated by our one-by-one prompt without encouragement. For a given “run" of the candidate selection, the suggestions are ranked by the expected reward, and then the $i$-th expected reward is averaged over all runs, for each $i\in[K]$.
We now turn our attention to the ability of pre-trained LLMs to explore large action spaces. Specifically, we leverage the inductive bias of an LLM to generate a small set of candidate actions from a text-based action space, before running an off-the-shelf MAB algorithm on top of the candidate set.
We consider two types of exploration tasks: answering an open-ended “philosophical" question (Q/A task) and suggesting a title for an arXiv research paper based on its abstract (arXiv task). Particular workloads within each task type are called explore puzzles.
The Q/A task (resp. arXiv task) is constructed as follows. We define the “best arm" as a contrarian answer generated by another LLM (resp. the actual title of the research paper). The expected reward $\mu(a)$ of an arm $a$ is the cosine similarity between this arm and the best arm in the embedding space. [^5] While cosine similarity ranges on $[-1,1]$, it was usually strictly positive in our experiments. In the (very rare) cases where it was negative we defined the expected reward as zero. Here, we generate sentence embeddings using the Sentence-BERT embedding model [35]. [^6] Robustness check: results are largely unchanged under the Universal Sentence Encoder [11], see Appendix B. The realized reward in a given round is an independent Bernoulli sample with the corresponding mean defined above.
Since action spaces are extremely large for these tasks, bandit algorithms which randomly subsample the action space catastrophically fail. If the embedding space, the distance notion therein, and the reward-distance relation were known to the algorithm/agent (assumptions we do not make), one could, in principle, apply “zooming algorithms" from Lipschitz bandits which gradually “zoom in" on better-performing regions of the metric space and may therefore achieve performance that scales advantageously over time [20, 21, 9, 37]. However, these algorithms suffer an exponential dependence on the dimension $d$ of the metric space (in the multiplicative constant in front of the regret rate), making them impractical in high-dimensional spaces such as ours.  [^7] Indeed, our embedding space has $d=384$.
4.1 Explore Puzzle: Open-Ended Questions
We used Gpt-4 to generate a dataset of $10$ open-ended questions with many reasonable answers, along with an intentionally contrarian answer for each question to serve as the ground truth. (E.g.,  “What does it mean to live a fulfilling life?” “Fulfillment comes from embracing discomfort.”) Each question-answer pair yields a task, as defined above.
We evaluate Gpt-4o as an exploration oracle on these tasks. We prompt it to suggest $K\in\{1,2,3,4,5,7,10\}$ candidate answers given a question. To grade the candidate set (the entire set, not just the best answer in this set), we run an off-the-shelf MAB algorithm for some large-but-realistic time horizon $T$ over these $K$ candidates. (We use the UCB1 algorithm [5] and $T=1000$.) We record the expected time-averaged reward $\mathtt{rew}:=\tfrac{1}{T}\,\sum_{t\in[T]}\mu(a_{t})$, where $a_{t}$ is the arm selected in round $t$. We repeat this process (both selecting the candidates and running the bandit algorithm) 10 times for a given task and $K$, and record the average $\mathtt{rew}$ over these runs, denoted by $\overline{\mathtt{rew}}=\overline{\mathtt{rew}}(\text{task},K)$.
[…]
To compare against not using an LLM as an exploration oracle, we also consider a baseline ($\mathtt{Random}$) in which the candidate set consists of $K$ points selected independently and uniformly at random in the embedding space.
We visualize our findings in Figure 6. We plot $\overline{\mathtt{rew}}(\text{task},K)$ against $K$, for one particular task (left) and averaged across all tasks (right). [^8] See Appendix B for similar plots for the $9$ other tasks. Each line corresponds to a particular prompting strategy or the $\mathtt{Random}$ baseline.
We find that all four prompting strategies perform similarly, with average reward between $0.5$ and $0.6$, and typically peak in performance around $K=3$ or $K=4$ suggestions. In contrast, the $\mathtt{Random}$ baseline catastrophically fails, with its average reward never exceeding $0.1$. We conclude that the LLM does succeed as an exploration oracle.
We discuss some additional findings below. First, we observe that the LLM-generated suggestions pass the “eye test”, in the sense that we get reasonable, yet spiritually and semantically different answers for a given question. For example, given the question “What is the role of technology in society”, the first $K=5$ suggestions generated by our one-by-one prompt are as follows:
  1. 1.
    Facilitates communication, innovation, and efficiency.
  2. 2.
    Transforms daily life and shapes culture.
  3. 3.
    Drives connectivity and enhances productivity.
  4. 4.
    Facilitates control and surveillance.
  5. 5.
    Disrupts traditional relationships and norms.
Second, we verify that the candidate suggestions are substantially different from one another. To this end, Figure 7 visualizes the spread of expected rewards within the candidate set. We consider $K=10$ suggestions generated by our one-by-one prompt without encouragement. For a given “run" of the candidate selection, the suggestions are ranked by the expected reward, and then the $i$-th expected reward is averaged over all runs, for each $i\in[K]$. We consider one question (left) and the average over all questions (right).
A more detailed comparison between the prompting strategies is perhaps not fruitful, at least within this type of experiments, because sentence embeddings are known to be somewhat imprecise at small scales (see Section B.4).
Figure 8: Algorithm’s average expected reward $\overline{\mathtt{rew}}(\text{category},K)$ (averaged over rounds and tasks), against $K$, the number of candidates. Each line corresponds to a prompting strategy or the Category-Only baseline. The shaded regions represent a $95\%$ confidence interval. A single arXiv category (“General Relativity and Quantum Cosmology", left), averages over 6 categories (right).
4.2 Explore Puzzle: arXiv Abstracts and Titles
We run similar experiments on a larger-scale dataset of paper titles and abstracts from arXiv.org. Using the arXiv API [4], we collect $10$ (paper abstract, paper title) pairs from each of the $41$ different arXiv categories across various academic disciplines. To minimize the likelihood that these papers appear in Gpt-4o’s training corpus, we only use papers uploaded to arXiv after June 2024. Each abstract-title pair yields a task, as discussed earlier.
We evaluate Gpt-4o as an exploration oracle for these tasks much like in in Section 4.1. Given an abstract, we prompt Gpt-4o to generate $K$ alternative titles, which are then used to instantiate a bandit algorithm. We use the same algorithm (UCB1) and time horizon $T=1000$. We record the expected time-averaged reward $\mathtt{rew}:=\tfrac{1}{T}\,\sum_{t\in[T]}\mu(a_{t})$ and compute the average over tasks within the same arXiv category, called $\overline{\mathtt{rew}}=\overline{\mathtt{rew}}(\text{category},K)$. We try “all-at-once" and “one-by-one" prompting strategies. [^9] We do not use “encouragement” on these tasks, since it does not appear to help much (if at all) in Section 4.1.
We visualize our findings in Figure 8, using the same conventions as Figure 6 and focusing on six arXiv categories. [^10] General relativity and quantum cosmology; computer vision and pattern recognition; statistics theory; biomolecules; signal processing; and general economics. For the other arXiv categories, we only collected the data for $K\in\{1,2,5\}$. To assess LLM’s ability to specialize to a task, we consider a stronger baseline, where the candidate arms are generated by Gpt-4o given only the category, not the abstract.
Category $K=1$ $K=2$ $K=5$
gr-qc 0.63 0.64 0.65
hep-ex 0.78 0.76 0.81
hep-lat 0.72 0.72 0.74
hep-ph 0.7 0.72 0.73
hep-th 0.64 0.69 0.68
math-ph 0.65 0.71 0.71
Table 1: Results for other arXiv categories: per-category rewards $\overline{\mathtt{rew}}(\text{category},K)$ for one-by-one prompting and $K\in\{1,2,5\}$.
Results for the other arXiv categories are provided via a per-category table with $\overline{\mathtt{rew}}(\text{category},K)$ for $K\in\{1,2,5\}$ and both prompts (Appendix B). In Table 1, we show them for (other) six arXiv categories and one-by-one prompting.
Similar to Section 4.1, we conclude that (1) both prompting strategies significantly outperform the baseline, and (2) their performance tends to improve modestly as $K$ increases.
5 Conclusions
We evaluate the potential of Gpt-4, Gpt-4o, and Gpt-3.5 to explore and exploit in decision-making tasks. We find that LLMs are useful as exploration oracles in large, semantically meaningful action spaces, where they effectively propose high-quality candidate actions. However, LLMs cannot yet robustly replace (let alone surpass) traditional algorithmic methods for exploitation, particularly in larger or more complex settings. While we suggest several helpful mitigations, they consistently underperform relative to a simple algorithmic baseline such as linear regression.
We highlight two directions for future work. First, LLMs trained to use tools like a calculator may be better at exploitation. However, it is unclear how much this would help in more complex scenarios, e.g.,  CB tasks with text-based contexts and actions, and which mitigations and/or prompting techniques would be needed. Second, while "zooming" bandit algorithms do not work for rich text-based action spaces (as discussed in Section 4), LLM-based exploration oracles may potentially help. The hope is to “zoom in" entirely in the space of “potentially relevant" actions (determined by the LLM), rather than in the space of all actions.
Appendix A Appendix for Section 3: LLMs as Exploitation Oracles
A.1 Prompts
In this section we give example prompts for each of our experimental setups.
“Buttons” prompt for the MAB puzzle:
[SYSTEM] You are in a room with 5 buttons labeled blue, green, red, yellow, purple. Each button is associated with a Bernoulli distribution with a fixed but unknown mean; the means for the buttons could be different. For each button, when you press it, you will get a reward that is sampled from the button’s associated distribution. Then you must pick the button with the highest empirical average, which must be exactly one of blue, green, red, yellow, or purple. You must provide your final answer immediately within the tags <Answer>COLOR <Answer>where COLOR is one of blue, green, red, yellow, or purple and with no text explanation.
[USER] The past rewards for each button are:
round $1$: blue button had reward $1$, green button had reward $1$, red button had reward 00, yellow button had reward $1$, purple button had reward 00
round $2$: blue button had reward 00, green button had reward $1$, red button had reward $1$, yellow button had reward $1$, purple button had reward 00
Which button do you choose? Remember, YOU MUST provide your final answer within the tags <Answer>COLOR <Answer>where COLOR is one of blue, green, red, yellow, or purple and with no text explanation.
“Adverts” prompt for the MAB puzzle:
[SYSTEM] You are recommendation engine that chooses advertisements to display to users when they visit your webpage. There are 5 advertisements you can choose from, named A, B, C, D, E. When a user visits the webpage you can choose an advertisement to display and you will observe whether the user would have clicked each of the ads. You model this by assuming that each advertisement has a certain click rate and users click on advertisements with their corresponding rates. I will show you the past clicks for each advertisement. Then you must pick the advertisement with the highest empirical click rate, which must be exactly one of A, B, C, D, or E. You must provide your final answer immediately and with no text explanation. within the tags <Answer>ADVERTISEMENT <Answer>where ADVERTISEMENT is one of A, B, C, D, or E.
[USER] The past clicks for each advertisement are:
round $1$: advertisement A was clicked, advertisement B was clicked, advertisement C was not clicked, advertisement D was clicked, advertisement E was clicked
round $1$: advertisement A was not clicked, advertisement B was clicked, advertisement C was clicked, advertisement D was clicked, advertisement E was not clicked
Which advertisement do you choose? Remember, YOU MUST provide your final answer within the tags <Answer>ADVERTISEMENT <Answer>where ADVERTISEMENT is one of A, B, C, D, or E and with no text explanation.
“Buttons” prompt for the numerical CB puzzle:
[SYSTEM] You are in a room with a television and 2 buttons labeled blue, green. Each button is associated with a Bernoulli distribution with an unknown mean; the means for the buttons could be different from each other and may depend on the list of numbers shown on the screen (i.e. the context). For each button, when you press it, you will get a reward that is sampled from the button’s associated distribution, conditioned on the numbers shown on the television screen. I will show you the past numbers shown on the screen and the corresponding rewards for each button. A new list of numbers will then appear on the screen and you must pick the next button in order to maximize your reward in this round only, which must be exactly one of blue or green. You must provide your final answer immediately within the tags <Answer> COLOR </Answer> where COLOR is one of blue or green and with no text explanation.
[USER] The past contexts and rewards for each button are:
In round $1$, the context was $[0.3,0.7]$. The blue button had reward $1$, the green button had reward $1$
In round $2$, the context was $[0.4,0.6]$. The blue button had reward 00, the green button had reward $1$
Which button do you choose? Remember, YOU MUST provide your final answer within the tags <Answer>COLOR <Answer>where COLOR is one of blue or green and with no text explanation.
Prompt for the text-based CB puzzle:
[SYSTEM] You are in a room with a table and a button. There may also be other objects in the room, which I will tell you about. You must then take one of the following actions: "pet animal", "leave room", "use tool", "eat food", "press button", after which you will receive some reward. The reward you receive is a random function of both the action you take and the information you receive about the objects in the room and time of day. Your goal is to maximize the expected reward you receive. I will show you the past history of play over 2 rounds. For each round, I will show you the state of the room and the corresponding rewards for each action. I will then tell you the current state of the room, and you must pick the next action in order to maximize your reward in this round only, which must be exactly one of "pet animal", "leave room", "use tool", "eat food", or "press button". Look for patterns in the data and try to estimate the reward of each action, given the information at your disposal. You must provide your final answer immediately within the tags <Answer>ACTION <Answer>where ACTION is one of "pet animal", "leave room", "use tool", "eat food", or "press button" and with no text explanation.
[USER] The past observations and outcomes for each action are:
Round $1$ had context time of day: morning, animal: bear, table item: chest, tool: key, food: apple, button color: red. "pet animal" had reward 00, "leave room" had reward $1$, "use tool" had reward $1$, "eat food" had reward 00, press button had reward 00
Round $2$ had context time of day: afternoon, animal: cat, table item: card, tool: hammer, food: cake, button color: orange. "pet animal" had reward $1$, "leave room" had reward 00, "use tool" had reward 00, "eat food" had reward $1$, press button had reward 00
The current state of the room is time of day: evening, animal: bear, table item: envelope, tool: key, food: nut, button color: red.
Which action do you choose? Remember, you must provide your final answer immediately within the tags <Answer>ACTION <Answer>where ACTION is one of "pet animal", "leave room", "use tool", "eat food", or "press button" and with no text explanation.
A.2 Additional MAB Figures
See Figure 9Figure 10Figure 11 for additional results in our MAB exploit puzzle.
Figure 9: Cumulative fraction correct for Gpt-4 in the MAB adverts puzzle.
Figure 10: Cumulative fraction correct for Gpt-4 with chain-of-thought reasoning in the MAB adverts puzzle.
Figure 11: Cumulative fraction correct for Gpt-3.5 in the MAB buttons puzzle.
A.3 Additional Details for Text-Based CB Puzzles
Each context contains a time of day (belonging to {morning, afternoon, evening, night}), an animal ({bear, dog, cat, None}), a tool ({key, letter opener, hammer, None}), a food item ({cake, apple, nut, None}), and a button with a particular color ({red, orange, yellow, green}). The actions in each round are “pet animal”, “leave room”, “use tool”, “eat food”, and “press button”.
We experimented with two reward functions: an “easy” reward function, where the expected rewards for each action are as follows:
  • The expected reward of petting the animal is 0.01 if the animal is a bear, 0.7 if the animal is a dog, and 0.4 if the animal is a cat. Otherwise, the expected reward if 0.5.
  • The expected reward for leaving the room is always 0.5.
  • The expected reward for using the tool is 0.75 if it is a key, 0.6 if it is a letter opener, 0.45 if it is a hammer, and 0.2 otherwise.
  • The expected reward for eating food is 0.8 if it is cake, 0.6 if it is an apple, 0.2 if it is a nut, and 0.3 otherwise.
  • The expected reward for pressing the button is 0.89 if it is green, 0.62 if it is yellow, 0.39 if it is orange, and 0.27 if it is red.
Our results under this reward function are summarized in Figure 12. We used hamming distance to implement our mitigations. Note that in higher-dimensional settings, distance in an embedding space may be used.
Figure 12: Performance of Gpt-4o with mitigations on the words CB puzzle with “easy” rewards.
The reward function we use in the main body is more complicated, and is detailed below:
  • The expected reward for petting the animal is 0.01 if it is a bear, 0.7 if it is a dog, 0.3 if it is a cat and the time of day is morning or afternoon, 0.7 if it is a cat and the time of day is evening or night, and otherwise 0.5.
  • The expected reward for leaving the room is always 0.5
  • If the animal is a bear, the expected reward for using the tool is 0.1. Otherwise, if the tool is a key and the table item is a chest, the expected reward is 0.9. Otherwise, it is 0.4
  • If the animal is a bear, the expected reward for eating food is always 0.5. Otherwise, the expected reward is 0.8 for cake, 0.6 for an apple, 0.2 for a nut, and 0.5 otherwise.
  • If the animal is a bear, the expected reward for pressing the button is 0.1. Otherwise if the button is green and the time of day is morning, or the button is yellow and the time of day is afternoon, or the button is orange and the time of day is evening, or the button is red and the time of day is night, then the expected reward is 0.9. In all other cases, the expected reward is 0.25.
Appendix B Appendix for Section 4: LLMs as Exploration Oracles
B.1 Prompts
In this section we give example prompts for each of our experimental setups.
“All-at-once” prompt with encouragement for open-ended question puzzle:
[SYSTEM] I will give you an open-ended question. Come up with 5 different candidate answers. Reply only with the 5 candidate answers, and put each candidate answer on a separate line. Each answer should only be a few words, skipping any introductory phrasing and going straight to the essence. Try to come up with answers that are very different in spirit from one another.
[USER] Here is the question: “What is the purpose of art?”
“One-by-one” prompt without encouragement for arXiv puzzle:
[SYSTEM] I will give you an abstract and some candidate titles for a paper. Come up with a new candidate title that is relevant to the abstract, but different from the other candidate titles. Reply only with the candidate title.
[USER] Here is the abstract: {abstract goes here}
Here are the other candidate titles: {previous suggestions go here}
B.2 Datasets
Our open-ended question dataset consists of the following $10$ questions and the corresponding “ground-truth” answers.
Questions:
  1. 0.
    What is the meaning of freedom?
  2. 1.
    How should we define success?
  3. 2.
    What is the role of technology in society?
  4. 3.
    What is the nature of reality?
  5. 4.
    What is the purpose of art?
  6. 5.
    What does it mean to live a fulfilling life?
  7. 6.
    How do cultural differences shape our understanding of morality?
  8. 7.
    What is the relationship between happiness and wealth?
  9. 8.
    How can we balance individuality and community in modern society?
  10. 9.
    What is the role of education in personal and societal growth?
Answers:
  1. 0.
    Freedom is an illusion shaped by societal norms and external influences.
  2. 1.
    Success should be defined as contributing to the greater good rather than personal achievement.
  3. 2.
    Technology disrupts the natural balance of society and often creates more problems than it solves.
  4. 3.
    Reality is subjective, varying entirely based on individual perception and experience.
  5. 4.
    The purpose of art is to challenge conventions and disrupt established ideas.
  6. 5.
    Fulfillment comes from embracing discomfort.
  7. 6.
    Cultural differences create moral superiority.
  8. 7.
    Wealth detracts from true happiness.
  9. 8.
    Individuality thrives when shaped by community.
  10. 9.
    Education’s purpose is to challenge authority.
Here is the list of paper titles we used in our arXiv dataset, along with their corresponding categories:
gr-qc
  1. 1.
    There is more to the de Sitter horizon than just the area
  2. 2.
    Mitigating cosmic variance in the Hellings-Downs curve: a Cosmic Microwave Background analogy
  3. 3.
    Calabi-Yau Feynman integrals in gravity: $\varepsilon$-factorized form for apparent singularities
  4. 4.
    QG from SymQRG: AdS3/CFT2 Correspondence as Topological Symmetry-Preserving Quantum RG Flow
  5. 5.
    Black hole solutions in theories of supergravity
  6. 6.
    Horndeski in motion
  7. 7.
    Wormholes from beyond
  8. 8.
    Regularizing the Pulsar Timing Array likelihood: A path towards Fourier Space
  9. 9.
    Solutions to the mode equation for a quantized massless scalar field outside a black hole that forms from the collapse of a null shell: Late-time behaviors and computation of the stress-energy tensor
  10. 10.
    Gravitational waves from regular black holes in extreme mass-ratio inspirals
hep-ex
  1. 1.
    Observation of the $K^{+}\rightarrow\pi^{+}\nu\bar{\nu}$ decay and measurement of its branching ratio
  2. 2.
    Test of lepton flavour universality in $W$-boson decays into electrons and $\tau$-leptons using $pp$ collisions at $\sqrt{s}=13$ TeV with the ATLAS detector
  3. 3.
    Searching for neutrino self-interactions at future muon colliders
  4. 4.
    Quantum Decoherence at ESSnuSB Experiment
  5. 5.
    Test of lepton flavour universality with $B^{+}\to K^{+}\pi^{+}\pi^{-}\ell^{+}\ell^{-}$ decays
  6. 6.
    Cross-section measurements for the production of a $W$-boson in association with high-transverse-momentum jets in $pp$ collisions at $\sqrt{s}$= 13 TeV with the ATLAS detector
  7. 7.
    Charmful two-body $\Omega_{b}$ decays in the light-front quark model
  8. 8.
    Observation of a spectral hardening in cosmic ray boron spectrum with the DAMPE space mission
  9. 9.
    New BaBar studies of high-order radiation and the new landscape of data-driven HVP predictions of the muon g-2
  10. 10.
    Toponium: the smallest bound state and simplest hadron in quantum mechanics
hep-lat
  1. 1.
    Quantum sampling on a quantum annealer for large volumes in the strong coupling limit for gauge group U(3)
  2. 2.
    Phase diagram of Rydberg atoms in a two-leg rectangular ladder
  3. 3.
    Graph Attention Hamiltonian Neural Networks: A Lattice System Analysis Model Based on Structural Learning
  4. 4.
    What do we know about the confinement mechanism?
  5. 5.
    Designing weight regularizations based on Lefschetz thimbles to stabilize complex Langevin
  6. 6.
    Likelihood of a zero in the proton elastic electric form factor
  7. 7.
    Real-Time Simulation of Asymmetry Generation in Fermion-Bubble Collisions
  8. 8.
    Investigating SU(3) with Nf=8 fundamental fermions at strong renormalized coupling
  9. 9.
    The determination of potential scales in 2+1 flavor QCD
  10. 10.
    Towards the phase diagram of fermions coupled with $SO(3)$ quantum links in $(2+1)$-D
hep-ph
  1. 1.
    Predictions for dimuon production in high-energy neutrino-proton collisions using the color dipole model
  2. 2.
    Extrapolating Jet Radiation with Autoregressive Transformers
  3. 3.
    Accurate Surrogate Amplitudes with Calibrated Uncertainties
  4. 4.
    Calabi-Yau Feynman integrals in gravity: $\varepsilon$-factorized form for apparent singularities
  5. 5.
    The causal structure of the quark propagator
  6. 6.
    Fuzzy Axions and Associated Relics
  7. 7.
    Non-Radial Oscillation Modes in Hybrid Stars with Hyperons and Delta Baryons: Full General Relativity Formalism vs. Cowling Approximation
  8. 8.
    Evidence for the Sombrero Galaxy as an Accelerator of the Highest-Energy Cosmic Rays
  9. 9.
    The cosmic history of Primordial Black Hole accretion and its uncertainties
  10. 10.
    Searching for neutrino self-interactions at future muon colliders
hep-th
  1. 1.
    There is more to the de Sitter horizon than just the area
  2. 2.
    Calabi-Yau Feynman integrals in gravity: $\varepsilon$-factorized form for apparent singularities
  3. 3.
    QG from SymQRG: AdS3/CFT2 Correspondence as Topological Symmetry-Preserving Quantum RG Flow
  4. 4.
    Geometrically constrained localized configurations engendering non-topological profile
  5. 5.
    The causal structure of the quark propagator
  6. 6.
    Entanglement Hamiltonian and orthogonal polynomials
  7. 7.
    Black hole solutions in theories of supergravity
  8. 8.
    Fuzzy Axions and Associated Relics
  9. 9.
    Celestial Mellin Amplitudes
  10. 10.
    Evidence for the Sombrero Galaxy as an Accelerator of the Highest-Energy Cosmic Rays
math-ph
  1. 1.
    QG from SymQRG: AdS3/CFT2 Correspondence as Topological Symmetry-Preserving Quantum RG Flow
  2. 2.
    Entanglement Hamiltonian and orthogonal polynomials
  3. 3.
    Fermi’s golden rule in tunneling models with quantum waveguides perturbed by Kato class measures
  4. 4.
    Semiclassical measure of the propagation between two topological insulators
  5. 5.
    On the Protection Against Noise for Measurement-Based Quantum Computation
  6. 6.
    Calculating Spectra by Sequential High-Pass Filtering
  7. 7.
    Validity of the stochastic Landau approximation for super-pattern forming systems with a spatial 1:3 resonance
  8. 8.
    Multi-component Hamiltonian difference operators
  9. 9.
    Emptiness Instanton in Quantum Polytropic Gas
  10. 10.
    Unitary $n$-correlations with restricted support in random matrix theory
nucl-ex
  1. 1.
    The evidence of $N=16$ shell closure and $\beta$-delayed neutron emission from $\wedge^{25}$
  2. 2.
    Isotopic Transparency in Central Xe+Sn Collisions at 100 MeV/nucleon
  3. 3.
    Detecting the Coupling of Axion Dark Matter to Neutron Spins at Spallation Sources via Rabi Oscillation
  4. 4.
    Likelihood of a zero in the proton elastic electric form factor
  5. 5.
    Nuclear structure and direct reaction studies in particle-$\gamma$ coincidence experiments at the FSU John D. Fox Superconducting Linear Accelerator Laboratory
  6. 6.
    Bottomonium-like states in proton collisions: Fragmentation and resummation
  7. 7.
    Towards a foundation model for heavy-ion collision experiments through point cloud diffusion
  8. 8.
    Study of the energy spectrum of alpha particles in an experiment on irradiation of a boron target with a proton beam at the Prometheus accelerator
  9. 9.
    Staking out the Proton Drip-Line of Thulium at the N=82 Shell Closure
  10. 10.
    Measurements of global and local spin polarization of $\wedge$ and $\bar{\wedge}$ in Au+Au collisions from the RHIC Beam Energy Scan
nucl-th
  1. 1.
    The causal structure of the quark propagator
  2. 2.
    Non-Radial Oscillation Modes in Hybrid Stars with Hyperons and Delta Baryons: Full General Relativity Formalism vs. Cowling Approximation
  3. 3.
    Isotopic Transparency in Central Xe+Sn Collisions at 100 MeV/nucleon
  4. 4.
    Quantum-Corrected Holographic Wilson Loop Correlators and Confinement
  5. 5.
    Dynamics of Hot QCD Matter 2024 – Bulk Properties
  6. 6.
    Spurious Isospin Breaking in the In-medium Similarity Renormalization Group
  7. 7.
    Likelihood of a zero in the proton elastic electric form factor
  8. 8.
    Born-Oppenheimer Renormalization group for High Energy Scattering: the Modified BFKL, or where did it all go?
  9. 9.
    Nuclear structure and direct reaction studies in particle-$\gamma$ coincidence experiments at the FSU John D. Fox Superconducting Linear Accelerator Laboratory
  10. 10.
    Bottomonium-like states in proton collisions: Fragmentation and resummation
quant-ph
  1. 1.
    Noisy initial-state qubit-channel metrology with additional undesirable noisy evolution
  2. 2.
    The State Preparation of Multivariate Normal Distributions using Tree Tensor Network
  3. 3.
    QG from SymQRG: AdS3/CFT2 Correspondence as Topological Symmetry-Preserving Quantum RG Flow
  4. 4.
    Entanglement Hamiltonian and orthogonal polynomials
  5. 5.
    Learning interactions between Rydberg atoms
  6. 6.
    The isoholonomic inequality and tight implementations of holonomic quantum gates
  7. 7.
    Fermi’s golden rule in tunneling models with quantum waveguides perturbed by Kato class measures
  8. 8.
    Temporal evolution of a forced optomechanical system with linear and quadratic field – mechanical oscillator couplings
  9. 9.
    Control of a Josephson Digital Phase Detector via an SFQ-based Flux Bias Driver
  10. 10.
    Commentary on the decomposition of universal multiport interferometers: how it works in practice
cs.AI
  1. 1.
    MaxInfoRL: Boosting exploration in reinforcement learning through information gain maximization
  2. 2.
    SepLLM: Accelerate Large Language Models by Compressing One Segment into One Separator
  3. 3.
    Stabilizing Reinforcement Learning in Differentiable Multiphysics Simulation
  4. 4.
    Revelations: A Decidable Class of POMDPs with Omega-Regular Objectives
  5. 5.
    Artificial Intelligence in Traffic Systems
  6. 6.
    The Impact of AI Assistance on Radiology Reporting: A Pilot Study Using Simulated AI Draft Reports
  7. 7.
    Can LLM Prompting Serve as a Proxy for Static Analysis in Vulnerability Detection
  8. 8.
    FSFM: A Generalizable Face Security Foundation Model via Self-Supervised Facial Representation Learning
  9. 9.
    Learning to Navigate in Mazes with Novel Layouts using Abstract Top-down Maps
  10. 10.
    SpeechPrune: Context-aware Token Pruning for Speech Information Retrieval
cs.CL
  1. 1.
    SepLLM: Accelerate Large Language Models by Compressing One Segment into One Separator
  2. 2.
    Making FETCH! Happen: Finding Emergent Dog Whistles Through Common Habitats
  3. 3.
    Semi-automated analysis of audio-recorded lessons: The case of teachers’ engaging messages
  4. 4.
    Virtual Agent-Based Communication Skills Training to Facilitate Health Persuasion Among Peers
  5. 5.
    How Private are Language Models in Abstractive Summarization?
  6. 6.
    Can LLM Prompting Serve as a Proxy for Static Analysis in Vulnerability Detection
  7. 7.
    SpeechPrune: Context-aware Token Pruning for Speech Information Retrieval
  8. 8.
    The Open Source Advantage in Large Language Models (LLMs)
  9. 9.
    LLM-RG4: Flexible and Factual Radiology Report Generation across Diverse Input Contexts
  10. 10.
    ExecRepoBench: Multi-level Executable Code Completion Evaluation
cs.CV
  1. 1.
    PanSplat: 4K Panorama Synthesis with Feed-Forward Gaussian Splatting
  2. 2.
    Causal Diffusion Transformers for Generative Modeling
  3. 3.
    CAP4D: Creating Animatable 4D Portrait Avatars with Morphable Multi-View Diffusion Models
  4. 4.
    Wonderland: Navigating 3D Scenes from a Single Image
  5. 5.
    Stabilizing Reinforcement Learning in Differentiable Multiphysics Simulation
  6. 6.
    Instruction-based Image Manipulation by Watching How Things Move
  7. 7.
    IDArb: Intrinsic Decomposition for Arbitrary Number of Input Views and Illuminations
  8. 8.
    UniLoc: Towards Universal Place Recognition Using Any Single Modality
  9. 9.
    CPath-Omni: A Unified Multimodal Foundation Model for Patch and Whole Slide Image Analysis in Computational Pathology
  10. 10.
    CG-Bench: Clue-grounded Question Answering Benchmark for Long Video Understanding
cs.LG
  1. 1.
    MaxInfoRL: Boosting exploration in reinforcement learning through information gain maximization
  2. 2.
    SepLLM: Accelerate Large Language Models by Compressing One Segment into One Separator
  3. 3.
    No More Tuning: Prioritized Multi-Task Learning with Lagrangian Differential Multiplier Methods
  4. 4.
    Stabilizing Reinforcement Learning in Differentiable Multiphysics Simulation
  5. 5.
    Extrapolating Jet Radiation with Autoregressive Transformers
  6. 6.
    Bilevel Learning with Inexact Stochastic Gradients
  7. 7.
    LLMs for Cold-Start Cutting Plane Separator Configuration
  8. 8.
    LeARN: Learnable and Adaptive Representations for Nonlinear Dynamics in System Identification
  9. 9.
    Thermodynamics-informed graph neural networks for real-time simulation of digital human twins
  10. 10.
    Memory-Reduced Meta-Learning with Guaranteed Convergence
cs.NE
  1. 1.
    Deep-learning-based identification of individual motion characteristics from upper-limb trajectories towards disorder stage evaluation
  2. 2.
    Speeding Up the NSGA-II With a Simple Tie-Breaking Rule
  3. 3.
    Optimal Gradient Checkpointing for Sparse and Recurrent Architectures using Off-Chip Memory
  4. 4.
    Runtime Analysis for Multi-Objective Evolutionary Algorithms in Unbounded Integer Spaces
  5. 5.
    Theoretical Analysis of Quality Diversity Algorithms for a Classical Path Planning Problem
  6. 6.
    Populating cellular metamaterials on the extrema of attainable elasticity through neuroevolution
  7. 7.
    Deployment Pipeline from Rockpool to Xylo for Edge Computing
  8. 8.
    Interlocking-free Selective Rationalization Through Genetic-based Learning
  9. 9.
    EVOS: Efficient Implicit Neural Training via EVOlutionary Selector
  10. 10.
    Brain-inspired Chaotic Graph Backpropagation for Large-scale Combinatorial Optimization
cs.RO
  1. 1.
    MaxInfoRL: Boosting exploration in reinforcement learning through information gain maximization
  2. 2.
    Stabilizing Reinforcement Learning in Differentiable Multiphysics Simulation
  3. 3.
    LeARN: Learnable and Adaptive Representations for Nonlinear Dynamics in System Identification
  4. 4.
    Backstepping Control of Tendon-Driven Continuum Robots in Large Deflections Using the Cosserat Rod Model
  5. 5.
    Learning to Navigate in Mazes with Novel Layouts using Abstract Top-down Maps
  6. 6.
    Emma-X: An Embodied Multimodal Action Model with Grounded Chain of Thought and Look-ahead Spatial Reasoning
  7. 7.
    Lightweight Decentralized Neural Network-Based Strategies for Multi-Robot Patrolling
  8. 8.
    Learning Human-Aware Robot Policies for Adaptive Assistance
  9. 9.
    Hardware-in-the-loop Simulation Testbed for Geomagnetic Navigation
  10. 10.
    Sonar-based Deep Learning in Underwater Robotics: Overview, Robustness and Challenges
cs.IT
  1. 1.
    Codes from $A_{m}$-invariant polynomials
  2. 2.
    BA-BFL: Barycentric Aggregation for Bayesian Federated Learning
  3. 3.
    Capacity of Hierarchical Secure Coded Gradient Aggregation with Straggling Communication Links
  4. 4.
    Wireless Environmental Information Theory: A New Paradigm towards 6G Online and Proactive Environment Intelligence Communication
  5. 5.
    Quantum search in a dictionary based on fingerprinting-hashing
  6. 6.
    Identification Over Binary Noisy Permutation Channels
  7. 7.
    Iterative Detection and Decoding for Clustered Cell-Free Massive MIMO Networks
  8. 8.
    Structured Sampling for Robust Euclidean Distance Geometry
  9. 9.
    Study of Iterative Detection and Decoding for Multiuser Systems and MMSE Refinements with Active or Passive RIS
  10. 10.
    Shannon information and integrated information: message and meaning
cs.CR
  1. 1.
    Can LLM Prompting Serve as a Proxy for Static Analysis in Vulnerability Detection
  2. 2.
    Efficient Layered New Bit-Flipping QC-MDPC Decoder for BIKE Post-Quantum Cryptography
  3. 3.
    But Can You Use It? Design Recommendations for Differentially Private Interactive Systems
  4. 4.
    Efficiently Achieving Secure Model Training and Secure Aggregation to Ensure Bidirectional Privacy-Preservation in Federated Learning
  5. 5.
    On Large Language Models in Mission-Critical IT Governance: Are We Ready Yet?
  6. 6.
    Just a Simple Transformation is Enough for Data Protection in Vertical Federated Learning
  7. 7.
    SeSeMI: Secure Serverless Model Inference on Sensitive Data
  8. 8.
    DB-PAISA: Discovery-Based Privacy-Agile IoT Sensing+Actuation
  9. 9.
    OTA-Key: Over the Air Key Management for Flexible and Reliable IoT Device Provision
  10. 10.
    Android App Feature Extraction: A review of approaches for malware and app similarity detection
cs.DS
  1. 1.
    Approximating the Top Eigenvector in Random Order Streams
  2. 2.
    Witty: An Efficient Solver for Computing Minimum-Size Decision Trees
  3. 3.
    Adaptive Manipulation for Coalitions in Knockout Tournaments
  4. 4.
    Counting Butterflies over Streaming Bipartite Graphs with Duplicate Edges
  5. 5.
    Quantum search in a dictionary based on fingerprinting-hashing
  6. 6.
    Regularized Dikin Walks for Sampling Truncated Logconcave Measures, Mixed Isoperimetry and Beyond Worst-Case Analysis
  7. 7.
    Proportionally Fair Matching via Randomized Rounding
  8. 8.
    Logarithmic Positional Partition Interval Encoding
  9. 9.
    New results for the detection of bicliques
  10. 10.
    Deterministic Even-Cycle Detection in Broadcast CONGEST
cs.HC
  1. 1.
    Virtual Agent-Based Communication Skills Training to Facilitate Health Persuasion Among Peers
  2. 2.
    The Impact of AI Assistance on Radiology Reporting: A Pilot Study Using Simulated AI Draft Reports
  3. 3.
    Combining Large Language Models with Tutoring System Intelligence: A Case Study in Caregiver Homework Support
  4. 4.
    But Can You Use It? Design Recommendations for Differentially Private Interactive Systems
  5. 5.
    LLMs Can Simulate Standardized Patients via Agent Coevolution
  6. 6.
    LLM-DaaS: LLM-driven Drone-as-a-Service Operations from Text User Requests
  7. 7.
    Private Yet Social: How LLM Chatbots Support and Challenge Eating Disorder Recovery
  8. 8.
    Task-Based Role-Playing VR Game for Supporting Intellectual Disability Therapies
  9. 9.
    Privacy-Preserving Brain-Computer Interfaces: A Systematic Review
  10. 10.
    Accurate, Robust and Privacy-Preserving Brain-Computer Interface Decoding
math.AG
  1. 1.
    Geometry of 3-dimensional del Pezzo fibrations in positive characteristic
  2. 2.
    The Mordell-Schinzel conjecture for cubic diophantine equations
  3. 3.
    The many faces of a logarithmic scheme
  4. 4.
    Lorentzian polynomials and the incidence geometry of tropical linear spaces
  5. 5.
    $p$-adic Local Langlands Correspondence
  6. 6.
    Real del Pezzo surfaces without points
  7. 7.
    Linearization problem for finite subgroups of the plane Cremona group
  8. 8.
    Groupes de monodromie finie des variétés abéliennes
  9. 9.
    Duality for Arithmetic $p$-adic Pro-étale Cohomology of Analytic Spaces
  10. 10.
    The external activity complex of a pair of matroids
math.AT
  1. 1.
    Digital $n-$Manifolds With Or Without Boundaries
  2. 2.
    Spatiotemporal Persistence Landscapes
  3. 3.
    Simplifications of finite spaces equipped with sheaves
  4. 4.
    Rational homotopy theory of operad modules through colored operads
  5. 5.
    Algebraic Topology Without Open Sets: A Net Approach to Homotopy Theory in Limit Spaces
  6. 6.
    The geometry of simplicial distributions on suspension scenarios
  7. 7.
    On the Last Kervaire Invariant Problem
  8. 8.
    Machine Proofs for Adams Differentials and Extension Problems among CW Spectra
  9. 9.
    Finite asymptotic dimension and the coarse assembly map
  10. 10.
    Modeling $(\infty,1)$-categories with Segal spaces
math.AP
  1. 1.
    Decay estimates for massive Dirac equation in a constant magnetic field
  2. 2.
    Semiclassical measure of the propagation between two topological insulators
  3. 3.
    Convex waves grazing convex obstacles to high order
  4. 4.
    A Note on Hyperbolic Relaxation of the Navier-Stokes-Cahn-Hilliard system for incompressible two-phase flow
  5. 5.
    Positive solutions to general semilinear overdetermined boundary problems
  6. 6.
    Capacitary measures in fractional order Sobolev spaces: Compactness and applications to minimization problems
  7. 7.
    Validity of the stochastic Landau approximation for super-pattern forming systems with a spatial 1:3 resonance
  8. 8.
    Spectral bounds for the operator pencil of an elliptic system in an angle
  9. 9.
    Infinite dimensional invariant tori for nonlinear Schrödinger equations
  10. 10.
    A Serrin-type over-determined problem for Hessian equations in the exterior domain
math.CT
  1. 1.
    Open Condensed Subgroups and Mackey’s Formula
  2. 2.
    The Relational Quotient Completion
  3. 3.
    Classification of localizing subcategories along t-structures
  4. 4.
    Categorification of modules and construction of schemes
  5. 5.
    Rational RG flow, extension, and Witt class
  6. 6.
    Intrinsically Correct Sorting in Cubical Agda
  7. 7.
    Single and multi-valued Hilbert-bundle renormings
  8. 8.
    Extended (tri)dendriform algebras, pre-Lie algebras and post-Lie algebras as companion structures of extended Rota-Baxter algebras
  9. 9.
    On The Telescopic Picard Group
  10. 10.
    Enhanced 2-categorical structures, two-dimensional limit sketches and the symmetry of internalisation
math.GR
  1. 1.
    $F$-birestriction monoids in enriched signature
  2. 2.
    Linearization problem for finite subgroups of the plane Cremona group
  3. 3.
    Salter’s question on the image of the Burau representation of $B_{4}$
  4. 4.
    Averaging groups
  5. 5.
    Enumerating Diagonalizable Matrices over $\mathbb{Z}_{p^{k}}$
  6. 6.
    The scale function for locally compact groups acting on non-positively curved spaces
  7. 7.
    A computational study of certain Weyl modules for type $G_{2}$ in characteristic 2
  8. 8.
    Left-Invariant Riemannian Distances on Higher-Rank Sol-Type Groups
  9. 9.
    Growth Rate Gap for Stable Subgroups
  10. 10.
    Computing Young’s Natural Representations for Generalized Symmetric Groups
math.NT
  1. 1.
    The Mordell-Schinzel conjecture for cubic diophantine equations
  2. 2.
    Simultaneous and multiplicative Diophantine approximation on missing-digit fractals
  3. 3.
    Codes from $A_{m}$-invariant polynomials
  4. 4.
    Generalised Fermat equation: a survey of solved cases
  5. 5.
    Groupes de monodromie finie des variétés abéliennes
  6. 6.
    Vanishing of Witten zeta function at negative integers
  7. 7.
    Popa’s "Recurrent Sequences" and Reciprocity
  8. 8.
    Duality for Arithmetic $p$-adic Pro-étale Cohomology of Analytic Spaces
  9. 9.
    About Eisenstein’s Theorem
  10. 10.
    On the packing dimension of weighted singular matrices on fractals
math.OC
  1. 1.
    Bilevel Learning with Inexact Stochastic Gradients
  2. 2.
    Memory-Reduced Meta-Learning with Guaranteed Convergence
  3. 3.
    On Differential Stability of a Class of Convex Optimization Problems
  4. 4.
    Convergence of trust-region algorithms in compact metric spaces
  5. 5.
    Eckstein-Ferris-Pennanen-Robinson duality revisited: paramonotonicity, total Fenchel-Rockallar duality, and the Chambolle-Pock operator
  6. 6.
    Capacitary measures in fractional order Sobolev spaces: Compactness and applications to minimization problems
  7. 7.
    A monotone block coordinate descent method for solving absolute value equations
  8. 8.
    Bivariate rational approximations of the general temperature integral
  9. 9.
    Toward a Unified Theory of Gradient Descent under Generalized Smoothness
  10. 10.
    A particle system approach towards the global well-posedness of master equations for potential mean field games of control
math.ST
  1. 1.
    Optimality of the Right-Invariant Prior
  2. 2.
    The entropic optimal (self-)transport problem: Limit distributions for decreasing regularization with application to score function estimation
  3. 3.
    Causal Invariance Learning via Efficient Optimization of a Nonconvex Objective
  4. 4.
    A partial likelihood approach to tree-based density modeling and its application in Bayesian inference
  5. 5.
    Dual Unscented Kalman Filter Architecture for Sensor Fusion in Water Networks Leak Localization
  6. 6.
    Learning Massive-scale Partial Correlation Networks in Clinical Multi-omics Studies with HP-ACCORD
  7. 7.
    Well-Posedness and Stability of the Stochastic OGTT Model
  8. 8.
    Posterior asymptotics of high-dimensional spiked covariance model with inverse-Wishart prior
  9. 9.
    Model checking for high dimensional generalized linear models based on random projections
  10. 10.
    The Stein-log-Sobolev inequality and the exponential rate of convergence for the continuous Stein variational gradient descent method
q-bio.BM
  1. 1.
    Category-Specific Topological Learning of Metal-Organic Frameworks
  2. 2.
    Applications of Knot Theory for the Improvement of the AlphaFold Protein Database
  3. 3.
    EquiFlow: Equivariant Conditional Flow Matching with Optimal Transport for 3D Molecular Conformation Prediction
  4. 4.
    FlowDock: Geometric Flow Matching for Generative Protein-Ligand Docking and Affinity Prediction
  5. 5.
    NeuralPLexer3: Physio-Realistic Biomolecular Complex Structure Prediction with Flow Models
  6. 6.
    COMET: Benchmark for Comprehensive Biological Multi-omics Evaluation Tasks and Language Models
  7. 7.
    Quadratic unconstrained binary optimization and constraint programming approaches for lattice-based cyclic peptide docking
  8. 8.
    High-dimensional Statistics Applications to Batch Effects in Metabolomics
  9. 9.
    Precise Antigen-Antibody Structure Predictions Enhance Antibody Development with HelixFold-Multimer
  10. 10.
    Sampling-based Continuous Optimization with Coupled Variables for RNA Design
q-bio.GN
  1. 1.
    BarcodeMamba: State Space Models for Biodiversity Analysis
  2. 2.
    VEPerform: a web resource for evaluating the performance of variant effect predictors
  3. 3.
    A robust, scalable K-statistic for quantifying immune cell clustering in spatial proteomics data
  4. 4.
    Can linguists better understand DNA?
  5. 5.
    A Misclassification Network-Based Method for Comparative Genomic Analysis
  6. 6.
    DNA Fragments in Crude Oil Reveals Earth’s Hidden History
  7. 7.
    Ancient DNA from 120-Million-Year-Old Lycoptera Fossils Reveals Evolutionary Insights
  8. 8.
    Emerging Challenges in Molecular Paleontology: Misapplication of Environmental DNA Fragments and Misconception of Deamination as a Key Criterion for In Situ DNA Identification
  9. 9.
    ProtGO: A Transformer based Fusion Model for accurately predicting Gene Ontology (GO) Terms from full scale Protein Sequences
  10. 10.
    DART-Eval: A Comprehensive DNA Language Model Evaluation Benchmark on Regulatory DNA
q-bio.QM
  1. 1.
    Deep-learning-based identification of individual motion characteristics from upper-limb trajectories towards disorder stage evaluation
  2. 2.
    Decoding Drug Discovery: Exploring A-to-Z In silico Methods for Beginners
  3. 3.
    BarcodeMamba: State Space Models for Biodiversity Analysis
  4. 4.
    FlowDock: Geometric Flow Matching for Generative Protein-Ligand Docking and Affinity Prediction
  5. 5.
    Reliable and superior elliptic Fourier descriptor normalization and its application software ElliShape with efficient image processing
  6. 6.
    MEATRD: Multimodal Anomalous Tissue Region Detection Enhanced with Spatial Transcriptomics
  7. 7.
    Cardiovascular Disease Detection By Leveraging Semi-Supervised Learning
  8. 8.
    Predictive Pattern Recognition Techniques Towards Spatiotemporal Representation of Plant Growth in Simulated and Controlled Environments: A Comprehensive Review
  9. 9.
    RAID-Database: human Responses to Affine Image Distortions
  10. 10.
    MiCull2 – simulating mastitis transmission through milking order
q-bio.PE
  1. 1.
    Asymmetric Interactions Shape Survival During Population Range Expansions
  2. 2.
    Quasispecies dynamics with time lags and periodic fluctuations in replication
  3. 3.
    Explicit modeling of density dependence in spatial capture-recapture models
  4. 4.
    Stochastic models in phylogenetic comparative methods: analytical properties and parameter estimation
  5. 5.
    Multivariate Aspects of Phylogenetic Comparative Methods
  6. 6.
    The expensive son hypothesis
  7. 7.
    Self-similarity in pandemic spread and fractal containment policies
  8. 8.
    Estimating excess mortality during the Covid-19 pandemic in Aotearoa New Zealand
  9. 9.
    An assessment of Alberta’s strategy for controlling mountain pine beetle outbreaks
  10. 10.
    Mountain pine beetle struggles with jack pine: A mechanistic explanation for slowed range expansion in Alberta
q-fin.CP
  1. 1.
    S&P 500 Trend Prediction
  2. 2.
    Simulation of square-root processes made simple: applications to the Heston model
  3. 3.
    From Votes to Volatility Predicting the Stock Market on Election Day
  4. 4.
    SusGen-GPT: A Data-Centric LLM for Financial NLP and Sustainability Report Generation
  5. 5.
    FinGPT: Enhancing Sentiment-Based Stock Movement Prediction with Dissemination-Aware and Context-Enriched LLMs
  6. 6.
    Reciprocity in Interbank Markets
  7. 7.
    Integrative Analysis of Financial Market Sentiment Using CNN and GRU for Risk Prediction and Alert Systems
  8. 8.
    Financial Fine-tuning a Large Time Series Model
  9. 9.
    Geometric Deep Learning for Realized Covariance Matrix Forecasting
  10. 10.
    Isogeometric Analysis for the Pricing of Financial Derivatives with Nonlinear Models: Convertible Bonds and Options
q-fin.PM
  1. 1.
    Cost-aware Portfolios in a Large Universe of Assets
  2. 2.
    PolyModel for Hedge Funds’ Portfolio Construction Using Machine Learning
  3. 3.
    Geometric Deep Learning for Realized Covariance Matrix Forecasting
  4. 4.
    LLMs for Time Series: an Application for Single Stocks and Statistical Arbitrage
  5. 5.
    A Joint Energy and Differentially-Private Smart Meter Data Market
  6. 6.
    Smart leverage? Rethinking the role of Leveraged Exchange Traded Funds in constructing portfolios to beat a benchmark
  7. 7.
    Correlation without Factors in Retail Cryptocurrency Markets
  8. 8.
    Turnover of investment portfolio via covariance matrix of returns
  9. 9.
    MILLION: A General Multi-Objective Framework with Controllable Risk for Portfolio Management
  10. 10.
    Dynamic ETF Portfolio Optimization Using enhanced Transformer-Based Models for Covariance and Semi-Covariance Prediction(Work in Progress)
q-fin.TR
  1. 1.
    Auto-Regressive Control of Execution Costs
  2. 2.
    FinGPT: Enhancing Sentiment-Based Stock Movement Prediction with Dissemination-Aware and Context-Enriched LLMs
  3. 3.
    Efficient and Verified Continuous Double Auctions
  4. 4.
    A Joint Energy and Differentially-Private Smart Meter Data Market
  5. 5.
    A theory of passive market impact
  6. 6.
    Uncertain Regulations, Definite Impacts: The Impact of the US Securities and Exchange Commission’s Regulatory Interventions on Crypto Assets
  7. 7.
    Ergodic optimal liquidations in DeFi
  8. 8.
    MarketGPT: Developing a Pre-trained transformer (GPT) for Modeling Financial Time Series
  9. 9.
    Calculating Profits and Losses for Algorithmic Trading Strategies: A Short Guide
  10. 10.
    Market Making without Regret
stat.AP
  1. 1.
    But Can You Use It? Design Recommendations for Differentially Private Interactive Systems
  2. 2.
    Efficient Bayesian inversion for simultaneous estimation of geometry and spatial field using the Karhunen-Loève expansion
  3. 3.
    Chopin: An Open Source R-language Tool to Support Spatial Analysis on Parallelizable Infrastructure
  4. 4.
    Spatial Cross-Recurrence Quantification Analysis for Multi-Platform Contact Tracing and Epidemiology Research
  5. 5.
    P3LS: Point Process Partial Least Squares
  6. 6.
    Missing data imputation for noisy time-series data and applications in healthcare
  7. 7.
    Balancing Accuracy and Costs in Cross-Temporal Hierarchies: Investigating Decision-Based and Validation-Based Reconciliation
  8. 8.
    Statistical Problems in the Diagnosis of Shaken Baby Syndrome/Abusive Head Trauma: Limitations to Algorithms and the Need for Reliable Data
  9. 9.
    CESAR: A Convolutional Echo State AutoencodeR for High-Resolution Wind Forecasting
  10. 10.
    Cardiovascular Disease Detection By Leveraging Semi-Supervised Learning
stat.ML
  1. 1.
    Generalization Analysis for Deep Contrastive Representation Learning
  2. 2.
    Multiplex Dirichlet stochastic block model for clustering multidimensional compositional networks
  3. 3.
    BetaExplainer: A Probabilistic Method to Explain Graph Neural Networks
  4. 4.
    Bayesian Surrogate Training on Multiple Data Sources: A Hybrid Modeling Strategy
  5. 5.
    Scalable Temporal Anomaly Causality Discovery in Large Systems: Achieving Computational Efficiency with Binary Anomaly Flag Data
  6. 6.
    Conditional Diffusion Models Based Conditional Independence Testing
  7. 7.
    Generalized Bayesian deep reinforcement learning
  8. 8.
    A partial likelihood approach to tree-based density modeling and its application in Bayesian inference
  9. 9.
    A Mapper Algorithm with implicit intervals and its optimization
  10. 10.
    Learning Massive-scale Partial Correlation Networks in Clinical Multi-omics Studies with HP-ACCORD
stat.TH
  1. 1.
    Optimality of the Right-Invariant Prior
  2. 2.
    The entropic optimal (self-)transport problem: Limit distributions for decreasing regularization with application to score function estimation
  3. 3.
    Causal Invariance Learning via Efficient Optimization of a Nonconvex Objective
  4. 4.
    A partial likelihood approach to tree-based density modeling and its application in Bayesian inference
  5. 5.
    Dual Unscented Kalman Filter Architecture for Sensor Fusion in Water Networks Leak Localization
  6. 6.
    Learning Massive-scale Partial Correlation Networks in Clinical Multi-omics Studies with HP-ACCORD
  7. 7.
    Well-Posedness and Stability of the Stochastic OGTT Model
  8. 8.
    Posterior asymptotics of high-dimensional spiked covariance model with inverse-Wishart prior
  9. 9.
    Model checking for high dimensional generalized linear models based on random projections
  10. 10.
    The Stein-log-Sobolev inequality and the exponential rate of convergence for the continuous Stein variational gradient descent method
eess.IV
  1. 1.
    Are the Latent Representations of Foundation Models for Pathology Invariant to Rotation?
  2. 2.
    Towards Physically-Based Sky-Modeling
  3. 3.
    Ant Nest Detection Using Underground P-Band TomoSAR
  4. 4.
    Ensemble Learning and 3D Pix2Pix for Comprehensive Brain Tumor Analysis in Multimodal MRI
  5. 5.
    Point Cloud-Assisted Neural Image Compression
  6. 6.
    Flex-PE: Flexible and SIMD Multi-Precision Processing Element for AI Workloads
  7. 7.
    Fast-staged CNN Model for Accurate pulmonary diseases and Lung cancer detection
  8. 8.
    High-speed and High-quality Vision Reconstruction of Spike Camera with Spike Stability Theorem
  9. 9.
    Data-driven Precipitation Nowcasting Using Satellite Imagery
  10. 10.
    Block-Based Multi-Scale Image Rescaling
eess.SP
  1. 1.
    Rate-Splitting Multiple Access for Integrated Sensing and Communications: A First Experimental Study
  2. 2.
    Soil moisture estimation of bare and vegetation-covered areas using a P/L/C-band SAR
  3. 3.
    Ant Nest Detection Using Underground P-Band TomoSAR
  4. 4.
    Scalable Data Transmission Framework for Earth Observation Satellites with Channel Adaptation
  5. 5.
    Sonar-based Deep Learning in Underwater Robotics: Overview, Robustness and Challenges
  6. 6.
    Evaluating the Efficacy of Vectocardiographic and ECG Parameters for Efficient Tertiary Cardiology Care Allocation Using Decision Tree Analysis
  7. 7.
    Acceleration and Parallelization Methods for ISRS EGN Model
  8. 8.
    On-the-Fly Interrogation of Mobile Passive Sensors from the Fusion of Optical and Radar Data
  9. 9.
    Capacity Analysis on OAM-Based Wireless Communications: An Electromagnetic Information Theory Perspective
  10. 10.
    Probabilistic GOSPA: A Metric for Performance Evaluation of Multi-Object Filters with Uncertainties
econ.EM
  1. 1.
    Moderating the Mediation Bootstrap for Causal Inference
  2. 2.
    VAR models with an index structure: A survey with new results
  3. 3.
    Treatment Evaluation at the Intensive and Extensive Margins
  4. 4.
    Forecasting realized covariances using HAR-type models
  5. 5.
    Do LLMs Act as Repositories of Causal Knowledge?
  6. 6.
    An overview of meta-analytic methods for economic research
  7. 7.
    A Neyman-Orthogonalization Approach to the Incidental Parameter Problem
  8. 8.
    Geometric Deep Learning for Realized Covariance Matrix Forecasting
  9. 9.
    A Kernel Score Perspective on Forecast Disagreement and the Linear Pool
  10. 10.
    The Global Carbon Budget as a cointegrated system
econ.GN
  1. 1.
    Multiplexing in Networks and Diffusion
  2. 2.
    Transition dynamics of electricity asset-owning firms
  3. 3.
    Binary or nonbinary? An evolutionary learning approach to gender identity
  4. 4.
    On Prior Confidence and Belief Updating
  5. 5.
    Strategically Acting on Information
  6. 6.
    Is Polarization an Inevitable Outcome of Similarity-Based Content Recommendations? – Mathematical Proofs and Computational Validation
  7. 7.
    Re-examining the social impact of silver monetization in the Ming Dynasty from the perspective of supply and demand
  8. 8.
    Delving into Youth Perspectives on In-game Gambling-like Elements: A Proof-of-Concept Study Utilising Large Language Models for Analysing User-Generated Text Data
  9. 9.
    Does Low Spoilage Under Cold Conditions Foster Cultural Complexity During the Foraging Era? – A Theoretical and Computational Inquiry
  10. 10.
    Emulating the Global Change Analysis Model with Deep Learning
B.3 Additional Results for Explore Puzzles
We ran a robustness check on the first six open-ended question experiments using the universal sentence encoder of Cer [11] as our embedding model. Our results remain largely unchanged, and are summarized in Figure 13.
Figure 13: Results averaged over the first six questions, for embeddings generated using the universal sentence encoder.
Below are the individual plots for the remaining $9$ questions using the Sentence-BERT encoder.
Figure 14: Results on Question 1
Figure 15: Results on Question 2
Figure 16: Results on Question 3
Figure 17: Results on Question 4
Figure 18: Results on Question 5
Figure 19: Results on Question 6
Figure 20: Results on Question 7
Figure 21: Results on Question 8
Figure 22: Results on Question 9
Table 2: Performance comparison for all-at-once on open-ended questions.
K=1 K=2 K=3 K=4 K=5 K=7 K=10
Q0: 0.39 0.4 0.4 0.4 0.41 0.42 0.56
Q1: 0.63 0.66 0.66 0.65 0.63 0.64 0.57
Q2: 0.41 0.43 0.41 0.4 0.36 0.41 0.41
Q3: 0.5 0.49 0.6 0.67 0.67 0.66 0.78
Q4: 0.33 0.46 0.4 0.43 0.38 0.42 0.42
Q5: 0.39 0.41 0.47 0.3 0.31 0.42 0.43
Q6: 0.45 0.78 0.78 0.79 0.78 0.78 0.75
Q7: 0.84 0.84 0.84 0.84 0.84 0.83 0.83
Q8: 0.39 0.41 0.47 0.36 0.37 0.37 0.39
Q9: 0.47 0.46 0.43 0.46 0.48 0.45 0.44
Table 3: Performance comparison for all-at-once with encouragement on open-ended questions.
K=1 K=2 K=3 K=4 K=5 K=7 K=10
Q0: 0.4 0.4 0.34 0.35 0.37 0.52 0.55
Q1: 0.45 0.51 0.66 0.61 0.58 0.56 0.59
Q2: 0.36 0.4 0.41 0.4 0.4 0.37 0.43
Q3: 0.48 0.42 0.41 0.49 0.4 0.6 0.49
Q4: 0.5 0.52 0.52 0.51 0.5 0.51 0.49
Q5: 0.4 0.39 0.39 0.43 0.45 0.4 0.41
Q6: 0.74 0.67 0.79 0.79 0.77 0.55 0.64
Q7: 0.69 0.82 0.78 0.79 0.79 0.77 0.8
Q8: 0.45 0.5 0.47 0.56 0.55 0.42 0.51
Q9: 0.44 0.39 0.37 0.4 0.47 0.46 0.47
Table 4: Performance comparison for one-by-one on open-ended questions.
K=1 K=2 K=3 K=4 K=5 K=7 K=10
Q0: 0.33 0.4 0.4 0.43 0.43 0.47 0.52
Q1: 0.62 0.62 0.62 0.62 0.61 0.61 0.6
Q2: 0.43 0.43 0.44 0.43 0.43 0.44 0.45
Q3: 0.58 0.57 0.59 0.61 0.59 0.6 0.6
Q4: 0.36 0.41 0.47 0.48 0.47 0.46 0.45
Q5: 0.39 0.48 0.49 0.48 0.48 0.49 0.49
Q6: 0.83 0.81 0.82 0.81 0.81 0.8 0.79
Q7: 0.8 0.81 0.81 0.81 0.81 0.8 0.79
Q8: 0.38 0.42 0.44 0.44 0.47 0.5 0.5
Q9: 0.45 0.47 0.47 0.46 0.45 0.45 0.43
Table 5: Performance comparison for one-by-one with encouragement on open-ended questions.
K=1 K=2 K=3 K=4 K=5 K=7 K=10
Q0: 0.35 0.37 0.5 0.54 0.55 0.58 0.59
Q1: 0.55 0.58 0.6 0.61 0.61 0.61 0.62
Q2: 0.45 0.45 0.44 0.44 0.45 0.45 0.42
Q3: 0.55 0.55 0.54 0.53 0.53 0.53 0.54
Q4: 0.33 0.49 0.48 0.49 0.49 0.48 0.46
Q5: 0.4 0.43 0.44 0.44 0.45 0.45 0.45
Q6: 0.83 0.83 0.82 0.82 0.81 0.81 0.8
Q7: 0.75 0.8 0.81 0.82 0.81 0.82 0.82
Q8: 0.35 0.4 0.45 0.47 0.49 0.5 0.53
Q9: 0.49 0.49 0.5 0.49 0.5 0.47 0.48
Table 6: Performance comparison for random actions on open-ended questions.
K=1 K=2 K=3 K=4 K=5 K=7 K=10
Q0: -0.01 -0.0 0.01 0.01 0.03 0.03 0.02
Q1: 0.01 0.01 0.02 0.02 0.04 0.04 0.03
Q2: 0.01 -0.0 0.03 0.02 0.02 0.03 0.02
Q3: 0.05 0.04 0.01 0.05 0.04 0.03 0.04
Q4: 0.02 0.01 0.03 0.02 0.02 0.04 0.01
Q5: -0.02 -0.01 0.02 0.03 0.02 0.03 0.05
Q6: 0.0 0.0 0.03 0.03 0.04 0.02 0.03
Q7: 0.01 0.05 0.03 0.04 0.04 0.02 0.04
Q8: -0.04 0.03 0.04 0.04 0.03 0.03 0.04
Q9: 0.04 0.0 0.04 0.03 0.04 0.03 0.04
Table 7: Performance for all-at-once on arXiv tasks.
K=1 K=2 K=5
gr-qc 0.63 0.64 0.68
hep-ex 0.81 0.81 0.83
hep-lat 0.72 0.72 0.72
hep-ph 0.7 0.76 0.75
hep-th 0.65 0.71 0.73
math-ph 0.64 0.73 0.74
nucl-ex 0.73 0.79 0.75
nucl-th 0.65 0.69 0.71
quant-ph 0.68 0.71 0.75
cs.AI 0.66 0.71 0.72
cs.CL 0.66 0.71 0.75
cs.CV 0.72 0.74 0.71
cs.LG 0.68 0.72 0.74
cs.NE 0.71 0.78 0.78
cs.RO 0.76 0.79 0.78
cs.IT 0.72 0.73 0.72
cs.CR 0.7 0.72 0.74
cs.DS 0.75 0.77 0.77
cs.HC 0.75 0.75 0.75
math.AG 0.7 0.78 0.78
math.AT 0.68 0.7 0.71
math.AP 0.7 0.79 0.78
math.CT 0.65 0.69 0.71
math.GR 0.73 0.77 0.76
math.NT 0.73 0.79 0.77
math.OC 0.77 0.79 0.76
math.ST 0.63 0.68 0.73
q-bio.BM 0.81 0.85 0.83
q-bio.GN 0.76 0.78 0.79
q-bio.QM 0.76 0.78 0.78
q-bio.PE 0.8 0.82 0.8
q-fin.CP 0.74 0.78 0.77
q-fin.PM 0.74 0.77 0.78
q-fin.TR 0.74 0.78 0.78
stat.AP 0.73 0.69 0.75
stat.ML 0.7 0.73 0.74
stat.TH 0.65 0.67 0.78
eess.IV 0.67 0.73 0.72
eess.SP 0.65 0.67 0.72
econ.EM 0.62 0.68 0.7
econ.GN 0.66 0.67 0.74
Table 8: Performance for one-by-one on arXiv tasks.
K=1 K=2 K=5
gr-qc 0.63 0.64 0.65
hep-ex 0.78 0.76 0.81
hep-lat 0.72 0.72 0.74
hep-ph 0.7 0.72 0.73
hep-th 0.64 0.69 0.68
math-ph 0.65 0.71 0.71
nucl-ex 0.72 0.74 0.76
nucl-th 0.64 0.67 0.71
quant-ph 0.7 0.71 0.71
cs.AI 0.7 0.74 0.73
cs.CL 0.69 0.71 0.75
cs.CV 0.73 0.74 0.77
cs.LG 0.67 0.72 0.72
cs.NE 0.74 0.76 0.77
cs.RO 0.78 0.79 0.78
cs.IT 0.76 0.75 0.75
cs.CR 0.72 0.71 0.74
cs.DS 0.75 0.78 0.78
cs.HC 0.72 0.72 0.73
math.AG 0.68 0.77 0.78
math.AT 0.63 0.69 0.7
math.AP 0.72 0.75 0.76
math.CT 0.62 0.71 0.73
math.GR 0.71 0.76 0.74
math.NT 0.73 0.75 0.73
math.OC 0.73 0.77 0.8
math.ST 0.63 0.66 0.7
q-bio.BM 0.84 0.84 0.87
q-bio.GN 0.69 0.76 0.75
q-bio.QM 0.76 0.8 0.77
q-bio.PE 0.8 0.79 0.82
q-fin.CP 0.71 0.72 0.76
q-fin.PM 0.67 0.77 0.73
q-fin.TR 0.72 0.74 0.75
stat.AP 0.73 0.72 0.79
stat.ML 0.74 0.75 0.77
stat.TH 0.61 0.64 0.72
eess.IV 0.71 0.72 0.74
eess.SP 0.64 0.66 0.68
econ.EM 0.66 0.66 0.67
econ.GN 0.67 0.69 0.69
Table 9: Performance for Category Only baseline on arXiv tasks.
K=1 K=2 K=5
gr-qc 0.28 0.32 0.36
hep-ex 0.25 0.35 0.38
hep-lat 0.31 0.31 0.38
hep-ph 0.25 0.26 0.32
hep-th 0.24 0.26 0.27
math-ph 0.22 0.3 0.3
nucl-ex 0.37 0.38 0.35
nucl-th 0.31 0.32 0.33
quant-ph 0.23 0.27 0.27
cs.AI 0.15 0.17 0.18
cs.CL 0.14 0.19 0.26
cs.CV 0.19 0.22 0.32
cs.LG 0.21 0.19 0.24
cs.NE 0.27 0.31 0.3
cs.RO 0.27 0.27 0.31
cs.IT 0.29 0.31 0.34
cs.CR 0.21 0.24 0.32
cs.DS 0.2 0.2 0.22
cs.HC 0.12 0.2 0.23
math.AG 0.32 0.33 0.33
math.AT 0.33 0.34 0.4
math.AP 0.19 0.23 0.32
math.CT 0.25 0.23 0.29
math.GR 0.25 0.28 0.34
math.NT 0.21 0.29 0.32
math.OC 0.2 0.27 0.3
math.ST 0.14 0.22 0.18
q-bio.BM 0.17 0.31 0.29
q-bio.GN 0.2 0.25 0.32
q-bio.QM 0.03 0.12 0.1
q-bio.PE 0.28 0.32 0.31
q-fin.CP 0.34 0.36 0.37
q-fin.PM 0.39 0.39 0.43
q-fin.TR 0.33 0.34 0.38
stat.AP 0.07 0.12 0.1
stat.ML 0.19 0.21 0.28
stat.TH -0.01 0.07 0.12
eess.IV 0.18 0.19 0.24
eess.SP 0.21 0.23 0.24
econ.EM 0.21 0.32 0.37
econ.GN 0.13 0.19 0.19
Figure 23: Results for arXiv category econ.GN
Figure 24: Results for arXiv category eess.SP
Figure 25: Results for arXiv category q-bio.BM
Figure 26: Results for arXiv category math.ST
B.4 Benchmarking Encoders
Here we benchmark the two encoders we use (Sentence-BERT and the universal sentence encoder) by measuring the cosine similarity between semantically similar/different words.
Table 10: Cosine similarity of different words.
Sentence-BERT Universal Sentence Encoder
dog, tacos: 0.25 0.24
Pittsburgh, tiki bar: 0.12 0.17
Honolulu, tiki bar: 0.30 0.25
Pittsburgh, Honolulu: 0.41 0.29
angel, devil: 0.48 0.54
machine learning, artificial intelligence: 0.70 0.58
war, peace: 0.61 0.49
love, hate: 0.49 0.59
love, affection: 0.62 0.56
war, battle: 0.74 0.57
machine learning, battle: 0.25 0.19
The similarity scores of both models in Table 10 suggest that while the embeddings produced by both embedding models are generally “in the ballpark” of what one would consider “similar”/“different”, they are still a somewhat coarse measure of distance, which may explain the similar performance of our different prompting strategies.

### v2

@@
 1 Introduction
-There has been significant interest in the machine learning community to apply recent advances in generative AI and large language models (LLMs) to solve important decision-making problems.
-Early work in this direction has already produced impressive agentic behavior in both virtual [e.g.,  41, 32] and physical-world environments [e.g.,  7].
-However, significant challenges remain.
+There’s been significant interest in the machine learning community to apply recent advances in generative AI and large language models (LLMs) to solve important decision-making problems.
+Early work in this direction has already produced impressive agentic behavior in both virtual (e.g.,  Wang et al., 2023; OpenAI, 2025) and physical-world environments (e.g.,  Black et al., 2024).
 Beyond generalization (needed for supervised learning), decision-making under uncertainty requires two additional capabilities: exploitation (making the best decision given the current data) and exploration (trying new options for long-term benefit).
-Balancing the two has been a subject of an enormous literature, see books [e.g.,  see books 38, 25, 3].
-A recent line of work [e.g.,  23, 31] evaluates the ability of LLMs to balance exploration and exploitation entirely in-context, i.e.,  by specifying the problem description, parameters, and history in the LLM prompt.
-Focused on simple tasks in reinforcement learning, these results have been mixed.
-Both papers show that LLMs fail to solve these tasks adequately out-of-the-box, but they can be prompted to do so, e.g.
-by providing various summary statistics in-context, or by fine-tuning on data from algorithmic baselines on similar problem instances.
-Unfortunately, such mitigations do not readily extend beyond simple settings.
-For example, succinct summary statistics do not exist for many decision-making tasks, and fine-tuning for specific tasks may be prohibitive due to cost or lack of sufficient training data.
-Motivated by these observations, we study the ability of Gpt-4 [2], Gpt-4o [18], and Gpt-3.5 [8] to explore and exploit in-context in silos, with an eye towards leveraging a pre-trained LLM (and the inductive bias therein) as a part of a larger decision-making agent.
+Balancing the two has led to an enormous literature (e.g.,  see books Slivkins, 2019; Lattimore and Szepesvári, 2020; Agarwal et al., 2020).
+A recent line of work (e.g.,  Krishnamurthy et al., 2024; Nie et al., 2024) evaluates the ability of LLMs to balance exploration and exploitation entirely in-context, i.e.,  specifying the problem description, parameters, and history in the LLM prompt.
+Focused on simple tasks in reinforcement learning (RL), these results are mixed.
+Both papers show that LLMs fail to solve these tasks adequately out-of-the-box, but they can be prompted to do so by providing succinct summary statistics in-context.
+However, such statistics do not exist beyond simple decision-making problems, e.g.,  for contextual bandits.
+RL-specific pre-training or fine-tuning (on data from algorithmic baselines on similar problem instances) tends to work well (see Related Work), but may be prohibitive due to cost or insufficient training data.
+Besides, using a commonly available frontier LLM would leverage its “generalist” intelligence and may be much easier in terms of logistics and required expertise.
+Motivated by these observations, we study the ability of LLMs to explore and exploit in-context in silos, with an eye towards leveraging a pre-trained LLM (and the inductive bias therein) as a part of a larger decision-making agent.
 We focus on (contextual) bandits, as a standard abstraction for the explore-exploit tradeoff.
+We experiment with a variety of LLMs: Gpt-4, Gpt-4o, Gpt-3.5, Qwen-2.5, Gemma-3, Mistral-7B, as well as DeepSeek-R1, a reasoning model.
+[^2] Several other LLMs did not work for our purposes: Llama-3.2-3B and Llama-3-8B would not follow instructions, as well as DeepSeek-R1-Distill-Llama-70B (a reasoning model); Qwen-2.5-Math-70B (another reasoning model that we tried) has a short context window.
+We restrict our GPT access to Gpt-4, Gpt-4o, and Gpt-3.5, as newer variants such as Gpt-4.1 and Gpt-4.5 are primarily optimized for efficiency (e.g.
+latency, cost, coding) rather than broad general reasoning, and Gpt-5 uses a reasoning backbone as a routing mechanism across sub-models.
 In Section 2, we evaluate LLMs as exploitation oracles for contextual bandits.
@@
 We find that in-context summary techniques are useful for improving performance, but LLMs with these mitigations still perform worse than a simple linear regression baseline.
-In Section 4, we evaluate LLMs as exploration oracles which suggest candidate actions within a large action space.
-To do so, we introduce a text-based multi-armed bandit task, where actions correspond to free-text answers to an open-ended question, and rewards are driven by the distance from some preselected answer in an embedding space (where the embedding is computed exogenously).
-Given the high dimensionality of the action space, traditional bandit approaches are inapplicable.
+In Section 3, we evaluate LLMs as exploration oracles which suggest candidate actions by discretizing a large action space.
+To do so, we introduce a text-based multi-armed bandit task, where actions correspond to free-text answers to an open-ended question, and rewards are driven by the distance from some preselected answer in an embedding space (the latter is computed exogenously).
+Given the high dimensionality of the action space, traditional discretization approaches typically used in continuous bandit problems are inapplicable.
 However, an LLM can generate a small set of candidate actions which can then be used to instantiate an off-the-shelf bandit algorithm.
@@
 Finally, we repeat our experiments on a larger-scale bandit task based on paper titles and abstracts from arXiv (where the goal is to find a suitable title for a given abstract), with similar findings.
-2 Related Work and Background
-Our results belong to a small, but quickly growing line of work on using pre-trained LLMs for in-context reinforcement learning (RL).
+Related work.
+Our results belong to a growing line of work on using pre-trained LLMs for in-context reinforcement learning (RL).
 Coda-Forno et al.
-[13], Krishnamurthy et al.
-[23], Nie et al.
-[31], Monea et al.
-[30], Xia et al.
-[43], Park et al.
-[33], Wu et al.
-[42] evaluate the ability of LLMs to solve various multi-armed bandit and contextual bandit tasks, and find that the current generation of LLMs largely fail to solve these tasks in-context.
+(2023); Krishnamurthy et al.
+(2024); Nie et al.
+(2024); Monea et al.
+(2024); Xia et al.
+(2024); Park et al.
+(2024); Wu et al.
+(2024) evaluate the ability of LLMs to solve various multi-armed bandit and contextual bandit tasks, and find that the current generation of LLMs largely fail to solve these tasks in-context.
 Indeed, positive findings are restricted to very simple tasks and/or require substantial mitigations (which in turn do not readily extend beyond simple settings).
 Xia et al.
-[43] use LLMs to solve dueling bandit tasks, and Park et al.
-[33] also evaluate the ability of LLMs to learn in games.
-While our paper is primarily concerned with whether LLMs succeed as algorithms, several others [e.g.,  36, 17, 14] use in-context bandits (and many other tasks) to study whether LLMs exhibit human-like behavior/biases in decision-making.
-A broader literature on in-context learning [8] aims to solve various tasks by providing all relevant information in the LLM prompt.
-The work on exemplar selection (selecting examples and other information to present in-context) [e.g.,  19, 48, 44, 40] is relevant to our exploitation experiments.
-A growing line of work aims to use LLMs as a part of a larger decision-making agent [e.g.,  28, 50, 49].
-Our exploration experiments take inspiration from the work on using LLMs as “action priors” inside of a larger RL algorithm [46, 10, 47, 16].
+(2024) use LLMs to solve dueling bandit tasks, and Park et al.
+(2024) also evaluate the ability of LLMs to learn in games.
+While our paper is primarily concerned with whether LLMs succeed as algorithms, several others (e.g.,  Schubert et al., 2024; Hayes et al., 2024; Coda-Forno et al., 2024) use in-context bandits (and many other tasks) to study whether LLMs exhibit human-like behavior/biases in decision-making.
+A broader literature on in-context learning (starting from Brown et al., 2020) aims to solve various tasks by providing all relevant information in the LLM prompt.
+The work on exemplar selection (selecting examples and other information to present in-context) (e.g.,  Khalifa et al., 2023; Zhang et al., 2022; Xiong et al., 2023; Tonglet et al., 2023) is relevant to our exploitation experiments.
+A growing line of work aims to use LLMs as a part of a larger decision-making agent (e.g.,  Li et al., 2024; Zhou et al., 2023; Zhao et al., 2024).
+Our exploration experiments take inspiration from the work on using LLMs as “action priors” inside of a larger RL algorithm (Yan et al., 2024; Carta et al., 2023; Yao et al., 2024; Hao et al., 2023).
 Much of this work falls under the proposer-verifier framework of Snell et al.
-[39], where an LLM proposes several possible sequences from which a verifier selects suitable candidates.
+(2024), where an LLM proposes several possible sequences from which a verifier selects suitable candidates.
 In comparison, our goal is a more systematic evaluation of LLMs’ abilities to explore large action spaces, in isolation from other components of the decision-making task.
-Finally, a parallel line of work trains transformers from scratch to solve various RL tasks [e.g.,  24, 26, 34, 45, 27].
-Multi-armed bandits.
-We consider tasks based on multi-armed bandits (MAB) and contextual bandits (CB), well-studied special cases of RL that abstract the explore-exploit tradeoff, see books [38, 25] for background.
+Finally, a parallel line of work trains transformers to solve various RL tasks (e.g.,  Laskin et al., 2022; Lin et al., 2023; Lee et al., 2024; Raparthy et al., 2023; Xu et al., 2022; Lehnert et al., 2024; Mukherjee et al., 2024).
+Our negative results in Section 2 provide an additional foundation for this work, emphasizing the shortcomings on frontier LLMs that are not custom-trained for RL.
+Background on multi-armed bandits (MAB).
+We consider tasks based on MAB and contextual bandits (CB), well-studied special cases of RL that abstract the explore-exploit tradeoff, see Slivkins (2019); Lattimore and Szepesvári (2020) for background.
 In MAB, there are $T$ rounds and $K$ arms.
-In each round $t\in[T]$, the learner chooses an action (arm) $a_{t}\in[K]$ and observes a noisy reward $r_{t}$ drawn from some sub-Gaussian reward distribution for this arm.
+In each round $t\in[T]$, the learner chooses an action (arm) $a_{t}\in[K]$ and observes reward $r_{t}$ drawn from some sub-Gaussian reward distribution for this arm.
 The reward distribution, and particularly its mean $\mu(a_{t})$, are unknown to the algorithm.
@@
 The learner’s goal is to balance exploration and exploitation to maximize cumulative reward.
-An “exploitation oracle" (which optimizes for the current round given the history) naturally plugs into standard bandit algorithms such as Epsilon-Greedy, Explore-then-Commit, and Follow-The-Perturbed-Leader.
-Typical implementations in CB involve model-based (e.g.,  linear) regression or cost-sensitive classification [38, Ch.8].
+An “exploitation oracle” (which optimizes for the current round given the history) naturally plugs into standard bandit algorithms such as Epsilon-Greedy, Explore-then-Commit, and Follow-The-Perturbed-Leader.
+Typical implementations in CB involve model-based (e.g.,  linear) regression or cost-sensitive classification (Slivkins, 2019, Ch.8).
 Designing CB exploitation oracles for their own sake, a.k.a.
-offline policy optimization, is well-studied [starting from, e.g.,  6, 15].
-An “exploration oracle" can suggest a set of candidate actions to discretize a large action space and initialize an off-the-shelf bandit algorithm.
-Ideally, the oracle would "zoom in" on a lower-dimensional subspace with better-performing actions.
-Our intuition for exploration oracles is grounded in the study of bandits with Lipschitz-like structure [38, Ch.4].
-In particular, the size vs.
-quality tradeoff in the discretization and the importance of “zooming in" are key issues in this literature [e.g.,  22, 20, 21, 9].
+offline policy optimization, is well-studied (starting from, e.g.,  Beygelzimer and Langford, 2009; Dudík et al., 2014).
+Usage of an LLM as an “exploration oracle” is closely related to the literature on dynamic pricing and Lipschitz bandits (e.g.,  Kleinberg and Leighton, 2003; Kleinberg et al., 2008, 2019; Bubeck et al., 2011), as we elaborate in Section 3.
 2 LLMs as exploitation oracles
-In this section, we evaluate the ability of LLMs to exploit in decision-making tasks with statistical uncertainty on the outcomes.
+Figure 1: MAB exploit puzzle for Gpt-4 (left), Gpt-4 with CoT (middle), and Gpt-3.5 with CoT (right), all with ”buttons” prompt.
+The following conventions apply to all figures in this section.
+Each line corresponds to a particular value of #rounds $T$ and plots $\mathrm{FracCorrect}(\epsilon,T)$ against empirical gap $\epsilon$ on the X-axis.
+The shaded band around the line represents a $95\%$ confidence interval.
+The dashed line is the number of tasks (“runs”) with empirical gap $\leq\epsilon$; the resp.
+Y-scale is on the right.
+We evaluate the ability of LLMs to exploit in decision-making tasks with statistical uncertainty on the outcomes.
 We present LLMs with in-context exploit tasks inspired by multi-armed bandits (MAB) and contextual bandits (CB).
@@
 These tasks are generated from some parameterized distributions called exploit puzzles.
-Our visualizations plot the fraction of correct answers against the task difficulty.
-3.1 MAB Exploit Puzzles
-Figure 1: MAB exploit puzzle for Gpt-4 (left), Gpt-4 with CoT (middle), and Gpt-3.5 with CoT (right), all with "buttons" prompt.
-The following conventions apply to all figures in this section.
-Each line corresponds to a particular value of $T$ and plots $\mathrm{FracCorrect}(\epsilon,T)$ against empirical gap $\epsilon$ on the X-axis.
-The shaded band around the line represents a $95\%$ confidence interval.
-The dashed line is the number of tasks (“runs") with empirical gap at most $\epsilon$; the corresponding Y-scale is on the right.
+As our results here are largely negative, we focus on relatively small tasks.
+We use synthetic data, to better control task size/difficulty and trace the “boundary of tractability”.
+We find that LLMs are consistently outperformed by a simple linear regression baseline, even on moderately-sized tasks.
+Due to computational constraints, our experiments focused on Gpt-4, Gpt-4o, and Gpt-3.5, with additional LLMs evaluated as a robustness check in Appendix A.4, with similar findings.
+MAB exploit puzzles.
+Our MAB-based experiments on Gpt-4 and Gpt-3.5 provide a partial explanation for why the current LLMs fail to solve MAB tasks in-context when presented with raw (non-summarized) history, as first observed by Krishnamurthy et al.
+(2024); Nie et al.
+(2024).
+Following these two papers, we try two prompts: one in which arms correspond to pushing different colored buttons and one where they correspond to showing different advertisements to users.
+The LLM is asked to choose the arm with the highest empirical reward in the next round.
+We also try chain-of-thought (CoT) prompts, for the total of 4 prompt designs: $\left\{\,\text{buttons, adverts}\,\right\}\times\left\{\,\text{CoT, no-CoT}\,\right\}$.
+See Appendix A for more details on our experimental setup.
 […]
-[…]
-Once an exploit task is generated, we measure its difficulty via empirical gap: the difference between the largest and second-largest average reward $\bar{r}_{t}(a):=\frac{1}{T}\sum_{t\in[T]}r_{t}(a)$ among all arms $a$.
-Intuitively, the smaller its empirical gap, the more difficult is the puzzle.
+Given a (realized) exploit task, we measure its difficulty via empirical gap $\epsilon$: the difference between the largest and second-largest average reward $\bar{r}_{t}(a):=\frac{1}{T}\sum_{t\in[T]}r_{t}(a)$ among all arms $a$.
+Intuitively, puzzle difficulty decreases with $\epsilon$ (as one needs to distinguish the top two arms).
+Empirical gap mirrors the “gap” between the top two expected rewards, a standard notion of difficulty in MAB.
 Let $S(\epsilon,T)$ be the set of all tasks with empirical gap at most $\epsilon$ and history size $T$.
-We measure an LLM’s performance over a given set $S$ of tasks as the fraction of tasks for which the LLM returns a “correct answer": an arm with the largest empirical reward; denoted by $\mathrm{FracCorrect}(S)$.
+We measure an LLM’s performance over a given set $S$ of tasks as the fraction of tasks for which the LLM returns a “correct answer”: an arm with the largest empirical reward; denoted by $\mathrm{FracCorrect}(S)$.
 We are interested in how $\mathrm{FracCorrect}$ varies depending on the difficulty level.
@@
 Performance tends to degrade (1) as the history length $T$ increases and (2) as the empirical gap decreases.
-While Gpt-4 generally performs much better than Gpt-3.5, we found that prompting the LLM to use chain-of-thought (CoT) reasoning provided a slight boost for Gpt-3.5, while hurting the performance of Gpt-4.
+(Our intuition for the former is that the LLM has more data points to process in context, and for the latter is that the bandit instance becomes more difficult.) While Gpt-4 generally performs much better than Gpt-3.5, we found that prompting the LLM to use chain-of-thought (CoT) reasoning provided a slight boost for Gpt-3.5, while hurting the performance of Gpt-4.333From CoT logs, we found that Gpt-4 was doing incorrect reasoning in a non-trivial fraction of runs.
 We found that performance was similar across all of our prompt designs (see Appendix A for more plots).
-3.2 CB Exploit Puzzles and Mitigations
-Figure 2: Gpt-4 succeeds on a small CB exploit puzzle (left) and fails on a slightly larger one (right).
-Figure 3: CB exploit puzzle with $d=K=2$ and $T=4000$: mitigations help substantially.
-Gpt-4 without CoT (left) and Gpt-4 with CoT (right).
-Note that providing the full history with this $T$ vastly exceeds the context window for Gpt-4, Gpt-4o, and Gpt-3.5.
-While the history in $K$-armed bandits can be summarized using $2K$ numbers (the average reward for each arm and number of times it has been played), such succinct summary statistics may not be readily available (or even exist) in more complicated decision-making tasks such as CB.
+Figure 2: Gpt-4 succeeds on a small CB exploit puzzle (left), but fails on a slightly larger one (right).
+CB exploit puzzles and mitigations.
+While the history in $K$-armed bandits can be summarized using $2K$ numbers (for each arm, the average reward and #plays) such succinct summary statistics may not be readily available (or even exist) in more complicated decision-making tasks such as CB.
 [^4] This consideration also motivates MAB exploit puzzles with raw (non-summarized) history, as a simpler special case of the general scenario when succinct summarization is unavailable.
-We focus on linear CB, a well-studied CB model [starting from 29, 12, 1] in which the expected reward of each arm is linear in the context.
-Specifically, given context $z\in\mathbb{R}^{d}$ and arm $a$, the expected reward is $\mu(z,a)=\langle z,\theta^{*}_{a}\rangle$, for some fixed (but unknown) parameters $\theta^{*}_{a}\in\mathbb{R}^{d}$.
-We consider a CB exploit puzzle parameterized by the number of arms $K$, context dimension $d$, and history size $T$.
+We focus on linear CB, where the expected reward of each arm $a$ is linear in the context $z\in\mathbb{R}^{d}$: $\mu(z,a)=\langle z,\theta^{*}_{a}\rangle$ for some fixed (but unknown) parameters $\theta^{*}_{a}\in\mathbb{R}^{d}$.
+[^5] Linear CB are well-studied, starting from Li et al.
+(2010); Chu et al.
+(2011); Abbasi-Yadkori et al.
+(2011).
+(We try non-linear CB, too.)
+We consider a CB exploit puzzle parameterized by #arms $K$, dimension $d$, and history size $T$.
 The tasks are constructed as follows.
@@
 Contexts $z_{t}$ are sampled independently and u.a.r.
-from $[-1,1]^{d}$.
-For simplicity, the history contains rewards of all arms $a$ in each round $t$, where the reward is generated as $\mu(z_{t},a)$ plus an independent unit-variance Gaussian.
-Given the history and a new context $z_{T+1}$ (drawn in the same way), the LLM is asked to select the action which appears best.
+from $1,1]^{d}$.
+For simplicity, the history contains rewards of all arms $a$ in each round $t$, where the reward equals $\mu(z_{t},a)$ plus an independent unit-variance Gaussian.
+Given the history and a new context $z_{T+1}$ (drawn in the same way), the LLM is asked to select the action for round $T+1$ which appears best.
 This gives one exploit task.
 We generate $N$ tasks for the same $K,T,d$.
-We present these exploit tasks to the LLM in-context using a modified “buttons” prompt, where contexts corresponds to “numbers on a screen" which affect the payoffs for pressing each button.
+We use a modified “buttons” prompt, where contexts correspond to “numbers on a screen” which affect the payoffs for each button.
 The prompt does not mention linearity of the CB instance (because such model-based information is typically unavailable in applications).
-For a particular exploit task, we now define a “correct answer" as an arm $a$ which maximizes expected reward $\mu(z_{T+1},a)$.
+Given an exploit task, we now define a “correct answer” as an arm $a$ which maximizes expected reward $\mu(z_{T+1},a)$.
 [^6] Note that it is unclear how to define an ”empirically best arm” given a CB history and the current context.
-As before, $\mathrm{FracCorrect}(S)$ is the fraction of correct answers in a given set $S$ of tasks.
+$\mathrm{FracCorrect}(S)$ is the fraction of correct answers in a given set $S$ of tasks.
 Likewise, the task difficulty is not easily defined in terms of the realized rewards.
@@
 We plot $\mathrm{FracCorrect}(\epsilon,T):=\mathrm{FracCorrect}\left(\,S(\epsilon,T)\,\right)$ against $\epsilon$.
-Figure 4: CB exploit puzzle with $d=K=5$ and $T=1000$: mitigations perform badly, but (mostly) much better than the no-mitigation baseline. Gpt-4o without CoT.
-[…]
+First, we find that Gpt-4 obtains near-perfect performance on a “small” puzzle with $K=2$ arms, context dimension $d=1$, and history size $T=50$ (Figure 2).
+However, its performance degrades as the problem size increases: see Figure 2 for $K=d=2$ and $T=100$.
+Moreover, limited prompt size may prevent processing larger histories.
+[^7] E.g.,  our LLM access points bottomed out at $T\approx\text{100-200}$ for Gpt-4 and $T\approx\text{1000-2000}$ for Gpt-4o.
+Figure 3: CB exploit puzzle with $d=K=2$ and $T=4000$: mitigations help substantially.
+Gpt-4 without CoT (left) and Gpt-4 with CoT (right).
+Note that providing the full history with this $T$ vastly exceeds the context window for Gpt-4, Gpt-4o, and Gpt-3.5.
 Motivated by these observations, we implement several natural mitigations inspired by the literature on exemplar selection for in-context learning (discussed in Related Work).
-The mitigations are as follows:
 1.
@@
 Report $\left(\,z^{*},\,a,\,\bar{r}(z^{*},a)\,\right)$ as a context-arm-reward triple, for each arm $a$ and each centeroid context $z^{*}$ among the $k^{\prime}$ centroids closest to $z_{T+1}$ (according to the $\ell_{2}$ metric).
-In all three mitigations, we do not explain the “mitigation strategy" in the prompt.
-That is, we present the reported context-arm-reward tuples as if it were the entire history, not mentioning nearest neighbors, clustering, or averaging.
+We do not explain the “mitigation strategy” in the prompt: we present the reported context-arm-reward tuples as if it were the entire history, without mentioning clustering or averaging.
 Figure 3 visualizes the performance of these mitigations on a slightly larger (but still relatively small) puzzle with $K=d=2$ and $T=4000$.
 We use Gpt-4 with and without CoT.
-We compare the mitigations against linear regression baseline (which is effectively an "upper bound", as the underlying CB instance is linear).
+We compare the mitigations against linear regression baseline (which is effectively an ”upper bound”, as the underlying CB instance is linear).
 Without CoT prompting, we find that all three mitigations achieve $\mathrm{FracCorrect}$ around $80\%-90\%$, although this dips to around $60\%-85\%$ when using CoT.
 In addition to (potentially) improving performance, mitigations can also offer a practical way to solve decision-making tasks using LLMs when the history is large; when $T=4,000$, our prompt vastly exceeds the context window of all models we had access to.
-Figure 5: Gpt-4o on the text-based CB exploit puzzle: some mitigations help, but are outperformed by linear regression.
+Figure 4: CB exploit puzzle with $d=K=5$ and $T=1000$: mitigations perform badly, but (mostly) much better than the no-mitigation baseline. Gpt-4o without CoT.
 However, current LLMs struggle to exploit on even moderately-sized problems, even with these mitigations.
@@
 Both our $k$-nearest and $k$-means, $k^{\prime}$-nearest mitigations significantly out-perform unmitigated Gpt-4o, but fall significantly short of the linear baseline.
-3.3 CB Exploit Puzzles (text-based)
+Figure 5: Left: Performance of Deepseek-R1 on our numerical CB puzzle. Right: Gpt-4o on the text-based CB exploit puzzle. Some mitigations help, but are outperformed by linear regression.
+Reasoning models are LLMs specifically trained to follow multi-step chains of thought to solve complex problems that require intermediate reasoning.
+We ran one such model, DeepSeek-R1-Distill-Qwen-32B (henceforth DeepSeek-R1), a 32B parameter reasoning model distilled from Qwen-2.5.
+Our results are shown in Figure 5 (left).
+As before, we find that DeepSeek-R1 fails to match the performance of the baseline, even when using the mitigations.
+High inference costs and small context windows prevented us from running larger reasoning models.
+CB exploit puzzles (text-based & non-linear).
 As a robustness check, we repeat our CB experiments on a text-based exploit puzzle.
@@
 See Appendix A for full details on our experimental setup.
-Figure 5 shows the performance of Gpt-4o (with mitigations) on this puzzle.
+Figure 5 (right) shows the performance of Gpt-4o (with mitigations) on this puzzle.
 While the reward function is non-linear (and thus the linear baseline only achieves  $70\%$ $\mathrm{FracCorrect}$), we find that all configurations are still significantly out-performed by the linear baseline.
+Our intuition for these findings is as follows: While LLMs are generally good with textual inputs, the reward data is numerical, so the LLM faces most/all of the same challenges as before (i.e.,  choosing a “solution strategy” and executing it correctly), exacerbated by non-linearity.
+Meanwhile, linear regression is known to often work fairly well even on non-linear data.
+Takeaways.
+While unmitigated performance drops as $d,K,T$ grow, increases in $T$ can largely be taken care of by our mitigations, whereas increases in $d$ and $K$ cannot.
+Indeed, taken together, our findings in this section suggest that while the current generation of LLMs are effective at exploiting in small-scale decision-making tasks, they struggle in larger settings.
 3 LLMs as exploration oracles
@@
 The shaded regions represent a $95\%$ confidence interval.
-Figure 7: Arm histogram for one particular question (“What is the meaning of freedom?”, left) and averaged over all questions (right).
-We consider $K=10$ suggestions generated by our one-by-one prompt without encouragement.
-For a given “run" of the candidate selection, the suggestions are ranked by the expected reward, and then the $i$-th expected reward is averaged over all runs, for each $i\in[K]$.
-We now turn our attention to the ability of pre-trained LLMs to explore large action spaces.
-Specifically, we leverage the inductive bias of an LLM to generate a small set of candidate actions from a text-based action space, before running an off-the-shelf MAB algorithm on top of the candidate set.
-We consider two types of exploration tasks: answering an open-ended “philosophical" question (Q/A task) and suggesting a title for an arXiv research paper based on its abstract (arXiv task).
+We now study the ability of LLMs to explore large action spaces.
+We leverage the inductive bias of an LLM to generate a small set of candidate actions from a text-based action space, before running an off-the-shelf MAB algorithm on this set.
+We refer to this LLM usage as exploration oracle.
+Our usage of LLMs for exploration is closely aligned with the idea of “discretization” in the study of dynamic pricing and Lipschitz bandits (e.g.,  Kleinberg and Leighton, 2003; Kleinberg et al., 2008, 2019; Bubeck et al., 2011; Slivkins, 2014).
+Faced with a very large action space, it may be advantageous to focus on a much smaller set of candidate actions.
+A naively constructed discretization, e.g., uniform with respect to some known notion of distance in the action space, may have a poor tradeoff between size (#candidates) and quality (of the best candidate), placing all but a few candidates in low-performing regions of the action space.
+A crucial theme in this literature is “smarter” discretization algorithms which gradually “zoom in” on better-performing regions of the action space.
+The intuition is that if the better-performing arms comprise a low-dimensional region of a high-dimensional action space, covering this region should require much fewer samples for the same “discretization quality”.
+Our hope is that LLMs can suggest candidate actions from this region, if there is sufficient semantic meaning encoded in the bandit task.
+We consider two types of exploration tasks: answering an open-ended “philosophical” question (Q/A task) and suggesting a title for an arXiv research paper based on its abstract (arXiv task).
 Particular workloads within each task type are called explore puzzles.
@@
 arXiv task) is constructed as follows.
-We define the “best arm" as a contrarian answer generated by another LLM (resp.
+We define the “best arm” $a^{*}$ as a contrarian answer generated by another LLM (resp.
 the actual title of the research paper).
-The expected reward $\mu(a)$ of an arm $a$ is the cosine similarity between this arm and the best arm in the embedding space.
-[^5] While cosine similarity ranges on $[-1,1]$, it was usually strictly positive in our experiments.
+The expected reward $\mu(a)$ of an arm $a$ is the cosine similarity between $a$ and $a^{*}$ in the embedding space.
+[^8] While cosine similarity ranges on [1,1][-1,1], it was usually strictly positive in our experiments.
 In the (very rare) cases where it was negative we defined the expected reward as zero.
-Here, we generate sentence embeddings using the Sentence-BERT embedding model [35].
-[^6] Robustness check: results are largely unchanged under the Universal Sentence Encoder [11], see Appendix B.
-The realized reward in a given round is an independent Bernoulli sample with the corresponding mean defined above.
-Since action spaces are extremely large for these tasks, bandit algorithms which randomly subsample the action space catastrophically fail.
-If the embedding space, the distance notion therein, and the reward-distance relation were known to the algorithm/agent (assumptions we do not make), one could, in principle, apply “zooming algorithms" from Lipschitz bandits which gradually “zoom in" on better-performing regions of the metric space and may therefore achieve performance that scales advantageously over time [20, 21, 9, 37].
-However, these algorithms suffer an exponential dependence on the dimension $d$ of the metric space (in the multiplicative constant in front of the regret rate), making them impractical in high-dimensional spaces such as ours.
-[^7] Indeed, our embedding space has $d=384$.
-4.1 Explore Puzzle: Open-Ended Questions
+Here, we generate sentence embeddings using the Sentence-BERT embedding model Reimers (2019).
+[^9] Our results are largely unchanged under the Universal Sentence Encoder Cer (2018), see Appendix B.
+The realized reward in a given round is an independent Bernoulli sample with the mean defined above.
+Since action spaces are extremely large for these tasks, standard bandit algorithms (e.g.,  Thompson Sampling, UCB1, etc.) catastrophically fail, and so do bandit algorithms which randomly subsample the action space.
+If the embedding space, the distance notion therein, and the reward-distance relation were known to the algorithm/agent (assumptions we do not make), one could, in principle, apply the “smarter” discretization algorithms mentioned above.
+However, the regret of these algorithms scales exponentially in the dimension $d$ of the metric space making them impractical in high-dimensional action spaces such as ours.
+(Indeed, our embedding space has $d=384$.)
+Like for our exploit puzzles, we highlight our results for GPT models in the main body and include the rest in Appendix B.5. Our findings are largely consistent across different LLMs.
+Explore puzzle: open-ended questions.
 We used Gpt-4 to generate a dataset of $10$ open-ended questions with many reasonable answers, along with an intentionally contrarian answer for each question to serve as the ground truth.
 (E.g.,  “What does it mean to live a fulfilling life?” “Fulfillment comes from embracing discomfort.”) Each question-answer pair yields a task, as defined above.
-We evaluate Gpt-4o as an exploration oracle on these tasks.
-We prompt it to suggest $K\in\{1,2,3,4,5,7,10\}$ candidate answers given a question.
-To grade the candidate set (the entire set, not just the best answer in this set), we run an off-the-shelf MAB algorithm for some large-but-realistic time horizon $T$ over these $K$ candidates.
-(We use the UCB1 algorithm [5] and $T=1000$.) We record the expected time-averaged reward $\mathtt{rew}:=\tfrac{1}{T}\,\sum_{t\in[T]}\mu(a_{t})$, where $a_{t}$ is the arm selected in round $t$.
-We repeat this process (both selecting the candidates and running the bandit algorithm) 10 times for a given task and $K$, and record the average $\mathtt{rew}$ over these runs, denoted by $\overline{\mathtt{rew}}=\overline{\mathtt{rew}}(\text{task},K)$.
-[…]
-To compare against not using an LLM as an exploration oracle, we also consider a baseline ($\mathtt{Random}$) in which the candidate set consists of $K$ points selected independently and uniformly at random in the embedding space.
+We evaluate Gpt-4o, Qwen-2.5, Gemma-3, and Mistral-7B as exploration oracles.
+We prompt them to suggest $K\in\{1,2,3,4,5,7,10\}$ candidate answers given a question.
+To grade the entire candidate set (not just the best answer), we run an off-the-shelf MAB algorithm for some large-but-realistic time horizon $T$ over these $K$ candidates.
+(We use the UCB1 algorithm (Auer et al., 2002) and $T=1000$.) We record the expected time-averaged reward $\mathtt{rew}:=\tfrac{1}{T}\,\sum_{t\in[T]}\mu(a_{t})$, where $a_{t}$ is the arm selected in round $t$.
+We repeat this process (candidate selection and running UCB1) 10 times for a given task and $K$, and record the average $\mathtt{rew}$ over these runs, denoted by $\overline{\mathtt{rew}}=\overline{\mathtt{rew}}(\text{task},K)$.
+Figure 7: Arm histogram for one particular question (“What is the meaning of freedom?”, left) and averaged over all questions (right).
+We consider $K=10$ suggestions generated by our one-by-one prompt without encouragement.
+For a given “run”, the $K$ suggestions are ranked by the expected reward, and then the $i$-th expected reward is averaged over all runs, for each $i\in[K]$.
+We try several prompting strategies.
+We either ask the LLM to generate suggestions “all-at-once” with temperature 0 or “one-by-one” with temperature $1$ (we repeatedly show the LLM the list of candidate answers so far and ask it to generate one more).
+We also experiment with explicitly prompting the LLM to provide a diverse set of candidate answers (“with encouragement”).
+Thus, we have 4 prompting strategies: $\left\{\,\text{all-at-once, one-by-one}\,\right\}\times\left\{\,\text{with, without}\,\right\}$ encouragement.
+To compare against not using an LLM as an exploration oracle, we also consider a baseline ($\mathtt{Random}$) in which the candidates are $K$ points selected independently and u.a.r.
+in the embedding space.
 We visualize our findings in Figure 6.
@@
 We conclude that the LLM does succeed as an exploration oracle.
-We discuss some additional findings below.
-First, we observe that the LLM-generated suggestions pass the “eye test”, in the sense that we get reasonable, yet spiritually and semantically different answers for a given question.
-For example, given the question “What is the role of technology in society”, the first $K=5$ suggestions generated by our one-by-one prompt are as follows:
-  1. 1.
-    Facilitates communication, innovation, and efficiency.
-  2. 2.
-    Transforms daily life and shapes culture.
-  3. 3.
-    Drives connectivity and enhances productivity.
-  4. 4.
-    Facilitates control and surveillance.
-  5. 5.
-    Disrupts traditional relationships and norms.
-Second, we verify that the candidate suggestions are substantially different from one another.
-To this end, Figure 7 visualizes the spread of expected rewards within the candidate set.
+We observe that the LLM-generated suggestions pass the “eye test”, in the sense that we get reasonable, yet spiritually and semantically different answers for a given question.
+For example, given the question “What is the role of technology in society”, the first $K=5$ suggestions generated by our one-by-one prompt are as follows: (1) Facilitates communication, innovation, and efficiency, (2) Transforms daily life and shapes culture, (3) Drives connectivity and enhances productivity, (4) Facilitates control and surveillance, (5) Disrupts traditional relationships and norms.
+Finally, we verify that the candidate suggestions are substantially different from one another.
+Figure 7 visualizes the spread of expected rewards within the candidate set.
 We consider $K=10$ suggestions generated by our one-by-one prompt without encouragement.
-For a given “run" of the candidate selection, the suggestions are ranked by the expected reward, and then the $i$-th expected reward is averaged over all runs, for each $i\in[K]$.
-We consider one question (left) and the average over all questions (right).
-A more detailed comparison between the prompting strategies is perhaps not fruitful, at least within this type of experiments, because sentence embeddings are known to be somewhat imprecise at small scales (see Section B.4).
+For a given “run” of the candidate selection, the suggestions are ranked by the expected reward, and then the $i$-th expected reward is averaged over all runs, for each $i\in[K]$.
+A more detailed comparison between the prompting strategies is not fruitful for these experiments, as sentence embeddings are known to be somewhat imprecise at small scales (see Section B.4).
 Figure 8: Algorithm’s average expected reward $\overline{\mathtt{rew}}(\text{category},K)$ (averaged over rounds and tasks), against $K$, the number of candidates.
@@
 The shaded regions represent a $95\%$ confidence interval.
-A single arXiv category (“General Relativity and Quantum Cosmology", left), averages over 6 categories (right).
-4.2 Explore Puzzle: arXiv Abstracts and Titles
+A single arXiv category (“General Relativity and Quantum Cosmology”, left), averages over 6 categories (right).
+Explore puzzle: arXiv abstracts and titles.
 We run similar experiments on a larger-scale dataset of paper titles and abstracts from arXiv.org.
-Using the arXiv API [4], we collect $10$ (paper abstract, paper title) pairs from each of the $41$ different arXiv categories across various academic disciplines.
-To minimize the likelihood that these papers appear in Gpt-4o’s training corpus, we only use papers uploaded to arXiv after June 2024.
+Using the arXiv API (arXiv.org, 2025), we collect $10$ abstract-title pairs from each of the $41$ different arXiv categories.
+To minimize the likelihood that these papers appear in Gpt-4o’s training corpus, we only use papers uploaded after June 2024 (these experiments were done in Sept’24).
 Each abstract-title pair yields a task, as discussed earlier.
-We evaluate Gpt-4o as an exploration oracle for these tasks much like in in Section 4.1.
-Given an abstract, we prompt Gpt-4o to generate $K$ alternative titles, which are then used to instantiate a bandit algorithm.
+We evaluate Gpt-4o as an exploration oracle for these tasks, much like for the Q/A tasks Given an abstract, we prompt Gpt-4o to generate $K$ alternative titles, which are then used to instantiate a bandit algorithm.
 We use the same algorithm (UCB1) and time horizon $T=1000$.
-We record the expected time-averaged reward $\mathtt{rew}:=\tfrac{1}{T}\,\sum_{t\in[T]}\mu(a_{t})$ and compute the average over tasks within the same arXiv category, called $\overline{\mathtt{rew}}=\overline{\mathtt{rew}}(\text{category},K)$.
-We try “all-at-once" and “one-by-one" prompting strategies.
-[^9] We do not use “encouragement” on these tasks, since it does not appear to help much (if at all) in Section 4.1.
+We record the expected time-averaged reward, $\mathtt{rew}:=\tfrac{1}{T}\,\sum_{t\in[T]}\mu(a_{t})$, and compute the average over tasks within the same arXiv category, $\overline{\mathtt{rew}}(\text{category},K)$.
+We try “all-at-once” and “one-by-one” prompting.
+[^11] We do not use “encouragement” on these tasks, since it does not help much (if at all) for the Q/A tasks.
 We visualize our findings in Figure 8 , using the same conventions as Figure 6 and focusing on six arXiv categories.
-[^10] General relativity and quantum cosmology; computer vision and pattern recognition; statistics theory; biomolecules; signal processing; and general economics.
-For the other arXiv categories, we only collected the data for $K\in\{1,2,5\}$.
+[^12] General relativity, computer vision, statistics theory, biomolecules, signal processing, general economics.
 To assess LLM’s ability to specialize to a task, we consider a stronger baseline, where the candidate arms are generated by Gpt-4o given only the category, not the abstract.
-Category $K=1$ $K=2$ $K=5$
-gr-qc 0.63 0.64 0.65
-hep-ex 0.78 0.76 0.81
-hep-lat 0.72 0.72 0.74
-hep-ph 0.7 0.72 0.73
-hep-th 0.64 0.69 0.68
-math-ph 0.65 0.71 0.71
-Table 1: Results for other arXiv categories: per-category rewards $\overline{\mathtt{rew}}(\text{category},K)$ for one-by-one prompting and $K\in\{1,2,5\}$.
-Results for the other arXiv categories are provided via a per-category table with $\overline{\mathtt{rew}}(\text{category},K)$ for $K\in\{1,2,5\}$ and both prompts (Appendix B).
-In Table 1, we show them for (other) six arXiv categories and one-by-one prompting.
-Similar to Section 4.1, we conclude that (1) both prompting strategies significantly outperform the baseline, and (2) their performance tends to improve modestly as $K$ increases.
+For the other arXiv categories, we provide a table with $\overline{\mathtt{rew}}(\text{category},K)$ for $K=1,2,5$ and both prompts (Appendix B).
+Like for the Q/A tasks, (1) both prompting strategies significantly outperform the baseline, and (2) their performance tends to improve modestly as $K$ increases.
 4 Conclusions
-We evaluate the potential of Gpt-4, Gpt-4o, and Gpt-3.5 to explore and exploit in decision-making tasks.
-We find that LLMs are useful as exploration oracles in large, semantically meaningful action spaces, where they effectively propose high-quality candidate actions.
-However, LLMs cannot yet robustly replace (let alone surpass) traditional algorithmic methods for exploitation, particularly in larger or more complex settings.
-While we suggest several helpful mitigations, they consistently underperform relative to a simple algorithmic baseline such as linear regression.
-We highlight two directions for future work.
-First, LLMs trained to use tools like a calculator may be better at exploitation.
-However, it is unclear how much this would help in more complex scenarios, e.g.,  CB tasks with text-based contexts and actions, and which mitigations and/or prompting techniques would be needed.
-Second, while "zooming" bandit algorithms do not work for rich text-based action spaces (as discussed in Section 4), LLM-based exploration oracles may potentially help.
-The hope is to “zoom in" entirely in the space of “potentially relevant" actions (determined by the LLM), rather than in the space of all actions.
+Our work adds to (and provides partial explanations for) the growing literature on in-context learning for decision-making tasks.
+We find that LLMs are useful as exploration oracles that propose high-quality candidate actions in large, semantically meaningful action spaces.
+However, current LLMs are not that good at exploitation, particularly in larger or more complex tasks (we trace the “boundary of tractability” in terms of task size/complexity).
+While we suggest several helpful mitigations, they consistently underperform relative to a simple linear regression, even in inherently non-linear tasks.
+Future work.
+LLMs trained to use tools like a calculator may be better at exploitation.
+However, it is unclear how much this would help in more complex scenarios, e.g.,  CB tasks with text-based contexts and actions, and which mitigations/prompting techniques would be needed.
+Second, while ”zooming” bandit algorithms do not work for rich text-based action spaces (discussed in Section 3), LLM-based exploration oracles may potentially help.
+The hope is to “zoom in” entirely in the space of “potentially relevant” actions (as determined by the LLM), rather than in the space of all actions.
 Appendix A Appendix for Section 2: LLMs as exploitation oracles
@@
 [USER] The past rewards for each button are:
-round $1$: blue button had reward $1$, green button had reward $1$, red button had reward 00, yellow button had reward $1$, purple button had reward 00
-round $2$: blue button had reward 00, green button had reward $1$, red button had reward $1$, yellow button had reward $1$, purple button had reward 00
+round $1$: blue button had reward $1$, green button had reward $1$, red button had reward 0, yellow button had reward $1$, purple button had reward 0
+round $2$: blue button had reward 0, green button had reward $1$, red button had reward $1$, yellow button had reward $1$, purple button had reward 0
 Which button do you choose? Remember, YOU MUST provide your final answer within the tags <Answer>COLOR <Answer>where COLOR is one of blue, green, red, yellow, or purple and with no text explanation.
@@
 A new list of numbers will then appear on the screen and you must pick the next button in order to maximize your reward in this round only, which must be exactly one of blue or green.
-You must provide your final answer immediately within the tags <Answer> COLOR </Answer> where COLOR is one of blue or green and with no text explanation.
+You must provide your final answer immediately within the tags ¡Answer¿ COLOR ¡/Answer¿ where COLOR is one of blue or green and with no text explanation.
 [USER] The past contexts and rewards for each button are:
-In round $1$, the context was $[0.3,0.7]$. The blue button had reward $1$, the green button had reward $1$
-In round $2$, the context was $[0.4,0.6]$. The blue button had reward 00, the green button had reward $1$
+In round $1$, the context was $[0.3,0.7]$. The blue button had reward $1$, the green button had reward 11
+In round $2$, the context was $[0.4,0.6]$. The blue button had reward 0, the green button had reward 11
 Which button do you choose? Remember, YOU MUST provide your final answer within the tags <Answer>COLOR <Answer>where COLOR is one of blue or green and with no text explanation.
@@
 There may also be other objects in the room, which I will tell you about.
-You must then take one of the following actions: "pet animal", "leave room", "use tool", "eat food", "press button", after which you will receive some reward.
+You must then take one of the following actions: ”pet animal”, ”leave room”, ”use tool”, ”eat food”, ”press button”, after which you will receive some reward.
 The reward you receive is a random function of both the action you take and the information you receive about the objects in the room and time of day.
@@
 For each round, I will show you the state of the room and the corresponding rewards for each action.
-I will then tell you the current state of the room, and you must pick the next action in order to maximize your reward in this round only, which must be exactly one of "pet animal", "leave room", "use tool", "eat food", or "press button".
+I will then tell you the current state of the room, and you must pick the next action in order to maximize your reward in this round only, which must be exactly one of ”pet animal”, ”leave room”, ”use tool”, ”eat food”, or ”press button”.
 Look for patterns in the data and try to estimate the reward of each action, given the information at your disposal.
-You must provide your final answer immediately within the tags <Answer>ACTION <Answer>where ACTION is one of "pet animal", "leave room", "use tool", "eat food", or "press button" and with no text explanation.
+You must provide your final answer immediately within the tags <Answer>ACTION <Answer>where ACTION is one of ”pet animal”, ”leave room”, ”use tool”, ”eat food”, or ”press button” and with no text explanation.
 [USER] The past observations and outcomes for each action are:
 Round $1$ had context time of day: morning, animal: bear, table item: chest, tool: key, food: apple, button color: red.
-"pet animal" had reward 00, "leave room" had reward $1$, "use tool" had reward $1$, "eat food" had reward 00, press button had reward 00
+”pet animal” had reward 0, ”leave room” had reward $1$, ”use tool” had reward $1$, ”eat food” had reward 0, press button had reward 0
 Round $2$ had context time of day: afternoon, animal: cat, table item: card, tool: hammer, food: cake, button color: orange.
-"pet animal" had reward $1$, "leave room" had reward 00, "use tool" had reward 00, "eat food" had reward $1$, press button had reward 00
+”pet animal” had reward $1$, ”leave room” had reward 0, ”use tool” had reward 0, ”eat food” had reward $1$, press button had reward 0
 The current state of the room is time of day: evening, animal: bear, table item: envelope, tool: key, food: nut, button color: red.
 Which action do you choose?
-Remember, you must provide your final answer immediately within the tags <Answer>ACTION <Answer>where ACTION is one of "pet animal", "leave room", "use tool", "eat food", or "press button" and with no text explanation.
+Remember, you must provide your final answer immediately within the tags <Answer>ACTION <Answer>where ACTION is one of ”pet animal”, ”leave room”, ”use tool”, ”eat food”, or ”press button” and with no text explanation.
 A.2 Additional MAB figures
@@
 In all other cases, the expected reward is 0.25.
+A.4 Robustness Check with Other Models
+In this appendix, we include results for Qwen2.5-7B-Instruct (Qwen-2.5), Gemma 3 12B (Gemma-3), and Mistral-7B-Instruct-v0.3 (Mistral-7B) on our exploit puzzles.
+Due to the time and costs associated with running our exploit puzzles, we prioritized (1) getting at least one additional set of results for each of our main puzzles (using Qwen-2.5) and (2) getting results for all of our models in our word-based puzzles.
+Our findings are the same across all models.
+Figure 13: Results for Qwen-2.5 in the setting of Figure 4. Only mitigations are shown, as the unmitigated history is too long to fit in the context window.
+Figure 14: Results for Mistral-7B in the setting of Figure 4. Only mitigations are shown, as the unmitigated history is too long to fit in the context window.
+Figure 15: Results for Qwen-2.5 in the setting of Figure 3. Only mitigations are shown, as the unmitigated history is too long to fit in the context window.
+Figure 16: Results for Qwen-2.5 in the setting of Figure 5 (right). We use a slightly smaller history length ($T=450$ instead of $T=500$) so that the full history can fit in the context window.
+Figure 17: Results for Mistral-7B in the setting of Figure 16. Only mitigations are shown, as the full history could not fit in the context window.
+Figure 18: Results for Gemma-3 in the setting of Figure 16. Only mitigations are shown, as the full history could not fit in the context window.
+Figure 19: Results for Qwen-2.5 in the setting of Figure 12. We use a slightly smaller history length ($T=450$ instead of $T=500$) so that the full history can fit in the context window.
+Figure 20: Results for Qwen-2.5 in the setting of Figure 12, using chain-of-thought prompting.
+We use a slightly smaller history length ($T=450$ instead of $T=500$) so that the full history can fit in the context window.
+Figure 21: Results for Mistral-7B in the setting of Figure 19. Only mitigations are shown, as the full history could not fit in the context window.
+Figure 22: Results for Gemma-3 in the setting of Figure 19. Only mitigations are shown, as the full history could not fit in the context window.
 Appendix B Appendix for Section 3: LLMs as exploration oracles
@@
   1. 1.
-    The evidence of $N=16$ shell closure and $\beta$-delayed neutron emission from $\wedge^{25}$
+    The evidence of $N=16$ shell closure and $\beta$-delayed neutron emission from $\wedge^{25}$F
   2. 2.
@@
   7. 7.
-    Popa’s "Recurrent Sequences" and Reciprocity
+    Popa’s ”Recurrent Sequences” and Reciprocity
   8. 8.
@@
 The similarity scores of both models in Table 9 suggest that while the embeddings produced by both embedding models are generally “in the ballpark” of what one would consider “similar”/“different”, they are still a somewhat coarse measure of distance, which may explain the similar performance of our different prompting strategies.
+B.5 Robustness Check with Other Models
+In this appendix, we include results for Qwen2.5-7B-Instruct (Qwen-2.5), Gemma 3 12B (Gemma-3), and Mistral-7B-Instruct-v0.3 (Mistral-7B) on our explore Q/A puzzles.
+As was the case with the GPT models, performance increased with $K$ when the candidate answers are generated one-by-one.
+When generating answers all-at-once, we observed that Qwen-2.5 and Mistral-7B had trouble generating concise answers, leading to a drop in performance as $K$ increases.
+Figure 37: Results for Qwen-2.5 in the setting of Figure 6.
+Figure 38: Results for Mistral-7B in the setting of Figure 6.
+Figure 39: Results for Gemma-3 in the setting of Figure 6.

### v3

@@
 We evaluate the ability of the current generation of large language models (LLMs) to help a decision-making agent facing an exploration-exploitation tradeoff.
-We use LLMs to explore and exploit in silos in various (contextual) bandit tasks.
-We find that while the current LLMs often struggle to exploit, in-context mitigations may be used to substantially improve performance for small-scale tasks.
-However even then, LLMs perform worse than a simple linear regression.
+While previous work has largely study the ability of LLMs to solve combined exploration-exploitation tasks, we take a more systematic approach and use LLMs to explore and exploit in silos in various (contextual) bandit tasks.
+We find that reasoning models show the most promise for solving exploitation tasks, although they are still too expensive or too slow to be used in many practical settings.
+Motivated by this, we study tool use and in-context summarization using non-reasoning models.
+We find that these mitigations may be used to substantially improve performance on medium-difficulty tasks, however even then, all LLMs we study perform worse than a simple linear regression, even in non-linear settings.
 On the other hand, we find that LLMs do help at exploring large action spaces with inherent semantics, by suggesting suitable candidates to explore.
 1 Introduction
-There’s been significant interest in the machine learning community to apply recent advances in generative AI and large language models (LLMs) to solve important decision-making problems.
+The machine learning community is increasingly interested to apply advances in generative AI and large language models (LLMs) to decision-making problems.
 Early work in this direction has already produced impressive agentic behavior in both virtual (e.g.,  Wang et al., 2023; OpenAI, 2025) and physical-world environments (e.g.,  Black et al., 2024).
-Beyond generalization (needed for supervised learning), decision-making under uncertainty requires two additional capabilities: exploitation (making the best decision given the current data) and exploration (trying new options for long-term benefit).
-Balancing the two has led to an enormous literature (e.g.,  see books Slivkins, 2019; Lattimore and Szepesvári, 2020; Agarwal et al., 2020).
+Beyond generalization (needed for supervised learning), decision-making under uncertainty requires two additional capabilities: exploitation (making best decisions given the current data) and exploration (for long-term benefit).
+Balancing the two has led to a large literature (e.g.,  Slivkins, 2019; Lattimore and Szepesvári, 2020; Agarwal et al., 2020).
 A recent line of work (e.g.,  Krishnamurthy et al., 2024; Nie et al., 2024) evaluates the ability of LLMs to balance exploration and exploitation entirely in-context, i.e.,  specifying the problem description, parameters, and history in the LLM prompt.
 Focused on simple tasks in reinforcement learning (RL), these results are mixed.
-Both papers show that LLMs fail to solve these tasks adequately out-of-the-box, but they can be prompted to do so by providing succinct summary statistics in-context.
-However, such statistics do not exist beyond simple decision-making problems, e.g.,  for contextual bandits.
+LLMs fail to solve these tasks adequately out-of-the-box, but they can be prompted to do so by providing succinct summary statistics in-context.
+However, such statistics do not exist beyond simple RL problems, e.g.,  for contextual bandits.
 RL-specific pre-training or fine-tuning (on data from algorithmic baselines on similar problem instances) tends to work well (see Related Work), but may be prohibitive due to cost or insufficient training data.
-Besides, using a commonly available frontier LLM would leverage its “generalist” intelligence and may be much easier in terms of logistics and required expertise.
+Besides, using a commonly available LLM would leverage its “generalist” intelligence and may be much easier in terms of logistics and required expertise.
 Motivated by these observations, we study the ability of LLMs to explore and exploit in-context in silos, with an eye towards leveraging a pre-trained LLM (and the inductive bias therein) as a part of a larger decision-making agent.
 We focus on (contextual) bandits, as a standard abstraction for the explore-exploit tradeoff.
-We experiment with a variety of LLMs: Gpt-4, Gpt-4o, Gpt-3.5, Qwen-2.5, Gemma-3, Mistral-7B, as well as DeepSeek-R1, a reasoning model.
-[^2] Several other LLMs did not work for our purposes: Llama-3.2-3B and Llama-3-8B would not follow instructions, as well as DeepSeek-R1-Distill-Llama-70B (a reasoning model); Qwen-2.5-Math-70B (another reasoning model that we tried) has a short context window.
-We restrict our GPT access to Gpt-4, Gpt-4o, and Gpt-3.5, as newer variants such as Gpt-4.1 and Gpt-4.5 are primarily optimized for efficiency (e.g.
-latency, cost, coding) rather than broad general reasoning, and Gpt-5 uses a reasoning backbone as a routing mechanism across sub-models.
+We experiment with many LLMs: Gpt-5.2, Gpt-5-mini, Gpt-5-nano, Gpt-4, Gpt-4o, Gpt-3.5, Qwen-2.5, Gemma-3, Mistral-7B, and DeepSeek-R1, a reasoning model.
+[^2] Some other LLMs did not work for our purposes: Llama-3.2-3B would not follow instructions.
+Two reasoning models, DeepSeek-R1-Distill-Llama-70B and Qwen-2.5-Math-70B have a context window that is too short.
 In Section 2, we evaluate LLMs as exploitation oracles for contextual bandits.
@@
 Our results here are mixed.
-We show that LLMs can effectively exploit in-context for small-sized problems, but their performance degrades when the problem becomes moderately sized.
-We find that in-context summary techniques are useful for improving performance, but LLMs with these mitigations still perform worse than a simple linear regression baseline.
-In Section 3, we evaluate LLMs as exploration oracles which suggest candidate actions by discretizing a large action space.
-To do so, we introduce a text-based multi-armed bandit task, where actions correspond to free-text answers to an open-ended question, and rewards are driven by the distance from some preselected answer in an embedding space (the latter is computed exogenously).
-Given the high dimensionality of the action space, traditional discretization approaches typically used in continuous bandit problems are inapplicable.
-However, an LLM can generate a small set of candidate actions which can then be used to instantiate an off-the-shelf bandit algorithm.
-We experiment with several prompting strategies, and find that they all lead to relatively good exploration.
-Finally, we repeat our experiments on a larger-scale bandit task based on paper titles and abstracts from arXiv (where the goal is to find a suitable title for a given abstract), with similar findings.
+We show that LLMs can effectively exploit in-context for small-sized tasks but their performance degrades when the tasks become moderately sized.
+We find that frontier reasoning models outperform non-reasoning models (fixing model size and provider), but tend to be slow and/or expensive.
+We investigate to mitigations: in-context summary techniques and tool use (namely, a Python code interpreter).
+We find that both mitigations improve performance in a linear task, but on on-linear tasks LLMs with these mitigations still perform worse than a simple linear regression baseline.
+In Section 3, we evaluate LLMs as an exploration oracle which suggests a small set of candidate actions by discretizing a large action space.
+(This set can then be used to instantiate an off-the-shelf bandit algorithm.) On various text-based multi-armed bandit tasks that we design, we find this approach far superior to a non-LLM baseline.
+Our first task is based on the MovieLens dataset (Harper and Konstan, 2015), where actions are movie recommendations, rewards are random draws parameterized by ground-truth movie ratings, and the baseline is a hand-crafted discretization based on movie genre.
+Second, we design a larger-scale bandit task based on paper titles/abstracts from arXiv: find a suitable title for a given abstract.
+Given the high dimensionality of the action space, discretization approaches from continuous bandit problems are inapplicable.
+Several prompting strategies all lead to relatively good exploration compared to natural baselines.
+Our third task is based on open-ended “philosophical” questions and contrarian answers.
 Related work.
@@
 While our paper is primarily concerned with whether LLMs succeed as algorithms, several others (e.g.,  Schubert et al., 2024; Hayes et al., 2024; Coda-Forno et al., 2024) use in-context bandits (and many other tasks) to study whether LLMs exhibit human-like behavior/biases in decision-making.
-A broader literature on in-context learning (starting from Brown et al., 2020) aims to solve various tasks by providing all relevant information in the LLM prompt.
+A broader literature on in-context learning (starting from Brown et al.
+(2020)) aims to solve various tasks by providing all relevant information in the LLM prompt.
 The work on exemplar selection (selecting examples and other information to present in-context) (e.g.,  Khalifa et al., 2023; Zhang et al., 2022; Xiong et al., 2023; Tonglet et al., 2023) is relevant to our exploitation experiments.
@@
 Background on multi-armed bandits (MAB).
-We consider tasks based on MAB and contextual bandits (CB), well-studied special cases of RL that abstract the explore-exploit tradeoff, see Slivkins (2019); Lattimore and Szepesvári (2020) for background.
+We consider tasks based on MAB and contextual bandits (CB), well-studied special cases of RL that abstract explore-exploit tradeoff (Slivkins, 2019; Lattimore and Szepesvári, 2020).
 In MAB, there are $T$ rounds and $K$ arms.
-In each round $t\in[T]$, the learner chooses an action (arm) $a_{t}\in[K]$ and observes reward $r_{t}$ drawn from some sub-Gaussian reward distribution for this arm.
-The reward distribution, and particularly its mean $\mu(a_{t})$, are unknown to the algorithm.
-In CB, the learner additionally observes a context $z_{t}$ before each round $t$, and the expected reward $\mu(z_{t},a_{t})$ depends on both the context and the arm.
+In each round $t\in[T]$, the learner chooses an action (arm) $a_{t}\in[K]$ and observes reward $r_{t}$ drawn from some (fixed, unknown) sub-Gaussian reward distribution $D(a_{t})$ with mean $\mu(a_{t})$.
+In CB, the learner also observes a context $z_{t}$ before each round $t$.
+The expected reward $\mu(z_{t},a_{t})$ depends on both the context and the arm.
 The learner’s goal is to balance exploration and exploitation to maximize cumulative reward.
@@
 Designing CB exploitation oracles for their own sake, a.k.a.
-offline policy optimization, is well-studied (starting from, e.g.,  Beygelzimer and Langford, 2009; Dudík et al., 2014).
+offline policy optimization, is well-studied (starting from, e.g.,  Beygelzimer and Langford (2009); Dudík et al.
+(2014)).
 Usage of an LLM as an “exploration oracle” is closely related to the literature on dynamic pricing and Lipschitz bandits (e.g.,  Kleinberg and Leighton, 2003; Kleinberg et al., 2008, 2019; Bubeck et al., 2011), as we elaborate in Section 3.
 2 LLMs as exploitation oracles
-Figure 1: MAB exploit puzzle for Gpt-4 (left), Gpt-4 with CoT (middle), and Gpt-3.5 with CoT (right), all with ”buttons” prompt.
+Figure 1: MAB exploit puzzle for Gpt-3.5 (left), Gpt-4 (middle), and Gpt-5-nano (right), all with the “buttons” prompt and temperature 11.
+Results for Gpt-5-mini are not visualized, as $\mathrm{FracCorrect}(\epsilon,T)=1$ for all $\epsilon$ and $T$.
 The following conventions apply to all figures in this section.
@@
 These tasks are generated from some parameterized distributions called exploit puzzles.
-As our results here are largely negative, we focus on relatively small tasks.
-We use synthetic data, to better control task size/difficulty and trace the “boundary of tractability”.
-We find that LLMs are consistently outperformed by a simple linear regression baseline, even on moderately-sized tasks.
-Due to computational constraints, our experiments focused on Gpt-4, Gpt-4o, and Gpt-3.5, with additional LLMs evaluated as a robustness check in Appendix A.4, with similar findings.
+Due to computational constraints, our experiments focus mostly on Gpt model family, with additional LLMs evaluated as a robustness check in Appendix B.4, with similar findings.
+Across all experiments, varying the model temperature did not significantly affect performance.
+We include results for temperature 0 and 1.
 MAB exploit puzzles.
-Our MAB-based experiments on Gpt-4 and Gpt-3.5 provide a partial explanation for why the current LLMs fail to solve MAB tasks in-context when presented with raw (non-summarized) history, as first observed by Krishnamurthy et al.
-(2024); Nie et al.
-(2024).
-Following these two papers, we try two prompts: one in which arms correspond to pushing different colored buttons and one where they correspond to showing different advertisements to users.
+We begin by testing the Gpt model family on simple MAB exploit puzzles.
+We find that performance significantly improves as the models progress from Gpt-3.5 to Gpt-4 to Gpt-5.
+In fact, Gpt-5-mini achieves perfect performance.
+Our experiments on Gpt-4 and Gpt-3.5 provide a partial explanation for the observation in (Krishnamurthy et al., 2024; Nie et al., 2024), that these models fail to solve end-to-end MAB tasks in-context when presented with raw (non-summarized) history.
+Following these two papers, we try two prompts, where arms correspond, resp., to pushing different colored buttons and to showing different ads to users.
 The LLM is asked to choose the arm with the highest empirical reward in the next round.
 We also try chain-of-thought (CoT) prompts, for the total of 4 prompt designs: $\left\{\,\text{buttons, adverts}\,\right\}\times\left\{\,\text{CoT, no-CoT}\,\right\}$.
-See Appendix A for more details on our experimental setup.
+See Appendix B.1 for more details on this setup.
 […]
+Figure 2: Gpt-5 (left) and Gpt-4 (right) model families on a numerical CB puzzle (all with temperature 11).
+Gpt-5 models generally perform better than Gpt-4 models of similar size, although letting Gpt-4 and Gpt-4o use a code interpreter improves performance.
 Given a (realized) exploit task, we measure its difficulty via empirical gap $\epsilon$: the difference between the largest and second-largest average reward $\bar{r}_{t}(a):=\frac{1}{T}\sum_{t\in[T]}r_{t}(a)$ among all arms $a$.
-Intuitively, puzzle difficulty decreases with $\epsilon$ (as one needs to distinguish the top two arms).
+Intuitively, puzzle difficulty decreases with $\epsilon$ (as it gets harder to distinguish the top two arms).
 Empirical gap mirrors the “gap” between the top two expected rewards, a standard notion of difficulty in MAB.
-Let $S(\epsilon,T)$ be the set of all tasks with empirical gap at most $\epsilon$ and history size $T$.
 We measure an LLM’s performance over a given set $S$ of tasks as the fraction of tasks for which the LLM returns a “correct answer”: an arm with the largest empirical reward; denoted by $\mathrm{FracCorrect}(S)$.
 We are interested in how $\mathrm{FracCorrect}$ varies depending on the difficulty level.
-Hence, we plot $\mathrm{FracCorrect}(\epsilon,T):=\mathrm{FracCorrect}\left(\,S(\epsilon,T)\,\right)$ against empirical gap $\epsilon$.
-We find that Gpt-4 and Gpt-3.5 do not perform well on these MAB exploit puzzles, see Figure 1.
-Performance tends to degrade (1) as the history length $T$ increases and (2) as the empirical gap decreases.
-(Our intuition for the former is that the LLM has more data points to process in context, and for the latter is that the bandit instance becomes more difficult.) While Gpt-4 generally performs much better than Gpt-3.5, we found that prompting the LLM to use chain-of-thought (CoT) reasoning provided a slight boost for Gpt-3.5, while hurting the performance of Gpt-4.333From CoT logs, we found that Gpt-4 was doing incorrect reasoning in a non-trivial fraction of runs.
-We found that performance was similar across all of our prompt designs (see Appendix A for more plots).
-Figure 2: Gpt-4 succeeds on a small CB exploit puzzle (left), but fails on a slightly larger one (right).
-CB exploit puzzles and mitigations.
-While the history in $K$-armed bandits can be summarized using $2K$ numbers (for each arm, the average reward and #plays) such succinct summary statistics may not be readily available (or even exist) in more complicated decision-making tasks such as CB.
+Hence, let $S(\epsilon,T)$ be the set of all tasks with empirical gap at most $\epsilon$ and history size $T$.
+We plot $\mathrm{FracCorrect}(\epsilon,T):=\mathrm{FracCorrect}\left(\,S(\epsilon,T)\,\right)$ against $\epsilon$.
+We find that while Gpt-4 and Gpt-3.5 do not perform well, Gpt-5 models do, see Figure 1.
+For Gpt-4 and Gpt-3.5, performance tends to degrade as (1) history size $T$ increases, with more data points to process in-context and (2) empirical gap decreases, making the bandit instance more difficult.
+CB exploit puzzles.
+While the history in $K$-armed bandits can be summarized with $2K$ numbers (for each arm, the average reward and #plays), such succinct summary statistics may not be readily available (or even exist) in more complicated decision-making tasks such as CB.
 [^3] This consideration also motivates MAB exploit puzzles with raw (non-summarized) history, as a simpler special case of the general scenario when succinct summarization is unavailable.
-We focus on linear CB, where the expected reward of each arm $a$ is linear in the context $z\in\mathbb{R}^{d}$: $\mu(z,a)=\langle z,\theta^{*}_{a}\rangle$ for some fixed (but unknown) parameters $\theta^{*}_{a}\in\mathbb{R}^{d}$.
+We begin with linear CB, where the expected reward of each arm $a$ is linear in the context $z\in\mathbb{R}^{d}$: $\mu(z,a)=\langle z,\theta^{*}_{a}\rangle$ for some fixed (but unknown) vector $\theta^{*}_{a}\in\mathbb{R}^{d}$.
 [^4] Linear CB are well-studied, starting from Li et al.
@@
 (2011).
-(We try non-linear CB, too.)
+(We turn to non-linear CB later in the section.)
 We consider a CB exploit puzzle parameterized by #arms $K$, dimension $d$, and history size $T$.
 The tasks are constructed as follows.
-We sample parameters $\theta_{a}\in[-1,1]^{d}$ and $\gamma_{a}\in[-0.25,0.25]$ independently and u.a.r.
+We sample parameters $\theta_{a}\in[-1,1]^{d}$ and $\gamma_{a}\in[-\frac{1}{4},\frac{1}{4}]$ independently u.a.r.
 for each arm $a$.
@@
 We generate a history of $T$ rounds $t\in[T]$.
-Contexts $z_{t}$ are sampled independently and u.a.r.
+Contexts $z_{t}$ are sampled independently u.a.r.
 from $1,1]^{d}$.
-For simplicity, the history contains rewards of all arms $a$ in each round $t$, where the reward equals $\mu(z_{t},a)$ plus an independent unit-variance Gaussian.
+For simplicity, the history contains rewards of all arms $a$ in all rounds $t$, where the reward equals $\mu(z_{t},a)$ plus an independent unit-variance Gaussian.
 Given the history and a new context $z_{T+1}$ (drawn in the same way), the LLM is asked to select the action for round $T+1$ which appears best.
@@
 The prompt does not mention linearity of the CB instance (because such model-based information is typically unavailable in applications).
-Given an exploit task, we now define a “correct answer” as an arm $a$ which maximizes expected reward $\mu(z_{T+1},a)$.
+Given an exploit task, we now define a “correct answer” as arm $a$ which maximizes expected reward $\mu(z_{T+1},a)$.
 [^5] Note that it is unclear how to define an ”empirically best arm” given a CB history and the current context.
 $\mathrm{FracCorrect}(S)$ is the fraction of correct answers in a given set $S$ of tasks.
-Likewise, the task difficulty is not easily defined in terms of the realized rewards.
-Instead, we focus on the effective gap: the difference in expected reward between the best and second-best arm given the current context $z=z_{T+1}$.
-That is, the difference between the largest and second-largest number among $\mu(a,z)$, $a\in[K]$.
-Intuitively, smaller effective gap corresponds to increased difficulty.
-We are interested in how $\mathrm{FracCorrect}$ varies with effective gap.
-In each plot, we fix the number of arms ($K$) and dimension ($d$), and let $S(\epsilon,T)$ be the set of all tasks with given $K,d,T$ and effective gap at most $\epsilon$.
+Task difficulty is also not easily defined in terms of realized rewards.
+Instead, we focus on effective gap $\epsilon$: the difference in expected reward $\mu(a,z)$ between the best and second-best arm given the current context $z=z_{T+1}$.
+Intuitively, smaller $\epsilon$ corresponds to increased difficulty.
+We study how $\mathrm{FracCorrect}$ varies with effective gap.
+In each plot, we let $S(\epsilon,T)$ be the set of all tasks with given $K,d,T$ and effective gap at most $\epsilon$.
 We plot $\mathrm{FracCorrect}(\epsilon,T):=\mathrm{FracCorrect}\left(\,S(\epsilon,T)\,\right)$ against $\epsilon$.
-First, we find that Gpt-4 obtains near-perfect performance on a “small” puzzle with $K=2$ arms, context dimension $d=1$, and history size $T=50$ (Figure 2).
-However, its performance degrades as the problem size increases: see Figure 2 for $K=d=2$ and $T=100$.
+Figure 3: CB exploit puzzle with $d=K=2$ and $T=4000$: mitigations help substantially.
+Gpt-4 without CoT (left) and Gpt-4 with CoT (right).
+(Full history with this $T$ vastly exceeds the context window for Gpt-4, Gpt-4o, and Gpt-3.5.)
+Figure 2 considers a setting with $d=K=2$ for the Gpt-5 family (left) and the Gpt-4 family (right).
+We find that larger models generally perform better than their smaller, distilled counterparts and, conditioned on model size, Gpt-5 models outperform Gpt-4 models.
+We also plot a linear regression baseline (blue dashed line), which effectively serves as an upper-bound on the performance of any model, due to the underlying linear structure of our CB puzzle.
+While not pictured, all models tend to perform strictly better on smaller problem instances.
+For example, Gpt-4 performed near-perfectly when $d=1$, $K=2$, and $T=50$.
+However, performance degrades as the problem size grows (e.g., see Gpt-4o in Figure 4).
 Moreover, limited prompt size may prevent processing larger histories.
 [^6] E.g.,  our LLM access points bottomed out at $T\approx\text{100-200}$ for Gpt-4 and $T\approx\text{1000-2000}$ for Gpt-4o.
-Figure 3: CB exploit puzzle with $d=K=2$ and $T=4000$: mitigations help substantially.
-Gpt-4 without CoT (left) and Gpt-4 with CoT (right).
-Note that providing the full history with this $T$ vastly exceeds the context window for Gpt-4, Gpt-4o, and Gpt-3.5.
-Motivated by these observations, we implement several natural mitigations inspired by the literature on exemplar selection for in-context learning (discussed in Related Work).
-1.
-$k$-nearest: Among the observed contexts, consider the distinct $k$ contexts closest to $z_{T+1}$, according to the $\ell_{2}$ metric.
-Limit the history reported in the prompt to (the rounds with) these $k$ contexts.
+Motivated by these observations, we implement two types of mitigations: tool use (providing the model with access to a Python interpreter to execute code in) and in-context summarization (inspired by the literature on exemplar selection for in-context learning, see Related Work).
+Evaluating these mitigations, we focus on the Gpt-4 family as (1) at the time of writing Gpt-5 models are not configured for tool use and (2) we found the Gpt-5 models to be either prohibitively expensive (Gpt-5.2) or prohibitively slow (Gpt-5-mini and Gpt-5-nano) on larger-scale experiments.
+See Figure 2 (right) for the performance of Gpt-4 and Gpt-4o with access to a code interpreter (CI).
+Both see significantly improved performance with CI, with Gpt-4 even surpassing Gpt-5.2.
+However, enabling the CI made Gpt-4 and Gpt-4o much slower and much more expensive.
+[^7] E.g., our Gpt-4 experiment with CI in Figure 2 cost approximately $\mathdollar 70$ and took 6+ hours to run.
+Our Gpt-4 experiment in Figure 2 without CI cost a few dollars and took under 5 minutes.
+We consider the following summarization mitigations:
+1. $k$-nearest: Among the observed contexts, pick distinct $k$ contexts closest to $z_{T+1}$, according to the $\ell_{2}$ metric. Limit the reported history to (the rounds with) these $k$ contexts.
 2.
@@
 Report $\left(\,z^{*},\,a,\,\bar{r}(z^{*},a)\,\right)$ as a context-arm-reward triple, for each arm $a$ and each centeroid context $z^{*}$ among the $k^{\prime}$ centroids closest to $z_{T+1}$ (according to the $\ell_{2}$ metric).
-We do not explain the “mitigation strategy” in the prompt: we present the reported context-arm-reward tuples as if it were the entire history, without mentioning clustering or averaging.
-Figure 3 visualizes the performance of these mitigations on a slightly larger (but still relatively small) puzzle with $K=d=2$ and $T=4000$.
-We use Gpt-4 with and without CoT.
-We compare the mitigations against linear regression baseline (which is effectively an ”upper bound”, as the underlying CB instance is linear).
-Without CoT prompting, we find that all three mitigations achieve $\mathrm{FracCorrect}$ around $80\%-90\%$, although this dips to around $60\%-85\%$ when using CoT.
-In addition to (potentially) improving performance, mitigations can also offer a practical way to solve decision-making tasks using LLMs when the history is large; when $T=4,000$, our prompt vastly exceeds the context window of all models we had access to.
-Figure 4: CB exploit puzzle with $d=K=5$ and $T=1000$: mitigations perform badly, but (mostly) much better than the no-mitigation baseline. Gpt-4o without CoT.
-However, current LLMs struggle to exploit on even moderately-sized problems, even with these mitigations.
-In Figure 4, we plot the performance of Gpt-4o with mitigations under various hyperparameters.
-While $k$-means (left) is almost as bad as random guessing, the $k$-nearest mitigations (center) achieve about $50\%$ $\mathrm{FracCorrect}$, and the $k$-means, $k^{\prime}$-nearest mitigations (right) obtain approximately $60\%$ $\mathrm{FracCorrect}$.
-Both our $k$-nearest and $k$-means, $k^{\prime}$-nearest mitigations significantly out-perform unmitigated Gpt-4o, but fall significantly short of the linear baseline.
-Figure 5: Left: Performance of Deepseek-R1 on our numerical CB puzzle. Right: Gpt-4o on the text-based CB exploit puzzle. Some mitigations help, but are outperformed by linear regression.
-Reasoning models are LLMs specifically trained to follow multi-step chains of thought to solve complex problems that require intermediate reasoning.
-We ran one such model, DeepSeek-R1-Distill-Qwen-32B (henceforth DeepSeek-R1), a 32B parameter reasoning model distilled from Qwen-2.5.
-Our results are shown in Figure 5 (left).
-As before, we find that DeepSeek-R1 fails to match the performance of the baseline, even when using the mitigations.
-High inference costs and small context windows prevented us from running larger reasoning models.
+We do not explain the “mitigation strategy” in the prompt: we present the context-arm-reward tuples as if it were the entire history without mentioning clustering or averaging.
+Figure 3 shows the performance of our mitigations on a slightly larger puzzle with $K=d=2$, $T=4000$.
+We use Gpt-4 with and without chain-of-thought (CoT) prompting.
+As before, we compare against a linear regression baseline.
+Without CoT, all three mitigations achieve $\mathrm{FracCorrect}$ $\approx 80\%-90\%$, although this dips to $\approx 60\%-85\%$ with CoT.
+Performance aside, the mitigations offer a practical way to handle large histories; e.g., when $T=4,000$, our prompt exceeds the context window of all models we had access to.
+Figure 4: CB exploit puzzle with $d=K=5$ and $T=1000$: in-context mitigations perform badly, but (mostly) much better than the no-mitigation baseline. Gpt-4o without CoT.
+However, in-context mitigations can only help so much.
+In Figure 4, we plot their performance under various hyperparameters, on Gpt-4o.
+While $k$-means (left) is almost as bad as random guessing, $k$-nearest (center) achieves $\approx 50\%$ $\mathrm{FracCorrect}$, and the $k$-means, $k^{\prime}$-nearest mitigation (right) obtains $\approx 60\%$.
+The latter two out-perform unmitigated Gpt-4o, but fall significantly short of the linear baseline.
 CB exploit puzzles (text-based & non-linear).
@@
 Rewards are still presented numerically, and are non-linear functions of both the context and action.
-See Appendix A for full details on our experimental setup.
-Figure 5 (right) shows the performance of Gpt-4o (with mitigations) on this puzzle.
+See Appendix B for full details on our experimental setup.
+Figure 5 shows the performance of Gpt-5.2 (left) and Gpt-4o (right) with mitigations on this puzzle.
 While the reward function is non-linear (and thus the linear baseline only achieves  $70\%$ $\mathrm{FracCorrect}$), we find that all configurations are still significantly out-performed by the linear baseline.
-Our intuition for these findings is as follows: While LLMs are generally good with textual inputs, the reward data is numerical, so the LLM faces most/all of the same challenges as before (i.e.,  choosing a “solution strategy” and executing it correctly), exacerbated by non-linearity.
+Interestingly, Gpt-4o with code interpreter access (orange) performs the worst out of all configurations.
+Our intuition for these findings is as follows: While LLMs are generally good with textual inputs, the reward data is numerical, so the LLM faces same challenges as before (choosing a “solution strategy” and executing it correctly), exacerbated by non-linearity.
 Meanwhile, linear regression is known to often work fairly well even on non-linear data.
 Takeaways.
-While unmitigated performance drops as $d,K,T$ grow, increases in $T$ can largely be taken care of by our mitigations, whereas increases in $d$ and $K$ cannot.
-Indeed, taken together, our findings in this section suggest that while the current generation of LLMs are effective at exploiting in small-scale decision-making tasks, they struggle in larger settings.
+Frontier models can now solve fairly complicated MAB exploit puzzles.
+In contrast, Gpt-4 and Gpt-3.5 struggle, partially explaining their failure on end-to-end MAB tasks in Krishnamurthy et al.
+(2024); Nie et al.
+(2024).
+In numerical CB puzzles, unmitigated performance drops as $d,K,T$ grow, but tool use can help and increases in $T$ can largely be taken care of by our in-context mitigations.
+In contrast, our mitigations were largely unhelpful in a non-linear, text-based CB puzzle.
+Taken together, our findings suggest that while current LLMs are effective at exploiting in small-scale, numerical decision-making tasks, they still struggle in more complicated settings.
+Figure 5: Performance of the Gpt-5.2 (left) and Gpt-4o on a text-based CB puzzle. In this more challenging non-linear setting, some mitigations help, but all are outperformed by linear regression.
 3 LLMs as exploration oracles
-Figure 6: Algorithm’s average expected reward $\overline{\mathtt{rew}}(\text{task},K)$ (averaged over rounds and over runs), against $K$, the number of candidates.
-Each line corresponds to a prompting strategy or the $\mathtt{Random}$ baseline.
-The shaded regions represent a $95\%$ confidence interval.
 We now study the ability of LLMs to explore large action spaces.
@@
 We refer to this LLM usage as exploration oracle.
-Our usage of LLMs for exploration is closely aligned with the idea of “discretization” in the study of dynamic pricing and Lipschitz bandits (e.g.,  Kleinberg and Leighton, 2003; Kleinberg et al., 2008, 2019; Bubeck et al., 2011; Slivkins, 2014).
-Faced with a very large action space, it may be advantageous to focus on a much smaller set of candidate actions.
-A naively constructed discretization, e.g., uniform with respect to some known notion of distance in the action space, may have a poor tradeoff between size (#candidates) and quality (of the best candidate), placing all but a few candidates in low-performing regions of the action space.
-A crucial theme in this literature is “smarter” discretization algorithms which gradually “zoom in” on better-performing regions of the action space.
-The intuition is that if the better-performing arms comprise a low-dimensional region of a high-dimensional action space, covering this region should require much fewer samples for the same “discretization quality”.
+This usage of LLMs is closely aligned with “discretization” in dynamic pricing and Lipschitz bandits (e.g.,  Kleinberg and Leighton, 2003; Kleinberg et al., 2008, 2019; Bubeck et al., 2011; Slivkins, 2014).
+Given a very large action space, it may help to focus on a much smaller set of candidate arms.
+A naively constructed discretization, e.g., uniform w.r.t.
+some known distance notion on the arms, may have a poor tradeoff between size (#candidates) and quality (of the best candidate), placing all but a few candidates in low-performing regions of the action space.
+A crucial theme in this literature is “smarter” discretization algorithms which gradually “zoom in” on better-performing arms.
+If the latter comprise a low-dimensional region of the action space, covering this region should require much fewer samples for the same “discretization quality”.
 Our hope is that LLMs can suggest candidate actions from this region, if there is sufficient semantic meaning encoded in the bandit task.
-We consider two types of exploration tasks: answering an open-ended “philosophical” question (Q/A task) and suggesting a title for an arXiv research paper based on its abstract (arXiv task).
+Non-LLM approaches appear hopeless for our tasks.
+Standard bandit algorithms (Thompson Sampling, UCB1, etc.) fail for a very large #arms, and so do bandit algorithms which randomly subsample the arms.
+Another approach embeds the arms’ text labels in $\mathbb{R}^{d}$.
+If a suitable reward-distance relation were known to the algorithm/agent (assumption we do not make), one could apply the “smarter” discretization algorithms mentioned above.
+However, their regret scales exponentially in $d$, making them impractical for state-of-art text embeddings (e.g., one we use has $d=384$).
+Our tasks.
+We consider stylized exploration tasks of the following form.
+In each task, we define an MAB instance with Bernoulli rewards and a very large #arms, each labeled with a text string.
+The mean rewards $\mu(\cdot)$ and the meaning of the arms’ labels are based on real-world data.
+We call an LLM once to generate a small representative subset of $K$ arms.
+To grade the entire set (not just the best arm therein), we run a standard bandit algorithm (UCB1; Auer et al.
+(2002)) over these $K$ arms, for $T=1000$ rounds.
+We record the algorithm’s average cumulative reward, “denoised” as $\mathtt{rew}:=\tfrac{1}{T}\,\sum_{t\in[T]}\mu(a_{t})$.
+We consider three task types: (1) recommend a movie, based on the MovieLens 100K dataset Harper and Konstan (2015), a standard benchmark for evaluating CB (MovieLens task), (2) suggest a title for an arXiv research paper given its abstract (arXiv task), and (3) answer an open-ended “philosophical” question (Q/A task).
 Particular workloads within each task type are called explore puzzles.
-The Q/A task (resp.
-arXiv task) is constructed as follows.
-We define the “best arm” $a^{*}$ as a contrarian answer generated by another LLM (resp.
-the actual title of the research paper).
-The expected reward $\mu(a)$ of an arm $a$ is the cosine similarity between $a$ and $a^{*}$ in the embedding space.
+Our findings are similar across all task types and all LLMs.
+The MovieLens 100K dataset contains ratings from 943 users across 1,682 movies.
+In our tasks, arms correspond to the movies titles.
+Each task corresponds to a particular user, with mean rewards given by this user’s ratings.
+Since every user did not rate every movie, we used a standard algorithm to complete the data matrix (the soft-impute low-rank matrix completion method from the fancyimpute Python package).
+We then normalize this matrix so that all values lie $[0,1]$.
+The arXiv task (resp.
+Q/A task) is constructed as follows.
+Each task corresponds to an abstract-title pair from arXiv (resp., a question-answer pair).
+The arms are all possible titles (resp., answers).
+The “best arm” $a^{*}$ is the actual title (resp.
+a contrarian answer generated by another LLM).
+The expected reward $\mu(a)$ of an arm $a$ is the cosine similarity between $a$ and $a^{*}$ in the embedding generated by the Sentence-BERT embedding model Reimers (2019).
 [^8] While cosine similarity ranges on [1,1][-1,1], it was usually strictly positive in our experiments.
 In the (very rare) cases where it was negative we defined the expected reward as zero.
-Here, we generate sentence embeddings using the Sentence-BERT embedding model Reimers (2019).
-[^9] Our results are largely unchanged under the Universal Sentence Encoder Cer (2018), see Appendix B.
-The realized reward in a given round is an independent Bernoulli sample with the mean defined above.
-Since action spaces are extremely large for these tasks, standard bandit algorithms (e.g.,  Thompson Sampling, UCB1, etc.) catastrophically fail, and so do bandit algorithms which randomly subsample the action space.
-If the embedding space, the distance notion therein, and the reward-distance relation were known to the algorithm/agent (assumptions we do not make), one could, in principle, apply the “smarter” discretization algorithms mentioned above.
-However, the regret of these algorithms scales exponentially in the dimension $d$ of the metric space making them impractical in high-dimensional action spaces such as ours.
-(Indeed, our embedding space has $d=384$.)
-Like for our exploit puzzles, we highlight our results for GPT models in the main body and include the rest in Appendix B.5. Our findings are largely consistent across different LLMs.
-Explore puzzle: open-ended questions.
-We used Gpt-4 to generate a dataset of $10$ open-ended questions with many reasonable answers, along with an intentionally contrarian answer for each question to serve as the ground truth.
-(E.g.,  “What does it mean to live a fulfilling life?” “Fulfillment comes from embracing discomfort.”) Each question-answer pair yields a task, as defined above.
-We evaluate Gpt-4o, Qwen-2.5, Gemma-3, and Mistral-7B as exploration oracles.
-We prompt them to suggest $K\in\{1,2,3,4,5,7,10\}$ candidate answers given a question.
-To grade the entire candidate set (not just the best answer), we run an off-the-shelf MAB algorithm for some large-but-realistic time horizon $T$ over these $K$ candidates.
-(We use the UCB1 algorithm (Auer et al., 2002) and $T=1000$.) We record the expected time-averaged reward $\mathtt{rew}:=\tfrac{1}{T}\,\sum_{t\in[T]}\mu(a_{t})$, where $a_{t}$ is the arm selected in round $t$.
-We repeat this process (candidate selection and running UCB1) 10 times for a given task and $K$, and record the average $\mathtt{rew}$ over these runs, denoted by $\overline{\mathtt{rew}}=\overline{\mathtt{rew}}(\text{task},K)$.
-Figure 7: Arm histogram for one particular question (“What is the meaning of freedom?”, left) and averaged over all questions (right).
-We consider $K=10$ suggestions generated by our one-by-one prompt without encouragement.
-For a given “run”, the $K$ suggestions are ranked by the expected reward, and then the $i$-th expected reward is averaged over all runs, for each $i\in[K]$.
-We try several prompting strategies.
-We either ask the LLM to generate suggestions “all-at-once” with temperature 0 or “one-by-one” with temperature $1$ (we repeatedly show the LLM the list of candidate answers so far and ask it to generate one more).
-We also experiment with explicitly prompting the LLM to provide a diverse set of candidate answers (“with encouragement”).
-Thus, we have 4 prompting strategies: $\left\{\,\text{all-at-once, one-by-one}\,\right\}\times\left\{\,\text{with, without}\,\right\}$ encouragement.
-To compare against not using an LLM as an exploration oracle, we also consider a baseline ($\mathtt{Random}$) in which the candidates are $K$ points selected independently and u.a.r.
-in the embedding space.
-We visualize our findings in Figure 6.
-We plot $\overline{\mathtt{rew}}(\text{task},K)$ against $K$, for one particular task (left) and averaged across all tasks (right).
-[^10] See Appendix B for similar plots for the $9$ other tasks.
-Each line corresponds to a particular prompting strategy or the $\mathtt{Random}$ baseline.
-We find that all four prompting strategies perform similarly, with average reward between $0.5$ and $0.6$, and typically peak in performance around $K=3$ or $K=4$ suggestions.
-In contrast, the $\mathtt{Random}$ baseline catastrophically fails, with its average reward never exceeding $0.1$.
-We conclude that the LLM does succeed as an exploration oracle.
-We observe that the LLM-generated suggestions pass the “eye test”, in the sense that we get reasonable, yet spiritually and semantically different answers for a given question.
-For example, given the question “What is the role of technology in society”, the first $K=5$ suggestions generated by our one-by-one prompt are as follows: (1) Facilitates communication, innovation, and efficiency, (2) Transforms daily life and shapes culture, (3) Drives connectivity and enhances productivity, (4) Facilitates control and surveillance, (5) Disrupts traditional relationships and norms.
-Finally, we verify that the candidate suggestions are substantially different from one another.
-Figure 7 visualizes the spread of expected rewards within the candidate set.
-We consider $K=10$ suggestions generated by our one-by-one prompt without encouragement.
-For a given “run” of the candidate selection, the suggestions are ranked by the expected reward, and then the $i$-th expected reward is averaged over all runs, for each $i\in[K]$.
-A more detailed comparison between the prompting strategies is not fruitful for these experiments, as sentence embeddings are known to be somewhat imprecise at small scales (see Section B.4).
-Figure 8: Algorithm’s average expected reward $\overline{\mathtt{rew}}(\text{category},K)$ (averaged over rounds and tasks), against $K$, the number of candidates.
+[^9] Our results are largely unchanged under the Universal Sentence Encoder Cer (2018), see Section C.4.
+As a robustness check, we try several prompting strategies (detailed in Appendix C.1).
+We either ask the LLM to generate suggestions “all-at-once” with temperature 0 or “1-by-1” with temperature $1$ (we repeatedly show the LLM the candidate arms so far and ask it to generate one more).
+We also try to explicitly encourage the LLM to provide a diverse candidate set.
+Thus, we have 4 prompting strategies: $\left\{\,\text{all-at-once, 1-by-1}\,\right\}\times\left\{\,\text{with, without}\,\right\}$ encouragement.
+However, all 4 strategies performed similarly.
+[^10] We only tried the 4 strategies for the arXiv and Q/A tasks.
+Figure 6: Average expected reward $\overline{\mathtt{rew}}(\text{category},K)$ (averaged over rounds and tasks), against $K$, the number of candidates.
 Each line corresponds to a prompting strategy or the Category-Only baseline.
@@
 A single arXiv category (“General Relativity and Quantum Cosmology”, left), averages over 6 categories (right).
-Explore puzzle: arXiv abstracts and titles.
-We run similar experiments on a larger-scale dataset of paper titles and abstracts from arXiv.org.
-Using the arXiv API (arXiv.org, 2025), we collect $10$ abstract-title pairs from each of the $41$ different arXiv categories.
-To minimize the likelihood that these papers appear in Gpt-4o’s training corpus, we only use papers uploaded after June 2024 (these experiments were done in Sept’24).
-Each abstract-title pair yields a task, as discussed earlier.
-We evaluate Gpt-4o as an exploration oracle for these tasks, much like for the Q/A tasks Given an abstract, we prompt Gpt-4o to generate $K$ alternative titles, which are then used to instantiate a bandit algorithm.
-We use the same algorithm (UCB1) and time horizon $T=1000$.
-We record the expected time-averaged reward, $\mathtt{rew}:=\tfrac{1}{T}\,\sum_{t\in[T]}\mu(a_{t})$, and compute the average over tasks within the same arXiv category, $\overline{\mathtt{rew}}(\text{category},K)$.
-We try “all-at-once” and “one-by-one” prompting.
-[^11] We do not use “encouragement” on these tasks, since it does not help much (if at all) for the Q/A tasks.
-We visualize our findings in Figure 8 , using the same conventions as Figure 6 and focusing on six arXiv categories.
-[^12] General relativity, computer vision, statistics theory, biomolecules, signal processing, general economics.
-To assess LLM’s ability to specialize to a task, we consider a stronger baseline, where the candidate arms are generated by Gpt-4o given only the category, not the abstract.
-For the other arXiv categories, we provide a table with $\overline{\mathtt{rew}}(\text{category},K)$ for $K=1,2,5$ and both prompts (Appendix B).
-Like for the Q/A tasks, (1) both prompting strategies significantly outperform the baseline, and (2) their performance tends to improve modestly as $K$ increases.
+Explore puzzle: MovieLens.
+We run a MovieLens task for each of the $943$ users, with $K\in\{10,19\}$.
+We re-run each task 5 times to denoise.
+For a given $K$, we record $\overline{\mathtt{rew}}(K)$: reward $\mathtt{rew}$, defined above, averaged over all users and runs.
+For baselines, we compare against running UCB1 for $T$ rounds over (1) a random subset of $K$ movies and (2) a random movie from each of the $19$ movie genres in the dataset.
+(Different random draws for each user.)
+Our results for Mistral-7B and Qwen-2.5 are in Table 1 Both LLMs significantly outperform all baselines for both values of $K$.
+We also tried running the experiment using Llama-3-8B and Gemma-3, but the list of all movies was too long to fit in the context window.
+Table 1: Average Reward for Qwen-2.5, Mistral-7B, and Baselines for $K\in\{10,19\}$ in the MovieLens Task.
+$\overline{\mathtt{rew}}$ $K=10$ $K=19$
+Qwen-2.5 $671.88$ $\mathbf{678.25}$
+Mistral-7B $\mathbf{735.38}$ $675.16$
+Random $570.39$ $599.73$
+Genre-based $617.27$
+The LLM has implicit knowledge about each movie which it may leverage when selecting the discretization.
+To mitigate overfitting to “popular” movies, we prompted the LLM to not take movie popularity into account.
+However, this change did not significantly affect the rewards.
+Table 2: Performance using all-at-once prompting on select arXiv tasks. Results are averaged over $10$ abstract-title pairs. The full table (Table 3) may be found in the Appendix.
+$K=1$ $K=2$ $K=5$
+gr-qc 0.63 0.64 0.68
+hep-ex 0.81 0.81 0.83
+hep-th 0.65 0.71 0.73
+math-ph 0.64 0.73 0.74
+nucl-ex 0.73 0.79 0.75
+Explore puzzle: arXiv.
+Using the arXiv API (arXiv.org, 2025), we collected $10$ abstract-title pairs from each of the $41$ arXiv categories.
+[^11] To minimize the likelihood that these papers appear in the training corpus, we only use papers uploaded after June’24 (these experiments were done in Sept’24).
+Each abstract-title pair yields a task, with $K=1,2,5$ suggestions.
+Each task is re-run 5 times to denoise.
+Given $K$, we record $\overline{\mathtt{rew}}(\text{category},K)$: reward $\mathtt{rew}$, as defined above, averaged over all tasks within the same arXiv category and all runs.
+We evaluate Gpt-4o as an exploration oracle for these tasks.
+We visualize our findings in Figure 6 , focusing on six arXiv categories.
+To assess the LLM’s ability to specialize to a task, we consider a baseline where the candidate arms are generated by Gpt-4o given only the category, not the abstract.
+For the other arXiv categories, we provide a table with $\overline{\mathtt{rew}}(\text{category},K)$ for $K=1,2,5$ (Table 2, Appendix C).
+Explore puzzle: Q/A task.
+We used Gpt-4 to generate a dataset of $10$ open-ended questions with many reasonable answers, along with an intentionally contrarian answer for each question to serve as the ground truth.
+(E.g.,  “What does it mean to live a fulfilling life?” “Fulfillment comes from embracing discomfort.”) Each question-answer pair yields a task, with $K\in\{1,2,3,4,5,7,10\}$ candidate answers.
+For a given $K$, we run each task 10 times and record $\overline{\mathtt{rew}}(\text{task},K)$: reward $\mathtt{rew}$ averaged over all runs.
+We evaluate Gpt-4o, Qwen-2.5, Gemma-3, and Mistral-7B as exploration oracles.
+We also consider a non-LLM baseline ($\mathtt{Random}$), which selects $K$ arms independently and u.a.r.
+in the embedding space (leveraging the AI-based embedding but suffering from its high dimension).
+The LLMs perform similarly, with $\overline{\mathtt{rew}}\approx$0.5-0.6.
+and typically work best for KK=3,4.
+see Appendix A.
+In contrast, $\mathtt{Random}$ catastrophically fails, with $\overline{\mathtt{rew}}<0.1$.
+Moreover, we observe that LLM-generated answers pass the “eye test” (they are reasonable, yet spiritually and semantically different), and yield substantially different expected rewards.
+We conclude that the LLM does succeed as an exploration oracle for these tasks.
 4 Conclusions
@@
 We find that LLMs are useful as exploration oracles that propose high-quality candidate actions in large, semantically meaningful action spaces.
-However, current LLMs are not that good at exploitation, particularly in larger or more complex tasks (we trace the “boundary of tractability” in terms of task size/complexity).
-While we suggest several helpful mitigations, they consistently underperform relative to a simple linear regression, even in inherently non-linear tasks.
+However, current LLMs are not that good at exploitation, particularly in larger or more complex tasks, although reasoning models (specifically, the Gpt-5 family of models) show the most promise.
+While we experiment with several mitigations including tool use and in-context summaries, they consistently underperform relative to a simple linear regression baseline, even in inherently non-linear tasks.
+We provide a systematic evaluation of LLMs’ capabilities as exploration oracles, in isolation from other components of a decision-making agent.
+We find robust success across several different task types, LLMs, and prompting strategies.
+In contrast, non-LLM baselines are often quite weak, absent known problem structure to be leveraged by an algorithm.
+LLMs’ continued difficulties at exploitation tasks should be contrasted with the rapid progress in LLMs and the extreme optimism associated with this progress.
+Ever since the Gpt-3.5 release in 2022 (and especially since Gpt-4 frenzy of early 2023), it was commonly believed that “generalist” frontier LLMs will soon master many/most relatively simple tasks: if not yet, then more likely in several months than in several years.
+However, our results show that these difficulties persist longer than anticipated.
 Future work.
-LLMs trained to use tools like a calculator may be better at exploitation.
-However, it is unclear how much this would help in more complex scenarios, e.g.,  CB tasks with text-based contexts and actions, and which mitigations/prompting techniques would be needed.
-Second, while ”zooming” bandit algorithms do not work for rich text-based action spaces (discussed in Section 3), LLM-based exploration oracles may potentially help.
-The hope is to “zoom in” entirely in the space of “potentially relevant” actions (as determined by the LLM), rather than in the space of all actions.
-Appendix A Appendix for Section 2: LLMs as exploitation oracles
-A.1 Prompts
+LLM-based exploration oracles may potentially help with “smart discretization” for rich text-based action spaces.
+Non-LLM bandit algorithms tend to fail in such scenarios, as discussed in Section 3.
+The hope is to “zoom in” entirely in the space of “potentially relevant” actions, rather than in the space of all actions.
+Second, while we experiment with tool use for exploitation tasks, we find that it does not improve performance across the board.
+One could investigate various approaches for tool use in these tasks, and whether/when they help.
+Acknowledgements
+We would like to thank Dylan Foster and Akshay Krishnamurthy for helpful conversations at early stages of the project.
+Appendix A Main Plots for the Q/A explore task
+Here we provide the key plots for the Q/A explore task, as defined in Section 3.
+As stated there, all LLMs we considered perform similarly, with $\overline{\mathtt{rew}}\approx$0.5-0.6.
+and typically work best for KK=3,4, whereas the $\mathtt{Random}$ baseline catastrophically fails, with $\overline{\mathtt{rew}}<0.1$.
+Recall that $\overline{\mathtt{rew}}(\text{task},K)$ is the average reward $\mathtt{rew}$ for a given task with a given $K$ (as defined in Section 3), averaged over all runs of this task.
+We visualize our findings in Figure 7.
+We plot $\overline{\mathtt{rew}}(\text{task},K)$ against $K$, for one particular task (left) and averaged across all tasks (right).
+Similar plots for the $9$ other tasks can be found in Section C.6.
+Figure 7: Average expected reward $\overline{\mathtt{rew}}(\text{task},K)$ (averaged over rounds and over runs), against $K$, #candidates.
+Each line corresponds to a prompting strategy or the $\mathtt{Random}$ baseline.
+The shaded regions represent a $95\%$ confidence interval.
+The LLM-generated answers pass the “eye test”: they are reasonable, yet spiritually and semantically different.
+E.g., given the question “What is the role of technology in society”, the first $K=5$ suggestions generated by our 1-by-1 prompt are: (1) Facilitates communication, innovation, and efficiency, (2) Transforms daily life and shapes culture, (3) Drives connectivity and enhances productivity, (4) Facilitates control and surveillance, (5) Disrupts traditional relationships and norms.
+Finally, we verify that the candidate suggestions are substantially different from one another in terms of their rewards, see Figure 8.
+Figure 8: Spread of expected rewards within the candidate set.
+Arm histogram for one particular question (“What is the meaning of freedom?”, left) and averaged over all questions (right).
+We consider $K=10$ suggestions generated by our one-by-one prompt without encouragement.
+For a given “run”, the $K$ suggestions are ranked by the expected reward, and then the $i$-th expected reward is averaged over all runs, for each $i\in[K]$.
+Appendix B Details for Section 2: LLMs as exploitation oracles
+B.1 Prompts
 In this section we give example prompts for each of our experimental setups.
@@
 Remember, you must provide your final answer immediately within the tags <Answer>ACTION <Answer>where ACTION is one of ”pet animal”, ”leave room”, ”use tool”, ”eat food”, or ”press button” and with no text explanation.
-A.2 Additional MAB figures
-See Figure 9Figure 10Figure 11 for additional results in our MAB exploit puzzle.
+B.2 Additional Results on MAB exploit puzzle
 Figure 9: Cumulative fraction correct for Gpt-4 in the MAB adverts puzzle.
@@
 Figure 11: Cumulative fraction correct for Gpt-3.5 in the MAB buttons puzzle.
-A.3 Additional details for text-based CB puzzles
+B.3 Additional details for text-based CB puzzles
 Each context contains a time of day (belonging to {morning, afternoon, evening, night}), an animal ({bear, dog, cat, None}), a tool ({key, letter opener, hammer, None}), a food item ({cake, apple, nut, None}), and a button with a particular color ({red, orange, yellow, green}).
-The actions in each round are “pet animal”, “leave room”, “use tool”, “eat food”, and “press button”.
-We experimented with two reward functions: an “easy” reward function, where the expected rewards for each action are as follows:
-  • The expected reward of petting the animal is 0.01 if the animal is a bear, 0.7 if the animal is a dog, and 0.4 if the animal is a cat. Otherwise, the expected reward if 0.5.
-  • The expected reward for leaving the room is always 0.5.
-  • The expected reward for using the tool is 0.75 if it is a key, 0.6 if it is a letter opener, 0.45 if it is a hammer, and 0.2 otherwise.
-  • The expected reward for eating food is 0.8 if it is cake, 0.6 if it is an apple, 0.2 if it is a nut, and 0.3 otherwise.
-  • The expected reward for pressing the button is 0.89 if it is green, 0.62 if it is yellow, 0.39 if it is orange, and 0.27 if it is red.
+The actions are “pet animal”, “leave room”, “use tool”, “eat food”, and “press button”.
+We considered two reward functions. One was “easy”, with expected rewards (ExpRev) as follows:
+  • ExpRev for petting the animal is 0.01 if the animal is a bear, 0.7 if the animal is a dog, and 0.4 if the animal is a cat. Otherwise, ExpRev if 0.5.
+  • ExpRev for leaving the room is always 0.5.
+  • ExpRev for using the tool is 0.75 if it is a key, 0.6 if it is a letter opener, 0.45 if it is a hammer, and 0.2 otherwise.
+  • ExpRev for eating food is 0.8 if it is cake, 0.6 if it is an apple, 0.2 if it is a nut, and 0.3 otherwise.
+  • ExpRev for pressing the button is 0.89 if it is green, 0.62 if it is yellow, 0.39 if it is orange, and 0.27 if it is red.
 Our results under this reward function are summarized in Figure 12.
@@
 The reward function we use in the main body is more complicated, and is detailed below:
-  • The expected reward for petting the animal is 0.01 if it is a bear, 0.7 if it is a dog, 0.3 if it is a cat and the time of day is morning or afternoon, 0.7 if it is a cat and the time of day is evening or night, and otherwise 0.5.
-  • The expected reward for leaving the room is always 0.5
-  • If the animal is a bear, the expected reward for using the tool is 0.1. Otherwise, if the tool is a key and the table item is a chest, the expected reward is 0.9. Otherwise, it is 0.4
-  • If the animal is a bear, the expected reward for eating food is always 0.5. Otherwise, the expected reward is 0.8 for cake, 0.6 for an apple, 0.2 for a nut, and 0.5 otherwise.
-  • If the animal is a bear, the expected reward for pressing the button is 0.1.
-Otherwise if the button is green and the time of day is morning, or the button is yellow and the time of day is afternoon, or the button is orange and the time of day is evening, or the button is red and the time of day is night, then the expected reward is 0.9.
-In all other cases, the expected reward is 0.25.
-A.4 Robustness Check with Other Models
+  • ExpRev for petting the animal is 0.01 if it is a bear, 0.7 if it is a dog, 0.3 if it is a cat and the time of day is morning or afternoon, 0.7 if it is a cat and the time of day is evening or night, and otherwise 0.5.
+  • ExpRev for leaving the room is always 0.5
+  • If the animal is a bear, ExpRev for using the tool is 0.1. Otherwise, if the tool is a key and the table item is a chest, ExpRev is 0.9. Otherwise, it is 0.4
+  • If the animal is a bear, ExpRev for eating food is always 0.5. Otherwise, ExpRev is 0.8 for cake, 0.6 for an apple, 0.2 for a nut, and 0.5 otherwise.
+  • If the animal is a bear, ExpRev for pressing the button is 0.1.
+Otherwise if the button is green and the time of day is morning, or the button is yellow and the time of day is afternoon, or the button is orange and the time of day is evening, or the button is red and the time of day is night, then ExpRev is 0.9.
+In all other cases, ExpRev is 0.25.
+B.4 Robustness Check with Other Models
 In this appendix, we include results for Qwen2.5-7B-Instruct (Qwen-2.5), Gemma 3 12B (Gemma-3), and Mistral-7B-Instruct-v0.3 (Mistral-7B) on our exploit puzzles.
@@
 Figure 22: Results for Gemma-3 in the setting of Figure 19. Only mitigations are shown, as the full history could not fit in the context window.
-Appendix B Appendix for Section 3: LLMs as exploration oracles
-B.1 Prompts
-In this section we give example prompts for each of our experimental setups.
-“All-at-once” prompt with encouragement for open-ended question puzzle:
+Figure 23: Performance of DeepSeek-R1 on our numerical CB puzzle.
+Appendix C Details for Section 3: LLMs as exploration oracles
+C.1 Prompts
+Prompt for the MovieLens experiments:
+[SYSTEM] You are a movie expert helping a user choose a movie.
+[USER] Here is a list of movies with their numeric IDs: …
+From this list, choose {K} movies.
+You don’t know what taste in movies the user has, so select a diverse set of movies from different genres such that they will most likely enjoy at least one of the movies you select.
+Respond ONLY with the {K} numeric IDs, one per line, with no extra text.
+Do NOT consider a movie’s popularity when deciding whether to select it.
+For the arXiv and Q/A tasks, we used a 2x2 space of prompts, as discussed in Section 3: $\left\{\,\text{all-at-once, 1-by-1}\,\right\}\times\left\{\,\text{with, without}\,\right\}$ encouragement.
+We provide two illustrative examples below.
+“All-at-once” prompt with encouragement for the Q/A task:
 [SYSTEM] I will give you an open-ended question.
@@
 [USER] Here is the question: “What is the purpose of art?”
-“One-by-one” prompt without encouragement for arXiv puzzle:
+“One-by-one” prompt without encouragement for the arXiv task:
 [SYSTEM] I will give you an abstract and some candidate titles for a paper.
@@
 Here are the other candidate titles: {previous suggestions go here}
-B.2 Datasets
+C.2 Dataset: Q/A task
 Our open-ended question dataset consists of the following $10$ questions and the corresponding “ground-truth” answers.
@@
     Education’s purpose is to challenge authority.
+C.3 Dataset: arXiv task
 Here is the list of paper titles we used in our arXiv dataset, along with their corresponding categories:
@@
     Emulating the Global Change Analysis Model with Deep Learning
-B.3 Additional results for explore puzzles
+C.4 Robustness Check: Q/A task with another embedding model
 We ran a robustness check on the first six open-ended question experiments using the universal sentence encoder of Cer [2018] as our embedding model.
@@
 Figure 24: Results averaged over the first six questions, for embeddings generated using the universal sentence encoder.
-Below are the individual plots for the remaining $9$ questions using the Sentence-BERT encoder.
-Figure 24: Results on Question 1
-Figure 25: Results on Question 2
-Figure 26: Results on Question 3
-Figure 27: Results on Question 4
-Figure 28: Results on Question 5
-Figure 29: Results on Question 6
-Figure 30: Results on Question 7
-Figure 31: Results on Question 8
-Figure 32: Results on Question 9
-Table 1: Performance comparison for all-at-once on open-ended questions.
-K=1 K=2 K=3 K=4 K=5 K=7 K=10
-Q0: 0.39 0.4 0.4 0.4 0.41 0.42 0.56
-Q1: 0.63 0.66 0.66 0.65 0.63 0.64 0.57
-Q2: 0.41 0.43 0.41 0.4 0.36 0.41 0.41
-Q3: 0.5 0.49 0.6 0.67 0.67 0.66 0.78
-Q4: 0.33 0.46 0.4 0.43 0.38 0.42 0.42
-Q5: 0.39 0.41 0.47 0.3 0.31 0.42 0.43
-Q6: 0.45 0.78 0.78 0.79 0.78 0.78 0.75
-Q7: 0.84 0.84 0.84 0.84 0.84 0.83 0.83
-Q8: 0.39 0.41 0.47 0.36 0.37 0.37 0.39
-Q9: 0.47 0.46 0.43 0.46 0.48 0.45 0.44
-Table 2: Performance comparison for all-at-once with encouragement on open-ended questions.
-K=1 K=2 K=3 K=4 K=5 K=7 K=10
-Q0: 0.4 0.4 0.34 0.35 0.37 0.52 0.55
-Q1: 0.45 0.51 0.66 0.61 0.58 0.56 0.59
-Q2: 0.36 0.4 0.41 0.4 0.4 0.37 0.43
-Q3: 0.48 0.42 0.41 0.49 0.4 0.6 0.49
-Q4: 0.5 0.52 0.52 0.51 0.5 0.51 0.49
-Q5: 0.4 0.39 0.39 0.43 0.45 0.4 0.41
-Q6: 0.74 0.67 0.79 0.79 0.77 0.55 0.64
-Q7: 0.69 0.82 0.78 0.79 0.79 0.77 0.8
-Q8: 0.45 0.5 0.47 0.56 0.55 0.42 0.51
-Q9: 0.44 0.39 0.37 0.4 0.47 0.46 0.47
-Table 3: Performance comparison for one-by-one on open-ended questions.
-K=1 K=2 K=3 K=4 K=5 K=7 K=10
-Q0: 0.33 0.4 0.4 0.43 0.43 0.47 0.52
-Q1: 0.62 0.62 0.62 0.62 0.61 0.61 0.6
-Q2: 0.43 0.43 0.44 0.43 0.43 0.44 0.45
-Q3: 0.58 0.57 0.59 0.61 0.59 0.6 0.6
-Q4: 0.36 0.41 0.47 0.48 0.47 0.46 0.45
-Q5: 0.39 0.48 0.49 0.48 0.48 0.49 0.49
-Q6: 0.83 0.81 0.82 0.81 0.81 0.8 0.79
-Q7: 0.8 0.81 0.81 0.81 0.81 0.8 0.79
-Q8: 0.38 0.42 0.44 0.44 0.47 0.5 0.5
-Q9: 0.45 0.47 0.47 0.46 0.45 0.45 0.43
-Table 4: Performance comparison for one-by-one with encouragement on open-ended questions.
-K=1 K=2 K=3 K=4 K=5 K=7 K=10
-Q0: 0.35 0.37 0.5 0.54 0.55 0.58 0.59
-Q1: 0.55 0.58 0.6 0.61 0.61 0.61 0.62
-Q2: 0.45 0.45 0.44 0.44 0.45 0.45 0.42
-Q3: 0.55 0.55 0.54 0.53 0.53 0.53 0.54
-Q4: 0.33 0.49 0.48 0.49 0.49 0.48 0.46
-Q5: 0.4 0.43 0.44 0.44 0.45 0.45 0.45
-Q6: 0.83 0.83 0.82 0.82 0.81 0.81 0.8
-Q7: 0.75 0.8 0.81 0.82 0.81 0.82 0.82
-Q8: 0.35 0.4 0.45 0.47 0.49 0.5 0.53
-Q9: 0.49 0.49 0.5 0.49 0.5 0.47 0.48
-Table 5: Performance comparison for random actions on open-ended questions.
-K=1 K=2 K=3 K=4 K=5 K=7 K=10
-Q0: -0.01 -0.0 0.01 0.01 0.03 0.03 0.02
-Q1: 0.01 0.01 0.02 0.02 0.04 0.04 0.03
-Q2: 0.01 -0.0 0.03 0.02 0.02 0.03 0.02
-Q3: 0.05 0.04 0.01 0.05 0.04 0.03 0.04
-Q4: 0.02 0.01 0.03 0.02 0.02 0.04 0.01
-Q5: -0.02 -0.01 0.02 0.03 0.02 0.03 0.05
-Q6: 0.0 0.0 0.03 0.03 0.04 0.02 0.03
-Q7: 0.01 0.05 0.03 0.04 0.04 0.02 0.04
-Q8: -0.04 0.03 0.04 0.04 0.03 0.03 0.04
-Q9: 0.04 0.0 0.04 0.03 0.04 0.03 0.04
+C.5 Additional Results: arXiv task
+Figure 25: Results for other arXiv categories
 Table 3: Performance for all-at-once on arXiv tasks.
@@
 econ.GN 0.13 0.19 0.19
-Figure 33: Results for arXiv category econ.GN
-Figure 34: Results for arXiv category eess.SP
-Figure 35: Results for arXiv category q-bio.BM
-Figure 36: Results for arXiv category math.ST
-B.4 Benchmarking encoders
+C.6 Additional Results: QA task
+While previously we presented the plot for Question 0 (in Figure 7), below are the individual plots for the remaining $9$ questions, in the same notation.
+Throughout, we used the Sentence-BERT encoder.
+[Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image]
+Table 6: Performance comparison for all-at-once on open-ended questions.
+K=1 K=2 K=3 K=4 K=5 K=7 K=10
+Q0: 0.39 0.4 0.4 0.4 0.41 0.42 0.56
+Q1: 0.63 0.66 0.66 0.65 0.63 0.64 0.57
+Q2: 0.41 0.43 0.41 0.4 0.36 0.41 0.41
+Q3: 0.5 0.49 0.6 0.67 0.67 0.66 0.78
+Q4: 0.33 0.46 0.4 0.43 0.38 0.42 0.42
+Q5: 0.39 0.41 0.47 0.3 0.31 0.42 0.43
+Q6: 0.45 0.78 0.78 0.79 0.78 0.78 0.75
+Q7: 0.84 0.84 0.84 0.84 0.84 0.83 0.83
+Q8: 0.39 0.41 0.47 0.36 0.37 0.37 0.39
+Q9: 0.47 0.46 0.43 0.46 0.48 0.45 0.44
+Table 7: Performance comparison for all-at-once with encouragement on open-ended questions.
+K=1 K=2 K=3 K=4 K=5 K=7 K=10
+Q0: 0.4 0.4 0.34 0.35 0.37 0.52 0.55
+Q1: 0.45 0.51 0.66 0.61 0.58 0.56 0.59
+Q2: 0.36 0.4 0.41 0.4 0.4 0.37 0.43
+Q3: 0.48 0.42 0.41 0.49 0.4 0.6 0.49
+Q4: 0.5 0.52 0.52 0.51 0.5 0.51 0.49
+Q5: 0.4 0.39 0.39 0.43 0.45 0.4 0.41
+Q6: 0.74 0.67 0.79 0.79 0.77 0.55 0.64
+Q7: 0.69 0.82 0.78 0.79 0.79 0.77 0.8
+Q8: 0.45 0.5 0.47 0.56 0.55 0.42 0.51
+Q9: 0.44 0.39 0.37 0.4 0.47 0.46 0.47
+Table 8: Performance comparison for one-by-one on open-ended questions.
+K=1 K=2 K=3 K=4 K=5 K=7 K=10
+Q0: 0.33 0.4 0.4 0.43 0.43 0.47 0.52
+Q1: 0.62 0.62 0.62 0.62 0.61 0.61 0.6
+Q2: 0.43 0.43 0.44 0.43 0.43 0.44 0.45
+Q3: 0.58 0.57 0.59 0.61 0.59 0.6 0.6
+Q4: 0.36 0.41 0.47 0.48 0.47 0.46 0.45
+Q5: 0.39 0.48 0.49 0.48 0.48 0.49 0.49
+Q6: 0.83 0.81 0.82 0.81 0.81 0.8 0.79
+Q7: 0.8 0.81 0.81 0.81 0.81 0.8 0.79
+Q8: 0.38 0.42 0.44 0.44 0.47 0.5 0.5
+Q9: 0.45 0.47 0.47 0.46 0.45 0.45 0.43
+Table 9: Performance comparison for one-by-one with encouragement on open-ended questions.
+K=1 K=2 K=3 K=4 K=5 K=7 K=10
+Q0: 0.35 0.37 0.5 0.54 0.55 0.58 0.59
+Q1: 0.55 0.58 0.6 0.61 0.61 0.61 0.62
+Q2: 0.45 0.45 0.44 0.44 0.45 0.45 0.42
+Q3: 0.55 0.55 0.54 0.53 0.53 0.53 0.54
+Q4: 0.33 0.49 0.48 0.49 0.49 0.48 0.46
+Q5: 0.4 0.43 0.44 0.44 0.45 0.45 0.45
+Q6: 0.83 0.83 0.82 0.82 0.81 0.81 0.8
+Q7: 0.75 0.8 0.81 0.82 0.81 0.82 0.82
+Q8: 0.35 0.4 0.45 0.47 0.49 0.5 0.53
+Q9: 0.49 0.49 0.5 0.49 0.5 0.47 0.48
+Table 10: Performance comparison for random actions on open-ended questions.
+K=1 K=2 K=3 K=4 K=5 K=7 K=10
+Q0: -0.01 -0.0 0.01 0.01 0.03 0.03 0.02
+Q1: 0.01 0.01 0.02 0.02 0.04 0.04 0.03
+Q2: 0.01 -0.0 0.03 0.02 0.02 0.03 0.02
+Q3: 0.05 0.04 0.01 0.05 0.04 0.03 0.04
+Q4: 0.02 0.01 0.03 0.02 0.02 0.04 0.01
+Q5: -0.02 -0.01 0.02 0.03 0.02 0.03 0.05
+Q6: 0.0 0.0 0.03 0.03 0.04 0.02 0.03
+Q7: 0.01 0.05 0.03 0.04 0.04 0.02 0.04
+Q8: -0.04 0.03 0.04 0.04 0.03 0.03 0.04
+Q9: 0.04 0.0 0.04 0.03 0.04 0.03 0.04
+C.7 Robustness Check with Other Models
+We include results for Qwen2.5-7B-Instruct (Qwen-2.5), Gemma 3 12B (Gemma-3), and Mistral-7B-Instruct-v0.3 (Mistral-7B) on our explore Q/A puzzles, all in the setting of Figure 7.
+As was the case with the GPT models, performance increased with $K$ when the candidate answers are generated one-by-one.
+When generating answers all-at-once, we observed that Qwen-2.5 and Mistral-7B had trouble generating concise answers, leading to a drop in performance as $K$ increases.
+[Uncaptioned image][Uncaptioned image][Uncaptioned image]
+C.8 Benchmarking encoders
 Here we benchmark the two encoders we use (Sentence-BERT and the universal sentence encoder) by measuring the cosine similarity between semantically similar/different words.
@@
 The similarity scores of both models in Table 11 suggest that while the embeddings produced by both embedding models are generally “in the ballpark” of what one would consider “similar”/“different”, they are still a somewhat coarse measure of distance, which may explain the similar performance of our different prompting strategies.
-B.5 Robustness Check with Other Models
-In this appendix, we include results for Qwen2.5-7B-Instruct (Qwen-2.5), Gemma 3 12B (Gemma-3), and Mistral-7B-Instruct-v0.3 (Mistral-7B) on our explore Q/A puzzles.
-As was the case with the GPT models, performance increased with $K$ when the candidate answers are generated one-by-one.
-When generating answers all-at-once, we observed that Qwen-2.5 and Mistral-7B had trouble generating concise answers, leading to a drop in performance as $K$ increases.
-Figure 37: Results for Qwen-2.5 in the setting of Figure 6.
-Figure 38: Results for Mistral-7B in the setting of Figure 6.
-Figure 39: Results for Gemma-3 in the setting of Figure 6.
+C.9 Explore Experiments on the Movie Lens Dataset
+Prompt:
+[SYSTEM] You are a movie expert helping a user choose a movie.
+[USER] Here is a list of movies with their numeric IDs:{movie_list_str}
+From this list, choose K movies.
+You don’t know what taste in movies the user has, so select a diverse set of movies from different genres such that they will most likely enjoy at least one of the movies you select.
+Respond ONLY with the K numeric IDs, one per line, with no extra text.
+Do NOT consider a movie’s popularity when deciding whether to select it.